### v1

# Online Decision-Focused Learning

Abstract
Decision-focused learning (DFL) is an increasingly popular paradigm for training predictive models whose outputs are used in decision-making tasks. Instead of merely optimizing for predictive accuracy, DFL trains models to directly minimize the loss associated with downstream decisions. This end-to-end strategy holds promise for tackling complex combinatorial problems; however, existing studies focus solely on scenarios where a fixed batch of data is available and the objective function does not change over time. We instead investigate DFL in dynamic environments where the objective function and data distribution evolve over time. This setting is challenging because the objective function has zero or undefined gradients—which prevents the use of standard first-order optimization methods—and is generally non-convex. To address these difficulties, we (i) regularize the objective to make it differentiable and (ii) make use of the optimism principle, based on a near-optimal oracle along with an appropriate perturbation. This leads to a practical online algorithm for which we establish bounds on the expected dynamic regret, both when the decision space is a simplex and when it is a general bounded convex polytope.
1 CMAP, CNRS, École polytechnique
2 Inria Saclay, Université Paris Saclay, LMO
3 Inria Paris, Ecole Normale Supérieure, PSL
1 Introduction.
Many real-world decision problems involve uncertainty, and a common approach to handling it is through the predict-then-optimize framework bertsimas2020predictive . First, a prediction model is trained on historical data; then, its output is fed into an optimization problem to guide decision-making. This natural strategy has been successfully applied in many operation research (OR) problems, ranging from supply chain management acimovic2015making ; fisher2016optimal ; ban2019big ; bertsimas2020predictive and revenue management farias2013nonparametric ; ferreira2016analytics ; cohen2017impact ; chen2022statistical to healthcare operation bertsimas2013fairness ; aswani2019data ; gupta2020maximizing ; rath2017integrated , see mivsic2020data for an extensive review. It is clear that this approach would yield optimal decisions if the predictions were perfectly accurate. However, in practice, prediction errors are inevitable—and even small inaccuracies can propagate through the optimization process, potentially leading to poor decisions.
To address this limitation, an approach known as decision-focused learning mandi2024decision —also referred to as smart predict-then-optimize elmachtoub2022smart or integrated learning-optimization sadana2025survey —has emerged. Instead of optimizing for prediction accuracy alone, this method trains the predictive model to directly minimize the downstream decision loss. By aligning the learning objective with the decision-making goal, it produces models that are more robust to prediction errors in practical applications. While decision-focused learning yields strong empirical performance donti2017task ; verma2022case ; verma2023restless ; wang2023scalable , theoretical development has so far been limited to the batch setting, where models are trained on pre-collected, independently and identically distributed (i.i.d.) data wilder2018melding ; mandi2022decision ; shah2022decision ; schutte2024robust . This assumption breaks down in many real-world scenarios involving dynamic environments cheung2019learning ; padakandla2020reinforcement and shifting data distributions lu2018learning ; quinonero2022dataset .
Online learning cesa2006prediction ; hazan2023introductiononlineconvexoptimization provides a general way to cope with such non-stationarity of data-generating processes. This framework considers a learner who makes decisions sequentially, at each round leveraging data collected from previous rounds to inform its next decision. Crucially, the objective function is allowed to vary over time, either in a stochastic or adversarial way. This provides a natural framework for the work that we present here, which extends decision-focused learning beyond the i.i.d., batch setting.
Contributions.
We develop a theoretical foundation for online decision-focused learning, enabling its application in non-stationary settings. This presents a significant technical challenge, as the inherent difficulties of decision-focused learning—such as the non-differentiability of the objective function or the lack of convexity of the losses due to the bi-level nature of the problem—compound those already present in online learning. The algorithm that we develop addresses these non-differentiability and non-convexity challenges by leveraging a regularization-based approach combined with perturbation-driven optimism. More precisely, our contributions are as follows:
  1. (i)
    We formalize the online decision-focused learning problem by assuming that at each round, a decision-maker seeks to solve a linear optimization problem over a polytope but does not have access to the true cost function. Thus the decision-maker has to predict the cost using whatever partial information that it has at hand. The cost function is then revealed, and the decision-maker updates its model in a decision-focused fashion. This results in a bi-level optimization problem, where the inner problem consists in making a decision and the outer problem involves optimizing the resulting decision cost.
  2. (ii)
    We present a novel algorithm—Decision-Focused Online Gradient Descent (DF-OGD) for solving this problem. Our approach relies on two main ingredients: (a) regularizing the inner problem described above to make the resulting decision differentiable, and (b) applying an original gradient-based method to the regularized outer problem, leveraging optimistic online learning rakhlin2013online together with perturbation techniques suggala2020online to handle the non-convexity of the problem.
  3. (iii)
    Based on a margin assumption originating from the classification and bandit literature, we provide convergence guarantees in the form of dynamic regret bounds (defined in Section 2) for DF-OGD, when the decision space is either the simplex or a general convex polytope.
  4. (iv)
    Finally, we compare the empirical performance of DF-OGD to that of a classical prediction-focused approach in an online knapsack experiment inspired by mandi2024decision . In our simulation, our method outperforms the standard PFL approach in terms of downstream decision loss, in spite of degraded prediction loss.
Additional related work.
Our work addresses the decision-focused learning problem, pioneered by elmachtoub2022smart . Several approaches have been proposed in the batch setting, among which differentiating the associated KKT conditions gould2016differentiating ; amos2017optnet ; donti2017task ; wilder2018melding ; mandi2020interior , smoothing via random perturbation berthet2020learning , building surrogate losses via duality elmachtoub2022smart or directional gradients huang2024decision and relying on pairwise ranking techniques mandi2022decision , see mandi2024decision ; sadana2025survey for additional references. However, the extension of decision-focused learning to the online setting remains unexplored. While several recent studies address online bi-level optimization shen2023online ; tarzanagh2024online ; lin2023non , none of them are applicable to decision-focused learning as they rely on restrictive smoothness assumptions, that are incompatible with the structure of decision-focused problems. In particular, the decision-focused objective is typically non-convex and features gradients that are either zero or undefined, owing to its underlying linear structure.
This motivates the use of different online methods, specifically tailored to address these problems. On the one hand, lack of differentiability is usually tackled through zero-th order methods heliou2020gradient ; frezat2023gradient , sub-gradient duchi2011adaptive , proximal dixit2019online or smoothing approaches abernethy2014online . On the other hand, methods to address non-convexity in online learning often rely on the existence of a near optimal oracle kalai2005efficient ; agarwal2019learning ; suggala2020online ; xu2024online or additional smoothness conditions lesage2020second ; ghai2022non . All those results are derived for static regret zinkevich2003online ; hazan2023introductiononlineconvexoptimization which compares the learner efficiency to the best static strategy (see Appendix A). An alternative approach to obtain guarantees in the online non-convex setting is to weaken static regret to local regret hazan2017efficient ; aydore2019dynamic ; zhuang2020no ; hallak2021regret , which is the sum of the objective gradient norms evaluated in the iterates over time. Minimizing local regret corresponds to encouraging convergence toward stationary points. However as we shall see, this approach is not sensible in our framework, as objective gradients are either zero or undefined. In this work, we focus on dynamic regret, defined in Section 2. Originally introduced in zinkevich2003online and later developed in hall2013dynamical ; besbes2015non ; zhao2020dyn ; zhao2021improved among others, this regret takes into account the evolution of the environment and is well suited to our framework. We also provide an analysis for the more classic static regret notion based on techniques from suggala2020online , see Appendix B.
Outline.
In Section 2, we introduce the online decision-focused problem, our regret notion and our assumptions. In Section 3, we present the DF-OGD algorithm, before analyzing its dynamic regret performance in the simplex and convex polytope cases in Section 4.1 and Section 4.2 respectively. Finally, we present our experiment in Section 5.
Notation.
For a differentiable map $g:\mathbb{R}^{m}\rightarrow\mathcal{Y}$ such that $\mathcal{Y}\subseteq\mathbb{R}$, $\nabla g(x)$ denotes the gradient of $g$ at $x\in\mathbb{R}^{m}$ and $\nabla^{2}g(x)$ denotes the Hessian of $g$. In the case where $\mathcal{Y}\subseteq\mathbb{R}^{d}$, $\nabla g(x)$ denotes the Jacobian of $g$. For two vectors $(v,w)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $v\succcurlyeq w$ means that $v_{i}\geqslant w_{i}$ for any $i\in[d]$, and $\mathopen{}\mathclose{{}\left\langle v,w}\right\rangle=v^{\mathsf{\scriptscriptstyle T}}w$ refers to the standard Euclidian inner product. Also, $\mathopen{}\mathclose{{}\left\|v}\right\|=\sqrt{\mathopen{}\mathclose{{}\left\langle v,v}\right\rangle}$ is the standard euclidian norm. Given a compact convex set $\Theta\subseteq\mathbb{R}^{d}$, $\Pi_{\Theta}$ denotes the orthogonal projection onto $\Theta$. For a matrix $M\in\mathbb{R}^{m\times d}$, $\mathopen{}\mathclose{{}\left\|M}\right\|_{\mathrm{op}}$ refers to its $\mathrm{L}^{2}$-operator norm. In the case $d=m$, $\lambda_{\min}(M)$ refers to its lowest real eigenvalue and $\lambda_{\max}(M)$ it largest real eigenvalue. For $(x,y)\in\mathbb{R}^{2}$, $x\propto y$ means that there exists $\lambda\in\mathbb{R}^{\star}$ such that $x=\lambda y$.
2 Framework
Sequential decision-making.
We consider an online decision-making problem over $T>0$ periods, defined for any $t\in[T]$ as
$\min_{w\in\mathcal{W}}\ \langle\bar{g}_{t}(X_{t}),w\rangle\;,$ (1)
where $X_{t}\in\mathcal{X}$ are covariates (possibly random), $\mathcal{W}=\mathrm{Conv}(v_{1},\ldots,v_{K})$ is a bounded convex polytope of $\mathbb{R}^{d}$ with non-empty interior and vertices $\{v_{i}\}_{i=1}^{K}$, and $\bar{g}_{t}:\mathcal{X}\to\mathcal{Z}$, $\mathcal{Z}\subset\mathbb{R}^{d}$ is a cost function, which for instance satisfies $\bar{g}_{t}(X_{t})=\mathbb{E}[Z_{t}\mid X_{t}]$ for some hidden state $Z_{t}\in\mathcal{Z}\subset\mathbb{R}^{d}$. Importantly, $X_{t}$ is revealed at the beginning of the round, but not $\bar{g}_{t}(X_{t})$ which is only available at the end of the round. While $\bar{g}_{t}(X_{t})$ is unknown to the decision-maker at the decision time, they have access to a family of models $g:\Theta\times\mathcal{X}\rightarrow\mathcal{Z}$, parameterized by $\Theta\subset\mathbb{R}^{m}$, to predict it. Then, the decision-maker seeks to jointly fit $\theta$ and solve the resulting decision problem. More precisely, the decision-maker considers a joint process $(\theta_{t},w_{t})_{t\in[T]}$, starting from $\theta_{1}\in\Theta$ and a history $\mathcal{H}_{0}=\emptyset$, and defined by the following recursion. At round $t\in[T]$, the decision-maker takes the best action given the collected history $\mathcal{H}_{t-1}$, the current estimate $\theta_{t}$ and a new feature $X_{t}$. Formally, they make the decision $w_{t}=w^{\star}_{t}(\theta_{t})$ where for any $\theta\in\Theta$,
$w_{t}^{\star}(\theta)\in\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\langle g(\theta,X_{t}),w\rangle\;.$ (2)
Then, they observe $\bar{g}_{t}(X_{t})$ and update their history $\mathcal{H}_{t}=\mathcal{H}_{t-1}\cup\{(X_{t},\bar{g}_{t}(X_{t}),\theta_{t},w^{\star}_{t}(\theta_{t}))\}$. Finally, they update their prediction parameter $\theta_{t+1}$ based on $\mathcal{H}_{t}$ through an algorithm $\mathsf{Alg}_{t}$. A central question for the decision-maker is then the choice of algorithms $\{\mathsf{Alg}_{t}\}_{t\in[T]}$ to fit their regression model, that is how to pick $\theta_{t}\in\Theta$ for each $t\in[T]$.
Online decision-focused learning.
From an online perspective, prediction-focused learning consists in selecting at each $t\in[T]$ an algorithm $\mathsf{Alg}_{t}$ to estimate $\operatorname*{\arg\!\min}_{\theta\in\Theta}\mathrm{R}_{t}(\theta)$, where $\mathrm{R}_{t}$ is a risk function based on the historical observations $\mathcal{H}_{t-1}$, typically chosen as an empirical risk of the form $\mathrm{R}_{t}(\theta)=\sum_{i=1}^{t}\ell(\bar{g}_{i}(X_{i}),g(\theta,X_{i}))$, for some loss function $\ell$ (e.g. cross-entropy or squared error).
Alternatively, we consider in this paper, the decision-focused learning approach, which selects parameters $\theta\in\Theta$ by directly minimizing the downstream objective instead of a risk function. Specifically, the decision-maker chooses $\mathsf{Alg}_{t}$ to solve $\operatorname*{\arg\!\min}_{\theta\in\Theta}\langle\bar{g}_{t}(X_{t}),w_{t}^{\star}(\theta)\rangle$ where $w_{t}^{\star}$ is defined in (2). From the previous formulation, we see that the decision-focused learning formulation corresponds to a bilevel optimization problem colson2007overview ; sinha2017review ; ji2021bilevel .
We emphasize that no stationary assumptions are made on the process $\{X_{t},\bar{g}_{t}(X_{t})\}_{t\in[T]}$, allowing an adversarial selection of the features $X_{t}$ and cost functions $\bar{g}_{t}$. To fully accommodate this flexibility, we assess the optimality of $\{\theta_{t}\}_{t\in[T]}$ with a dynamic regret. It compares the learner’s performance against a sequence of best decisions that may vary over time. More precisely, let
$f_{t}:\theta\in\Theta\mapsto\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}\right\rangle\;,$ (3)
denotes the loss incurred when taking an action based on the prediction parameter $\theta_{t}\in\Theta$. In line with zinkevich2003online ; haddouche2024optimisticallytemperedonlinelearning , we define the following dynamic regret:
$\mathfrak{R}_{T}=\sum_{t\in[T]}F_{t}(\theta_{t})-\sum_{t\in[T]}\inf_{\theta\in\Theta}F_{t}(\theta)\;,\quad\text{where}\quad F_{t}:\theta\mapsto\mathbb{E}\mathopen{}\mathclose{{}\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}\right]\;.$ (4)
We also consider a static regret and derive associated bounds in Appendix B. We make the following mild regularity assumptions for the rest of the analysis.
H 1.
(i) $\Theta$ is a compact and convex set with diameter $D_{\Theta}<\infty$. (ii) For any $\theta\in\Theta$ and $t\in[T]$, $\nabla_{\theta}g(\theta,X_{t})$ is continuously differentiable and $\mathopen{}\mathclose{{}\left\|\nabla_{\theta}g(\theta,X_{t})}\right\|_{\mathrm{op}}\leqslant G<\infty$ almost-surely. (iii) For any $t\in[T]$, $\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|\leqslant D_{\mathcal{Z}}<\infty$ almost-surely.
It is common in online learning to assume boundedness of the parameter space, model gradient and prediction space boyd2004convex ; bishop2006pattern . Note that in the well-specified setting, i.e., $\bar{g}_{t}=g(\bar{\theta}_{t},\cdot)$ for some $\bar{\theta}_{t}$, these two former conditions automatically imply the latter. We emphasize however that we do not assume this realizability condition, in contrast to most of the literature on decision-focused learning bennouna2024addressing .
Without further assumptions on the sequence of costs $\bar{g}_{t}$ and models $g$, the problem can still be made arbitrarily hard. Therefore, we make an assumption coming from the classification mammen1999smooth ; tsybakov2004optimal and bandit literature zeevi2009woodroofe ; perchet2013multi about the margins of the cost function. Recall that we denote by $\{v_{i}\}_{i=1}^{K}$ the vertices of $\mathcal{W}$. We define for any $x\in\mathcal{X}$, $u_{i}(\theta,x)=\mathopen{}\mathclose{{}\left\langle g(\theta,x),v_{i}}\right\rangle$ for $i\in[d]$, and assume that
H 2.
There exists $C_{0}>0$ and $\beta\in\mathopen{}\mathclose{{}\left[0,1}\right]$, such that almost surely for any $t\in[T]$, $\theta\in\Theta$ and $\varepsilon\in\mathopen{}\mathclose{{}\left[0,1}\right]$,
$\mathbb{P}\mathopen{}\mathclose{{}\left(\inf_{j\neq I_{t}(\theta)}\{u_{j}(\theta,X_{t})-u_{I_{t}(\theta)}(\theta,X_{t})\}\geqslant\varepsilon\,\mid\mathcal{H}_{t-1}}\right)\geqslant 1-C_{0}\varepsilon^{\beta}\;,$
where $I_{t}(\theta)\in\operatorname*{\arg\!\min}_{i\in[K]}u_{i}(\theta,X_{t})$.
H 2 controls how difficult it is to solve problem (2) since it determines the objective gap between the optimal vertex $v_{I_{t}(\theta)}\in\mathcal{W}$ and the other vertices. In other words, it quantifies how identifiable is the optimal vertex, $\varepsilon=0$ meaning it is not distinguishable from the others. Remark that H 2 is only related to the distribution of the features $\{X_{t}\}_{t\in[T]}$. If it is satisfied for $\beta\geqslant 1$, then it is automatically satisfied for $\beta=1$ and therefore we do not loose any generality restraining $\beta$ to $\mathopen{}\mathclose{{}\left[0,1}\right]$. This assumption is critical in our analysis, as it allows to bound the expected distance between the actual optimal decision $w^{\star}_{t}(\theta)$ and the regularized approximation $\tilde{w}_{t}(\theta)$ introduced in (5) below.
3 An online algorithm for decision-focused learning.
A natural approach to minimize the regret (4) would be to use an Online Gradient Descent (OGD) scheme zinkevich2003online for the algorithms $\{\mathsf{Alg}_{t}\}_{t\in[T]}$, which is standard in online learning. However, since in (2), $w^{\star}_{t}(\theta)$ minimizes a linear function over the convex polytope $\mathcal{W}$, $\theta\mapsto w_{t}^{\star}(\theta)$ ranges in the finite set of vertices $\{v_{1},\ldots,v_{K}\}$, thus $\nabla w^{\star}_{t}(\theta)$ is either zero or undefined.
To address this issue, following wilder2018melding , we propose to add a regularizer $\mathcal{R}$ to the objective function in (2). Accordingly, we define for any $\theta\in\Theta$,
$\tilde{w}_{t}(\theta)\in\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\{\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),w}\right\rangle+\alpha_{t}\mathcal{R}(w)\}\;,$ (5)
which is a regularized approximation of $w^{\star}_{t}$. When this surrogate is continuously differentiable—which holds for our choices of $\mathcal{R}$ in Section 4.1 and Section 4.2—we can use $\nabla\tilde{w}_{t}(\theta)$ in our algorithmic routine in place of $\nabla w^{\star}_{t}(\theta)$. This approach amounts to minimizing the surrogate
$\tilde{f}_{t}:\theta\mapsto\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle\;,$ (6)
instead of $f_{t}$, which admits gradients of the form
$\nabla\tilde{f}_{t}(\theta)=\nabla\tilde{w}_{t}(\theta)^{\mathsf{\scriptscriptstyle T}}\bar{g}_{t}(X_{t})\;.$
Note that $\tilde{f}_{t}$ is not necessarily convex (even when $\theta\mapsto g(\theta,X)$ is linear) because it involves the parametrized minimizer $\tilde{w}_{t}(\theta)$. This prevents us from directly using known online convex optimization algorithms hazan2023introductiononlineconvexoptimization . However, a recent line of research kalai2005efficient ; agarwal2019learning ; suggala2020online ; xu2024online has developed online algorithms in the non-convex case, provided the losses are Lipschitz-continuous—which is the case with our choices of regularizers in Section 4.1 and Section 4.2 (see Lemmas 2 and 7). These studies combine near-optimal oracles with perturbation techniques to establish sublinear bounds on the expected regret. In line with this literature, we also assume to have access to a near-optimal oracle. In the following assumption, we define for any $t\in[T]$, $\tilde{F}_{t}:\theta\mapsto\mathbb{E}[\tilde{f}_{t}(\theta)\,|\,\mathcal{H}_{t-1}]$.
H 3.
There exists $\xi>0$ such that for any $t\in[T]$, upon observing $\mathcal{H}_{t}$, we can compute $\vartheta_{t}\in\Theta$ satisfying almost surely:
$\tilde{F}_{t}(\vartheta_{t})\leqslant\inf_{\theta\in\Theta}\tilde{F}_{t}(\theta)+\xi\;.$ (7)
One way to design such oracles is to apply Stochastic Gradient Descent amari1993backpropagation to the differentiable function $\tilde{f}_{t}$. In this case, $\xi\geqslant 0$ captures both the error made by the offline minimization method generating $\vartheta_{t}$, and the gap between the local optimum to which it converges and the global optimum. In particular, $\xi$ vanishes for benign loss landscapes. Finally, note that in the case where $\tilde{w}_{t}(\theta)$ can only be approximated, then $\xi$ can also encompass the resulting error.
Leveraging the regularization and perturbation approaches previously mentioned, we now present our method, referred to as DF-OGD, in Algorithm 1.
Algorithm 1 Decision-Focused Online Gradient Descent (DF-OGD)
1:  Input: horizon $T>0$, initialization $\theta_{1}\in\Theta$ and history $\mathcal{H}_{0}=\emptyset$.
2:  for each $t\in\{1,\ldots,T\}$ do
3:     Observe $X_{t}\in\mathcal{X}$ and play $w^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\mathopen{}\mathclose{{}\left\langle g(\theta_{t},X_{t}),w}\right\rangle\;.$
4:     Observe $\bar{g}_{t}(X_{t})\in\mathcal{Z}$ and update the history $\mathcal{H}_{t}$.
5:     Observe $\vartheta_{t}\in\Theta$ satisfying (7).
6:     Draw $\delta_{t}\sim\text{Unif}([0,1])$, compute $u_{t}=\vartheta_{t}+\delta_{t}(\theta_{t}-\vartheta_{t})$ and ~$\tilde{\nabla}_{t}(u_{t})=\nabla\tilde{w}_{t}(u_{t})^{\mathsf{\scriptscriptstyle T}}\bar{g}_{t}(X_{t})\;.$
7:     Update $\theta_{t+1}=\Pi_{\Theta}(\theta_{t}-\eta_{t}\tilde{\nabla}_{t}(u_{t}))\;.$
8:  end for
At each time $t\in[T]$, Algorithm 1 first makes a decision $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ based on the current parameters $\theta_{t}\in\Theta$. Then, it observes $\bar{g}_{t}(X_{t})\in\mathcal{Z}$, determines the surrogate objective $\tilde{f}_{t}$ and computes the oracle $\vartheta_{t}$ from H 3. Next, it determines the gradient $\nabla\tilde{w}_{t}(u_{t})$, evaluated in a point $u_{t}$ drawn uniformly at random in $[\theta_{t},\vartheta_{t}]$. From there, it updates $\theta_{t+1}$ through the classical gradient step scheme of OGD, with the noticeable difference that the gradient is evaluated in $u_{t}$ instead of $\theta_{t}$.
Algorithm 1 requires determining $w^{\star}$, which can be efficiently achieved by classic numerical solvers in reasonable settings. However, we show in Appendix E that only being able to determine $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ such that $\mathopen{}\mathclose{{}\left\langle g(\theta_{t},X_{t}),\underline{w}_{t}(\theta_{t})}\right\rangle-\mathopen{}\mathclose{{}\left\langle g(\theta_{t},X_{t}),w^{\star}_{t}(\theta_{t})}\right\rangle\leqslant\kappa$ for some $\kappa>0$ only shifts the regret bounds of the next section by $\kappa$.
Note that while DF-OGD relies on OGD, it is possible to use other algorithmic approaches such as Follow the Perturbed Leader (FTPL) suggala2020online , see DF-FTPL in Appendix B for which we provide a static regret analysis.
4 A theoretical study of convergence
We now provide convergence guarantees for Algorithm 1 in the form of expected dynamic regret bounds in two settings. First, we deal with the case where $\mathcal{W}$ is the simplex. Second, we consider the general setting where $\mathcal{W}$ is any bounded convex polytope with non-empty interior.
4.1 Online decision-focused learning on the simplex
We first focus on the case where $\mathcal{W}$ is the simplex of $\mathbb{R}^{d}$, that is
[…]
This setting encompasses various decision-making scenarios such as portfolio selection li2014online or mixed strategy design in multiplayer games syrgkanis2015fast . In this case, we choose the negative entropy
$\mathcal{R}_{0}:w\mapsto\sum_{i\in[d]}w_{i}\ln(w_{i})\;,$
as the regularizer in (5). The minimizer $\tilde{w}_{t}(\theta)\in\mathcal{W}_{0}$ in (5) with $\mathcal{R}=\mathcal{R}_{0}$ can easily be shown to be the softmax mapping, that is it satisfies for any $i\in[d]$:
$\tilde{w}_{t,i}(\theta)=\frac{\exp(-\alpha_{t}^{-1}g_{i}(\theta,X_{t}))}{\sum_{k\in[d]}\exp(-\alpha_{t}^{-1}g_{k}(\theta,X_{t}))}\;.$ (8)
It is clear from this expression that $\tilde{w}_{t}$ is differentiable, and that for any $\theta\in\Theta$:
$\nabla\tilde{w}_{t}(\theta)=-\frac{1}{\alpha_{t}}\mathopen{}\mathclose{{}\left[\mathrm{diag}[\tilde{w}_{t}(\theta)]-\tilde{w}_{t}(\theta)\tilde{w}_{t}(\theta)^{\mathsf{\scriptscriptstyle T}}}\right]\nabla_{\theta}g(\theta,X_{t})\;.$ (9)
The regularized minimizer (8) and its Jacobian (9) can readily be used in Algorithm 1. In this case, the regret (4) incurred by Algorithm 1 can be bounded as follows.
Proposition 1.
Assume H 1, H 2, H 3, $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and that Algorithm 1 is instantiated with a non-increasing sequence $\mathopen{}\mathclose{{}\left(\eta_{t}}\right)_{t\in[T]}$. Then, denoting $P_{T}=\sum_{t=2}^{T}\mathopen{}\mathclose{{}\left\|\vartheta_{t}-\vartheta_{t-1}}\right\|$,
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}G\eta_{t}}{32T\alpha_{t}^{2}}$
$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2\ln(d)C_{0}\{\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d)\}]+\xi\;.$
The term $P_{T}$–which ultimately depends on the sequence $(X_{1},\bar{g}_{1}(X_{1})),\ldots,(X_{t},\bar{g}_{T}(X_{T}))$– controls the variation of the problem through time. Proposition 1 makes clear that keeping the dynamic regret (4) low becomes difficult when the revealed cost vectors vary drastically over time. Our guarantee does not depend on the dimension $m$ of $\Theta$, a notable strength stemming from the optimistic approach. This is sharp contrast with non-optimistic algorithms, such as DF-FTPL in Appendix B, which suffers a $m^{2}$ factor in its associated static regret bound (Proposition 4).
In addition to the usual trade-off associated with the learning rate $\eta_{t}$, Proposition 1 highlights a new trade-off introduced by the regularization strength $\alpha_{t}$. As $\alpha_{t}$ increases, the second term decreases as the mapping $\tilde{f}_{t}$ becomes smoother in the sense that its Lipschitz constant decreases (see Lemma 2). However, the third term grows larger as $\tilde{w}_{t}(\theta)$ becomes a poorer approximation of $w^{\star}_{t}(\theta)$. It is therefore natural to wonder how to pick the sequences $(\eta_{t})_{t\in[T]}$ and $(\alpha_{t})_{t\in[T]}$ to address these trade-offs.
In the following corollary, we define for any $t\in[T]$,
$P_{t}=\begin{cases}\sum_{s=1}^{t}\mathopen{}\mathclose{{}\left\|\vartheta_{s}-\vartheta_{s-1}}\right\|&\text{if }t\geqslant 2\;,\\ 0&\text{otherwise}\end{cases}\;.$
We prove that under a mild assumption regarding the growth of $(P_{t})_{t\in[T]}$, we can pick $(\alpha_{t})_{t\in[T]}$ and $(\eta_{t})_{t\in[T]}$ in a fully adaptative way to obtain an explicit regret bound.
Corollary 1.
Assume H 1,H 2, H 3, $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and $(t^{-1}(1+P_{t}))_{t\in[T]}$ is non-increasing almost surely. Then, using Algorithm 1 with $\alpha_{t}\propto t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto t^{-3/4}(1+P_{t})^{3/4}$ for any $t\in[T]$ gives:
$\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]}{T}=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\ln(d)\mathopen{}\mathclose{{}\left(\frac{1+\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{T}}\right)^{1/4}+\xi}\right)\;,$
where $\tilde{\mathcal{O}}$ contains polynomial dependencies in $\ln(1/\alpha),\ln(\ln(d))$.
Note that the resulting expected regret bound only depends logarithmically on the dimension of the decision space.
We stress that the assumption about $(t^{-1}(1+P_{t}))_{t\in[T]}$ being non-increasing can be weakened to the existence of a non-increasing sequence $(r_{t})_{t\in[T]}$ such that $t^{-1}P_{t}\leqslant r_{t}$ for any $t\in[T]$ without affecting the result. This assumption is merely technical, and we believe that Corollary 1 would hold without it up to slight modifications. Finally, under the setting of Corollary 1, we remark that $\eta_{t}$ and $\alpha_{t}$ are computable on the fly as they only require the knowledge of $P_{t}$, which depends itself on observed variables $(\vartheta_{s})_{s\in[T]}$. This differs, for instance, from the dynamic regret bound of OGD in (hazan2023introductiononlineconvexoptimization, , Chapter 10) which requires the knowledge of $P_{T}$ before training to reach their optimal rate.
We conclude by mentioning that in Appendix D, we consider a stronger version of H 2 which allows us to derive a counterpart of Proposition 1 and Corollary 1 for a stronger notion of dynamic regret.
4.2 Generic polytope
In Section 4.1, the very specific structure of the simplex $\mathcal{W}_{0}$ hinted us to choose $\mathcal{R}$ to be the negative entropy. However many problems, such as shortest-path, min-cost-flow lin2014quickest or unit commitment problems saravanan2013solution , require optimizing on a generic polytope. In what follows, we go back to the generic case where $\mathcal{W}=\mathcal{W}_{1}$ is a bounded, convex polytope with non-empty interior. We can write $\mathcal{W}_{1}$ as the intersection of $n$ half-spaces, denoting by $n>0$ its number of faces, that is
$\mathcal{W}_{1}=\{w\in\mathbb{R}^{d}:\>A^{\mathsf{\scriptscriptstyle T}}w-b\preccurlyeq 0\}\;,$
where $A\in\mathbb{R}^{d\times n},b\in\mathbb{R}^{n}$. In what follows, we assume that $\mathcal{W}_{1}$ is not degenerated, i.e. $AA^{\mathsf{\scriptscriptstyle T}}\in\mathbb{R}^{d\times d}$ is full-rank.
To tackle this more general structure, our strategy is to use a regularizer $\mathcal{R}$ such that $\tilde{w}_{t}$ always lies in the strict interior of $\mathcal{W}_{1}$, in which case the Jacobian map $\theta\mapsto\nabla\tilde{w}_{t}(\theta)$ can be determined by the first order condition and the implicit function theorem. To this end, we propose to use log-barriers as the regularizer in (5), that is
$\mathcal{R}_{1}:w\mapsto-\sum_{i=1}^{n}\ln(b_{i}-A_{i}^{\mathsf{\scriptscriptstyle T}}w)\;,$ (10)
where $A_{i}\in\mathbb{R}^{d}$ is the $i$-th column of $A$. With this choice of regularization, we show in Lemma 6 of Section F.2 that:
$\nabla\tilde{w}_{t}(\theta)=-\alpha_{t}^{-1}\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}(A_{i}^{\mathsf{\scriptscriptstyle T}}\tilde{w}_{t}(\theta)-b_{i})^{-2}A_{i}A_{i}^{\mathsf{\scriptscriptstyle T}}}\right)^{-1}\nabla_{\theta}g(\theta,X_{t})\;,$ (11)
which can be readily used in Algorithm 1. Note that (11) involves inverting a sum of rank-one matrices—a common operation in interior point methods—which can be performed efficiently; see jiang2020faster . With this choice, we can establish the following dynamic regret bound.
Proposition 2.
Assume H 1, H 2, H 3, $\mathcal{W}=\mathcal{W}_{1}$, $AA^{\mathsf{\scriptscriptstyle T}}$ is full-rank, $\mathcal{R}=\mathcal{R}_{1}$ and Algorithm 1 is instantiated with a non-increasing sequence $\mathopen{}\mathclose{{}\left(\eta_{t}}\right)_{t\in[T]}$. Then,
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{}\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}^{2}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}\right)^{2}$
$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2n\max_{k\in[K]}\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{1}C_{0}(\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d))]+\xi\;,$
where $P_{T}$ is defined as in Proposition 1.
Proposition 2 closely resembles to Proposition 1, but with an additional linear dependence on the number of faces $n>0$. This can be seen as the cost of operating over a more complex admissible space than the simplex. As before, it is possible to choose sequences $(\alpha_{t})_{t\in[T]}$ and $(\eta_{t})_{t\in[T]}$ to balance the trade-off discussed in Section 4.1 in a fully adaptative way.
Corollary 2.
Assume H 1, H 2, H 3, $\mathcal{W}=\mathcal{W}_{1}$, $\mathcal{R}=\mathcal{R}_{1}$ and $(t^{-1}(1+P_{t}))_{t\in[T]}$ is non-increasing almost surely. Then, using Algorithm 1 with $\alpha_{t}\propto t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto t^{-3/4}(1+P_{t})^{3/4}$ for any $t\in[T]$ leads to :
$\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]}{T}=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(n\frac{1+\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{T}^{1/4}+\xi}\right)\;.$
Note that since we assume $AA^{\mathsf{\scriptscriptstyle T}}$ to be full rank, we have in particular that $n\geqslant d$, which is a deteriorated dependence compared to the $\ln(d)$ factor of Corollary 1.
5 Experiments
00111122223333444455550011112222333344445555Value of Item 1Value of Item 2
Figure 1: Figure 2 from mandi2024decision .
Setting.
[…]
After having made their decision, the decision-maker observe the true item costs $\bar{g}_{t}(X_{t})=\mathbb{E}[v_{t}\,|\,X_{t}]$, and update $\theta_{t+1}$ for the next round based on this feedback. Note that this setting can directly be mapped in the simplex example of Section 4.2, since $a^{\star}_{t}(\theta_{t})=w^{\star}_{t}(\theta_{t})$ with $w^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{w\in\mathcal{W}_{0}}\mathopen{}\mathclose{{}\left\langle w,g(\theta_{t},X_{t})}\right\rangle$, $\mathcal{W}_{0}$ being the simplex of $\mathbb{R}^{K}$.
[…]
Synthetic data.
We instantiate the previous knapsack problem with the following synthetic data. For any $t\in[T]$, we draw $X_{t}\in\mathbb{R}^{K\times p}$ with correlated rows, and generate $v_{t}\in[0,1]^{k}$ as follows:
$v_{t}=\alpha\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon_{t}\;,$ (12)
where $\alpha>0$, $\varepsilon_{t}\sim\mathrm{N}(0,I_{K})$ is a Gaussian noise and $\theta^{\star}_{t}\in\mathbb{R}^{p}$ is a parameter which follows an autoregressive recursion:
$\theta^{\star}_{t}=\frac{1}{2}\theta^{\star}_{t-1}+\frac{1}{2}\zeta_{t}\,,\quad\text{where}\quad\zeta_{t}\sim\mathrm{N}(0,I_{p})\;.$
This is a challenging data generating process, since $\bar{g}_{t}(X_{t})=\sin^{4}((2X_{t}^{\mathsf{\scriptscriptstyle T}}\theta^{\star}_{t})^{-1})$ is non-stationary and highly non-linear, and features are correlated. Equation 12 is discussed more in detail in Appendix C. To predict $\hat{v}_{t}$ from $X_{t}$, we assume that the decision-maker has access to a class of linear predictors of the form $g:(\theta,X)\mapsto X\theta$.
Comparison between DFL and PFL.
In this framework, we compare DF-OGD  (Algorithm 1) to a Prediction-Focused Online Gradient Descent (PF-OGD) described in Algorithm 2, which minimizes the loss $\ell^{\mathrm{mse}}:(\theta,v,X)\mapsto\mathopen{}\mathclose{{}\left\|v-X\theta}\right\|^{2}$.
Algorithm 2 Prediction-Focused Online Gradient Descent (PF-OGD).
1:  Input: horizon $T>0$, initialization $\theta_{1}\in\Theta$.
2:  for each $t\in\{1,\ldots,T\}$ do
3:     Observe $X_{t}$, predict $\hat{v}_{t}=g(\theta_{t},X_{t})$
4:     Play $a^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{i\in[K]}\hat{v}_{t}(i)\;,$ observe $\bar{g}_{t}(X_{t})$.
5:     Update $\theta_{t+1}=\Pi_{\Theta}(\theta_{t}-\eta\nabla\ell^{\mathrm{mse}}(\theta_{t},\bar{g}_{t}(X_{t}),X_{t}))\;.$
6:  end for
In Figure 2, we plot on the left-hand panel the average cumulated cost $t\mapsto t^{-1}\sum_{i=1}^{t}v_{t}(a^{\star}_{t}(\theta_{t}))$ incurred by DF-OGD and PF-OGD over 70 runs of $T=2000$ timesteps, as well as the associated 95% confidence intervals. On the right-hand panel, we plot the average Mean Square Error resulting from the sequence of prediction parameters $(\theta_{t})_{t\in[T]}$. Interestingly, DF-OGD produces parameters that induce better decisions as indicated by the lower average cumulated loss, but degraded predictions as suggested by the constantly higher MSE. This is in line with the intuition according to which the decision-focused approach sacrifices accuracy for the sake of better downstream decisions.
Figure 2: Average cumulated cost (left) and prediction errors (right) of DF-OGD and PF-OGD.
We conclude this section by mentioning that we perform additional numerical experiments to study how the relative performances of DFL and PFL are affected by model mispecification in Appendix C. This experiment empirically supports the idea that DFL outperforms PFL when models are clearly mispecified.
6 Conclusion.
Decision-focused learning is a promising approach for better leveraging prediction in real-life decision making processes. While the theoretical analysis of decision-focused learning is currently restricted to the batch setting, we extend it to the online framework. This allows us to consider non-stationary data processes and varying objective functions. We propose an algorithm based on regularization, perturbation and optimism which only requires picking a regularizer along with a sequence of regularization coefficients and learning rates. In particular, we show how to pick these hyper-parameters in two canonical setups, namely when the decision space is the simplex, and when it is a general bounded, convex polytope with non-empty interior. In both cases, we provide theoretical guarantees for our algorithm, in the form of upper bounds on the expected average dynamic regret. Finally, we compare the empirical performances of our approach to a more standard prediction-focused learning procedure in a knapsack experiment with a mispecified prediction model. In this experiment, our method outperforms the PFL one, which supports the idea that DFL leads to decision that are more robust to model mispecification.
We believe that this work can be extended in several ways. First, it would be interesting to study whether meaningful guarantees could be obtained under weaker assumptions than ours. Second, we believe that alternative smoothing techniques could be successfully employed, such as Moreau-Yosida transform and proximal operator, or random perturbation and Fenchel conjugate techniques berthet2020learning . Analyzing the theoretical performance of these methods would be of significant interest. Third, it would be valuable to investigate less adversarial environments, such as those involving i.i.d. data, to explore whether stronger theoretical guarantees could be obtained and whether novel algorithmic designs might emerge. Finally, implementing our method in more ambitious experimental settings would be of great interest from a practical point of view.
Appendix A A brief recap on online learning
Online learning (OL) is a paradigm in which data is processed sequentially, either because the practitionner does not collect all data prior to analysis or because the dataset dynamically evolves through time, or simply because handling batch of data is numerically too demanding. From the seminal work of [77], which proposed an online version of the celebrated gradient descent algorithm, OL has been intensively studied (see [36] for an overview). The classical performance criterion of an online learning algorithm is the static regret. Given a sequence of loss functions $(\ell_{t}:\mathcal{K}\rightarrow\mathbb{R})_{t\geqslant 1}$, the static regret compares the efficiency of a sequence of decisions $\hat{\mu}=(\hat{\mu}_{t})_{t\geqslant 1}$ to the best fixed strategy for $T>0$:
$\operatorname{Regret}_{T}^{\mathrm{S}}(\hat{\mu})=\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\inf_{\mu_{0}\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mu_{0}),$ (13)
Classical upper bounds on static regret involve a sub-linear rate. For instance, [77] proposed a $\mathcal{O}(\sqrt{T})$ bound for Online Gradient Descent (OGD) which is valid for convex losses.
Static regret may not be sufficient to assert the efficiency of an online algorithm as the class of static strategies is limited compared to all possible strategies. Hence the notion of dynamic regret introduced by [77] and further developed by [34]. For any sequence $\hat{\mu}$ of predictors and any sequence $\mu$ of dynamic strategies, the dynamic regret is given by
$\operatorname{Regret}_{T}(\hat{\mu},\mu)=\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\sum_{t=1}^{T}\ell_{t}(\mu_{t})\,,\qquad T\geq 1\,.$
Dynamic regret has attracted a lot of interest recently, especially when the comparator sequence is set to $(\inf_{\mu\in\mathcal{K}}\ell_{t}(\mu))_{t\geqslant 1}$ (worst-case dynamic regret, as in [13, 74]) but also for any comparator sequence (universal dynamic regret, as in [75]). Note that dynamic regret is a stronger notion that static regret, since it compares the performance of $\hat{\mu}$ to the best dynamic strategy ($\inf_{\mu\in\mathcal{K}}\ell_{t}(\mu)$), while static regret relies on the best static strategy $(\inf_{\mu\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mu))$.
Appendix B An online decision-focused algorithm with static regret guarantees
In our work, we mainly focused on upper bounds of the dynamic regret, given its versatility and its ability to accommodate non-stationary environments. However, static regret as defined in (13) is a more common measure of performance. In this subsection, we show that it is possible to extend our analysis to static regret based on techniques from [64]. More precisely, we propose in what follows a decision-focused version of the Follow-The-Perturbed-Leader (FTPL) algorithm for nonconvex losses. We first introduce some necessary background in Section B.1, and apply it to Decision-Focused learning in Section B.2.
B.1 Background on FTPL.
Approximate Optimization Oracle.
The recent work of [64] proposed an online algorithm for nonconvex losses $(\ell_{t})_{t\in[T]}$ with static regret guarantees. They rely on an approximate optimization oracle $\mathcal{O}$ which takes as input any function $\ell$, a $d$-dimensional vector $\sigma$ and returns an approximate minimizer of $\ell-(\sigma,\cdot)$. An optimization oracle is called "( $\xi,\chi$ )-approximate optimization oracle" if it returns $\mu^{*}\in\mathcal{K}$ such that
$\ell\mathopen{}\mathclose{{}\left(\mu^{*}}\right)-\mathopen{}\mathclose{{}\left\langle\sigma,\mu^{*}}\right\rangle\leqslant\inf_{\mu\in\mathcal{K}}[\ell(\mu)-\langle\sigma,\mu\rangle]+\mathopen{}\mathclose{{}\left(\xi+\chi\|\sigma\|_{1}}\right)\;.$ (14)
We denote the output of such an optimization oracle by $\mu\star=\mathcal{O}_{\xi,\chi}(\ell-\langle\sigma,\cdot\rangle)$. Note the property (14) is very close from the assumption H 3 we made in Section 3 on the expert sequence $(\vartheta_{t})_{t\in[T]}$ (in this case $\chi=0$ as no $\sigma$ is involved).
Follow The Perturbed Leader.
Given access to an ($\xi,\chi$ )-approximate optimization oracle, [64] study the FTPL algorithm which is described by the following recursion. Starting from $\hat{\mu}_{1}$, at each time steps $t$, $\hat{\mu}_{t}\in\mathcal{K}$ is updated as follows:
$\hat{\mu}_{t+1}=\mathcal{O}_{\xi,\chi}\mathopen{}\mathclose{{}\left(\sum_{i=1}^{t-1}\ell_{i}-\mathopen{}\mathclose{{}\left\langle\sigma_{t},\cdot}\right\rangle}\right)$ (15)
where $\sigma_{t}\in\mathbb{R}^{d}$ is a random perturbation such that $\sigma_{t,j}$, the j-th ccoordinate of $\sigma_{t}$, is sampled from $\operatorname{Exp}(\eta)$, the exponential distribution with parameter $\eta$.
Following this update route, the addition of an exponential noise allowing to handle the non-convexity of the losses, they reach the following static regret bound.
Proposition 3 (Theorem 1 of [64]).
Let $D$ be the diameter of $\mathcal{K}$. Suppose that $\ell_{t}$ is $L$-Lipschitz w.r.t L1 norm, for all $t\in[T]$. For any fixed $\eta$, FTPL (Equation 15) with access to a $(\xi,\chi)$-approximate" optimization oracle satisfies the following static regret bound:
$\frac{1}{T}\operatorname{Regret}_{T}^{\mathrm{S}}(\hat{\mu})\leqslant\mathcal{O}\mathopen{}\mathclose{{}\left(\eta m^{2}DL^{2}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi mL}\right),$
where $\operatorname{Regret}_{T}^{\mathrm{S}}(\hat{\mu})$ is defined as in (13), and $\mathbb{E}$ denotes the expectation over $\sigma_{1}\cdots\sigma_{T}$.
B.2 Application to Decision-Focused Learning.
We now show how to adapt the techniques and results from Section B.1 to our online decision-focused problem. First, we instantiate the static regret defined (13) in our framework. For any $T>0$, loss sequence $(f_{t})_{t\in[T]}$ as defined in (3), and algorithms $(\mathsf{Alg}_{t})_{t\in[T]}$ outputting $(\theta_{t})_{t\in[T]}$, we define for any $\theta\in\Theta$:
$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)=\sum_{t\in[T]}F_{t}(\theta_{t})-\sum_{t\in[T]}F_{t}(\theta)\;.$ (16)
where $F_{T}:\theta\rightarrow\mathbb{E}\mathopen{}\mathclose{{}\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}\right]$ is the conditional expectation of the loss $f_{t}$ with respect to the history $\mathcal{H}_{t-1}$.
Note that Proposition 3 holds for any sequence of lipschitz losses, and that the surrogate losses ($\tilde{f}_{t}$) are indeed Lipschitz for our choice of regularizers in Section 4.1 and Section 4.2 (see Lemmas 2 and 7). from this observation, we derive a non-optimistic online algorithm following the same pattern as DF-OGD (Algorithm 1), but which leverages FTPL to enjoy the guarantee proposition 3.
Algorithm 3 Decision-Focused Follow The Perturbed Leader DF-FTPL
1:  Input: horizon $T>0$, initialization $\theta_{1}\in\Theta$.
2:  for each $t\in\{1,\ldots,T\}$ do
3:     Observe $X_{t}\in\mathcal{X}$ and play $w^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\mathopen{}\mathclose{{}\left\langle g(\theta_{t},X_{t}),w}\right\rangle\;.$
4:     Observe $\bar{g}_{t}(X_{t})\in\mathcal{Z}$ and update the history $\mathcal{H}_{t}$.
5:     Draw $\sigma_{t}\in\mathbb{R}^{d}$ such that for all $j\in[d]$, $\sigma_{j,t}\sim\mathrm{Exp}(\eta)$.
6:     Update
$\theta_{t+1}=\mathcal{O}_{\xi,\chi}\mathopen{}\mathclose{{}\left(\sum_{i=1}^{t}\tilde{f}_{i}-\mathopen{}\mathclose{{}\left\langle\sigma_{t},\cdot}\right\rangle}\right)$
7:  end for
In Algorithm 3, recall that $\tilde{f}_{i}$ are the surrogate losses defined in Equation 5. At each time $t\in[T]$, Algorithm 3 first makes a decision $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ based on the current parameters $\theta_{t}\in\Theta$. Then, it observes $\bar{g}_{t}(X_{t})\in\mathcal{Z}$, determines the surrogate objective $\tilde{f}_{t}$. Next, it updates $\theta_{t+1}$ following the FTPL algorithm applied to the surrogate losses $(\tilde{f}_{t})_{t\in[T]}$.
In order to obtain static regret bounds, we follow a similar route than in Section 4.1: we work on the simplex $\mathcal{W}_{0}$ and use the negative entropy $\mathcal{R}_{0}$ as regularizer to define $(\tilde{f}_{t})_{t\in[T]}$. We refer to Section 4.1 for a precise definition of these notions. We then have the following result.
Proposition 4.
Assume H 1, H 2 and having access to an ($\xi,\chi$ )-approximate optimization oracle $\mathcal{O}$. Fix for any $t$, $\alpha_{t}=\alpha$ in (6), $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and that DF-FTPL (Algorithm 3) is instantiated with $\eta>0$. Then, for any $\theta\in\Theta$:
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{S}}(\theta)}\right]$ $\leqslant\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi m\frac{1}{\alpha}+\alpha\ln(d)}\right)\;,$
where $\mathbb{E}$ denotes the expectation on both data and the intrinsic randomness of DF-FTPL and $\tilde{\mathcal{O}}$ contains polynomial dependency in $\ln(1/\alpha),\ln(\ln(d))$.
Thus, it is possible to have a decision-focused algorithm with static regret guarantees. A notable strength of Algorithm 3 compared to Algorithm 1 is that it does not require any optimism to run. However, this comes at a major cost: a dependency in the dimension $m^{2}$ of $\Theta$ which is not involved in Proposition 1 thanks to the optimistic approach.
The proof of Proposition 4 is deferred to Section F.3.
Appendix C Additional material on our numerical experiment.
In this section, we provide more detail about the experimental setup in Section 5 and perform additional numerical simulations.
Details on the experimental setup.
We start by explaining more precisely how the data used in our experiment are drawn. First, for any $t\in[T]$, $X_{t}\in\mathbb{R}^{K\times p}$ is constructed as follows: (i) we generate a Toeplitz covariance matrix $\Sigma=(\rho^{\mid i-j\mid})_{(i,j)\in[K]^{2}}$ for some $\rho\in(0,1)$, (ii) apply a Cholesky decomposition to it: $\Sigma=LL^{\mathsf{\scriptscriptstyle T}}$, and (iii) generate a matrix $\bar{X}_{t}$ with standard Gaussian entries. Then, $X_{t}$ is defined as $X_{t}=L\bar{X}_{t}$. This introduces correlation between features, which makes convergence harder for an online linear model. Second, we generate $v_{t}\in\mathbb{R}^{K}$ as follows:
$v_{t}=\min\mathopen{}\mathclose{{}\left(\max\mathopen{}\mathclose{{}\left(\tilde{v}_{t},0}\right),1}\right)\quad\text{with}\quad\tilde{v}_{t}=\alpha\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon\;.$
In the above equation, $\alpha>0$ is a constant, $\theta_{t}^{\star}$ follows the following autoregressive recursion:
$\theta^{\star}_{t}=\frac{1}{2}\theta^{\star}_{t-1}+\frac{1}{2}\zeta_{t}\,,\quad\text{where}\quad\zeta_{t}\sim\mathrm{N}(0,I_{K})\;,$
and $\varepsilon_{t}\sim\mathrm{N}(0,I_{K})$ is Gaussian noise. The fact that $\theta^{\star}_{t}$ varies throughout time and that the relationship between $v_{t}$ and $X_{t}$ is highly linear makes learning hard for a linear model.
We now present the values used for the different parameters in our experiment. We consider $K=5$ items and $p=10$ features. The horizon is set to $T=2000$. For each plot, we run $N=70$ times DF-OGD and PF-OGD. The reported error bars are 95% Gaussian confidence intervals (from the CLT). In (12), $\alpha=45$ and the correlation coefficient for $X_{t}$ is $\rho=0.8$. Our algorithms are instantiated with the following parameters (optimized through a grid search procedure). First, DF-OGD runs with $(\alpha_{t})$ and $(\eta_{t})$ as in Corollary 1. The oracle from H 2 is obtained through a SGD algorithm performing $t_{\vartheta}=10$ steps at each iteration with a learning rate set to $\eta_{\vartheta}=0.01$, while being initialized at $\vartheta_{t-1}$. On the other hand, PF-OGD runs with a learning rate $\eta=0.01$. Our code mostly builds on pytorch and numpy.
Deviation from realizability.
As argued in Section 1, prediction-focused learning would be optimal if the prediction model made no mistake, since in this case problem (2) would perfectly align with (1). As a consequence, one would expect prediction-focused learning to perform very well in the realizable setting, that is when for any $t\in[T]$, there exists $\bar{\theta}_{t}\in\Theta$ such that $\bar{g}_{t}(X_{t})=g(\bar{\theta}_{t},X_{t})$. On the flip side, it is likely to struggle in a mispecified setting. This is in contrast with DF-OGD, whose theoretical performance analysis does not highlight any peculiar dependence on realizability. We therefore hypothetize that PFL should outperform DFL in the realizable case, and DFL should take the upper hand when the prediction model is highly mispecified. To test this hypothesis, we run an alternative experiment where we interpolate between the well-specified, linear case, and ill-specified, non-linear one. More precisely, for $\gamma\in[0,1]$, we generate $(X_{t},v^{(\gamma)}_{t})_{t\in[T]}$ as follows:
$v^{(\gamma)}_{t}=(1-\gamma)X_{t}\theta^{\star}_{t}+\gamma\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon_{t}\;.$
In Figure 3, we make $\gamma$ vary from 0 to 1. At each value, we plot the average cumulated cost gap between DF-OGD and PF-OGD, averaged over $50$ runs with horizon $T=1000$. This experiment provides empirical support to the previous conjecture, namely, DFL becomes a competitive option when the prediction model is ill-specified.
Figure 3: Average cumulated cost gap PFL - DFL as a function of $\gamma$.
Appendix D A stronger version of H 2 for a stronger notion of regret
In this section, we present a stronger version of H 2, displayed in H 4. We show that under this new assumption, our guarantees still hold for a stronger notion of regret.
Here, we focus on the simplex setting of Section 4.1: we take $\mathcal{W}=\mathcal{W}_{0}$ the simplex and $\mathcal{R}=\mathcal{R}_{0}$ the negative entropy. This section show that at the cost of slightly strengthening this assumption, we may obtain regret bounds for DF-OGD with a stronger version of regret, which we define below.
A novel dynamic regret.
We define the pseudo dynamic regret as, for any $t$ any loss sequence $(f_{t})_{t\in[T]}$,
$\mathfrak{R}_{T}^{\mathrm{d}}:=\sum_{t\in[T]}f_{t}(\theta_{t})-\sum_{t\in[T]}\inf_{\theta\in\Theta}f_{t}(\theta).$
This notion is exactly what appeared in [77]. Note that since $\mathbb{E}\mathopen{}\mathclose{{}\left[\inf_{\theta\in\Theta}h(\theta,X)}\right]\leqslant\inf_{\theta\in\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[h(\theta,X)}\right]$, we immediately obtain that:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right].$
In this sense, obtaining a guarantee on $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right]$ is stronger than on $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$.
Strengthening H 2.
We propose the following assumption:
H 4.
There exist $C_{0}>0$ and $\beta\in\mathopen{}\mathclose{{}\left[0,1}\right]$, such that for any $t\in[T]$ and $\varepsilon\in\mathopen{}\mathclose{{}\left[0,1}\right]$,
$\mathbb{P}\mathopen{}\mathclose{{}\left(\forall\theta\in\Theta,\>\inf_{j\neq I_{t}(\theta)\;}u_{j}(\theta,X_{t})-u_{I_{t}(\theta)}(\theta,X_{t})\geqslant\varepsilon\,\mid\,\mathcal{H}_{t-1}}\right)\geqslant 1-C_{0}\varepsilon^{\beta}\;,$
where $I_{t}(\theta)\in\operatorname*{\arg\!\min}_{i\in[K]}u_{i}(\theta,X_{t})$.
This reinforcement of H 2 comes from the fact that we took $\forall\theta\in\Theta$ inside the probability and not outside. Based on this assumption we adapt Lemma 3 to get the following lemma.
Lemma 1.
Assume H 4, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $\theta\in\Theta$ and $t\in[T]$,
$\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}\right\|_{1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)\;.$
Lemma 1 shows that assuming H 4 instead of H 2 transform the control of the supremum of an expectation of a supremum (Lemma 3) into one on the expectation of a supremum.
Exploiting Lemma 1, we prove a novel regret bound involving the deterministic dynamic regret.
Proposition 5.
Assume H 1, H 4, H 3, $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and that Algorithm 1 is instantiated with a non-increasing sequence $\mathopen{}\mathclose{{}\left(\eta_{t}}\right)_{t\in[T]}$. Then, denoting $P_{T}=\sum_{t=1}^{T-1}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|$,
$\frac{1}{T}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}G\eta_{t}}{32T\alpha_{t}^{2}}$
$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2\ln(d)C_{0}\{\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d)\}]+\xi\;,$
where $\mathbb{E}$ denotes the expectation on both data and the intrinsic randomness of DF-OGD.
From Proposition 5 we can recover the same convergence rate than Corollary 1 and also perform a similar study than in Section 4.2 to have results for a general polytope. The tradeoff highlighted is that strengthening H 2 into H 4 is the price to pay to control the stronger notion of regret to $\mathfrak{R}_{T}^{\mathrm{d}}$.
The proofs of Lemma 1 and Proposition 5 are deferred to appendix Section F.4.
Appendix E Regret bound with an approximate optimal solution.
In this section, we show that when we cannot access $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ but only $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ such that $\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),\underline{w}_{t}(\theta_{t})-w^{\star}_{t}(\theta_{t})}\right\rangle\leqslant\kappa$ for $\kappa>0$, our regret bounds are only shifted by $\kappa$. The following proof is for the general convex polytope case. The simplex case is established in the exact same way.
Proposition 6.
Assume the assumptions from Proposition 2 and that at each step, $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ is replaced by $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ which satisfies
$\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),\underline{w}_{t}(\theta_{t})-w^{\star}_{t}(\theta_{t})}\right\rangle\leqslant\kappa\;,$
where $\kappa>0$. Then,
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$ $\leqslant\frac{D_{\Theta}^{2}-D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{}\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}\right)^{2}$
$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2n\max_{k\in[K]}\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{1}C_{0}(\ln(1\alpha_{t}^{-1})+(1-\beta)^{2}\ln^{2}(\alpha_{t}\ln d))]+\xi+\kappa\;.$
The proof Proposition 6 is deferred to Section F.5.
Appendix F Proofs.
F.1 Proofs of Section 4.1.
Lemma 2.
Assume H 1, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}$-Lipchitz almost surely, with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$.
Proof.
Let $(\theta,\theta^{\prime})\in\Theta^{2}$. We have:
$\mathopen{}\mathclose{{}\left|\tilde{f}_{t}(\theta)-\tilde{f}_{t}(\theta^{\prime})}\right|$ $=\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)-\tilde{w}_{t}(\theta^{\prime})}\right\rangle$
$\leqslant\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|\mathopen{}\mathclose{{}\left\|\tilde{w}_{t}(\theta)-\tilde{w}_{t}(\theta^{\prime})}\right\|$
$=\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|\int_{t=0}^{1}\mathopen{}\mathclose{{}\left\|\nabla\tilde{w}_{t}(\theta+t(\theta^{\prime}-\theta))(\theta-\theta^{\prime})}\right\|\,\mathrm{d}t$
[…]
Now, denoting $\zeta=\theta+t(\theta^{\prime}-\theta)$, we have:
$\mathopen{}\mathclose{{}\left\|\nabla\tilde{w}_{t}(\zeta)}\right\|_{\mathrm{op}}$ $=\alpha_{t}^{-1}\mathopen{}\mathclose{{}\left\|\mathopen{}\mathclose{{}\left(\mathrm{diag}[\tilde{w}_{t}(\zeta)]-\tilde{w}_{t}(\zeta)\tilde{w}_{t}(\zeta)^{\mathsf{\scriptscriptstyle T}}}\right)\nabla_{\theta}g(\zeta,X_{t})}\right\|_{\mathrm{op}}$
$\leqslant\alpha_{t}^{-1}\mathopen{}\mathclose{{}\left\|\mathrm{diag}[\tilde{w}_{t}(\zeta)]-\tilde{w}_{t}(\zeta)\tilde{w}_{t}(\zeta)^{\mathsf{\scriptscriptstyle T}}}\right\|_{\mathrm{op}}\mathopen{}\mathclose{{}\left\|\nabla_{\theta}g(\zeta,X_{t})}\right\|_{\mathrm{op}}$
[…]
Since $\mathrm{diag}[\tilde{w}_{t}(\zeta)]-\tilde{w}_{t}(\zeta)\tilde{w}_{t}(\zeta)^{\mathsf{\scriptscriptstyle T}}$ is symmetric, its operator norm equals its largest eigenvalue, denoted $\lambda_{\max}$. It then follows that:
$\mathopen{}\mathclose{{}\left\|\mathrm{diag}[\tilde{w}_{t}(\zeta)]-\tilde{w}_{t}(\zeta)\tilde{w}_{t}(\zeta)^{\mathsf{\scriptscriptstyle T}}}\right\|_{\mathrm{op}}$ $=\lambda_{\max}(\mathrm{diag}[\tilde{w}_{t}(\zeta)]-\tilde{w}_{t}(\zeta)\tilde{w}_{t}(\zeta)^{\mathsf{\scriptscriptstyle T}})$
$\leqslant\max_{i\in[d]}\sum_{j\in[d]}\mathopen{}\mathclose{{}\left|\mathrm{diag}[\tilde{w}_{t}(\zeta)]_{i,j}-\tilde{w}_{t}(\zeta)\tilde{w}_{t}(\zeta)^{\mathsf{\scriptscriptstyle T}}_{i,j}}\right|$
$\leqslant\max_{i\in[d]}\{\tilde{w}_{t}(\zeta)_{i,i}(1-\tilde{w}_{t}(\zeta)_{i,i})+\tilde{w}_{t}(\theta)_{i,i}\sum_{j\neq i}\tilde{w}_{t}(\zeta)_{i,j}\}$
$\leqslant\frac{1}{4}+1=\frac{5}{4}\;,$
where the two last inequalities hold because $\tilde{w}_{t}\in[0,1]^{d}$. Therefore, we obtain from (18) that $\mathopen{}\mathclose{{}\left\|\nabla\tilde{w}_{t}(\zeta)}\right\|_{\mathrm{op}}\leqslant 5G(4\alpha_{t})^{-1}$, and plugging this in (17) yields:
$\mathopen{}\mathclose{{}\left|\tilde{f}_{t}(\theta)-\tilde{f}_{t}(\theta^{\prime})}\right|\leqslant\frac{5GD_{\mathcal{Z}}}{4\alpha_{t}}\int_{0}^{1}\mathopen{}\mathclose{{}\left\|\theta-\theta^{\prime}}\right\|\,\mathrm{d}t=\frac{5GD_{\mathcal{Z}}}{4\alpha_{t}}\mathopen{}\mathclose{{}\left\|\theta-\theta^{\prime}}\right\|\;.$
Lemma 3.
Assume H 2, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $\theta\in\Theta$ and $t\in[T]$, for any $\theta\in\theta$
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}\right\|_{1}\mid\mathcal{H}_{t-1}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)\;.$
Proof.
Let $\theta\in\Theta$ and $t\in[T]$. Since $\mathcal{W}=\mathcal{W}_{0}$, denoting for any $j\in[d]$ $e_{j}$ the $j$-th element of the canonical basis, we have $u_{j}(\theta,X_{t})=\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),e_{j}}\right\rangle$. We recall that we denote $u_{I_{t}(\theta)}(\theta,X_{t})=\min_{j\in[d]}u_{j}(\theta,X_{t})$. We first prove the following lemma.
Lemma 4.
Assume that there exists $\varepsilon>0$ such that $u_{I_{t}(\theta)}(\theta,X_{t})+\varepsilon\leqslant u_{j}(\theta,X_{t})\,$ for any $j\in\{1,\ldots,K\}\setminus\{I_{t}(\theta)\}$. Then,
$\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\leqslant\frac{2\alpha_{t}\ln(d)}{\varepsilon}\;.$
Proof.
In this proof, $I_{t}(\theta)$ from H 2 is denoted $I_{t}$ to lighten notation. We have by assumption:
$\langle g(\theta,X_{t}),\tilde{w}_{t}(\theta)\rangle$ $=\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),\sum_{j=1}^{d}\tilde{w}_{t,j}(\theta)\,e_{j}}\right\rangle$
$=u_{I_{t}}(\theta,X_{t})\tilde{w}_{t,I_{t}}(\theta)+\sum_{j\neq I_{t}}u_{j}(\theta,X_{t})\tilde{w}_{t,j}(\theta)$
$\geqslant u_{I_{t}}(\theta,X_{t})\tilde{w}_{t,I_{t}}(\theta)+(u_{I_{t}}(\theta,X_{t})+\varepsilon)\sum_{j\neq I_{t}}\tilde{w}_{t,j}(\theta)$
$=u_{I_{t}}(\theta,X_{t})\tilde{w}_{t,I_{t}}(\theta)+(u_{I_{t}}(\theta,X_{t})+\varepsilon)\mathopen{}\mathclose{{}\left(1-\tilde{w}_{t,I_{t}}(\theta)}\right)\;.$ (19)
The last equality holds because $\mathopen{}\mathclose{{}\left\|\tilde{w}_{t}(\theta)}\right\|_{1}=1$. On the other hand, using the facts that $\mathcal{R}_{0}\geqslant 0$, $\tilde{w}_{t}(\theta)=\operatorname*{\arg\!\min}\langle g(\theta,X_{t}),w\rangle+\alpha_{t}\mathcal{R}_{0}(w)$ and $u_{I_{t}}(\theta,X_{t})=\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),w^{\star}_{t}(\theta)}\right\rangle$ yields:
$\langle g(\theta,X_{t}),\tilde{w}_{t}(\theta)\rangle$ $\leqslant\langle g(\theta,X_{t}),\tilde{w}_{t}(\theta)\rangle+\alpha_{t}\mathcal{R}_{0}(\tilde{w}_{t}(\theta))$
$\leqslant\langle g(\theta,X_{t}),w^{\star}_{t}(\theta)\rangle+\alpha_{t}\mathcal{R}_{0}(w^{\star}_{t}(\theta))$
$=u_{I_{t}}(\theta,X_{t})+\alpha_{t}\mathcal{R}_{0}(w^{\star}_{t}(\theta))$
$\leqslant u_{I_{t}}(\theta,X_{t})+\alpha_{t}\ln(d).$ (20)
The last line holds because $\mathcal{R}_{0}(w)\leqslant\ln(d)$ for any $w\in\mathcal{W}_{0}$. Combining (19) and (20) gives:
$u_{I_{t}}(\theta,X_{t})\tilde{w}_{t,I_{t}}(\theta)+(u_{I_{t}}+\varepsilon)\mathopen{}\mathclose{{}\left(1-\tilde{w}_{t,I_{t}}(\theta)}\right)\leqslant u_{I_{t}}(\theta,X_{t})+\alpha_{t}\ln(d).$
Re-organizing the terms then yields $\tilde{w}_{t,I_{t}}(\theta)\geqslant 1-\frac{\alpha_{t}\ln(d)}{\varepsilon}$ and again, as $\mathopen{}\mathclose{{}\left\|\tilde{w}_{t}(\theta)}\right\|_{1}=1$, we have: $\sum_{j\neq I_{t}}\tilde{w}_{t,j}(\theta)\leqslant\frac{\alpha_{t}\ln(d)}{\varepsilon}$.
Combining those inequalities yields:
$\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}$ $=\|e_{I_{t}}-\tilde{w}_{t}(\theta)\|_{1}=1-\tilde{w}_{t,I_{t}}(\theta)+\sum_{j\neq I_{t}}\tilde{w}_{t,j}(\theta)\leqslant\frac{2\alpha_{t}\ln(d)}{\varepsilon}.$
To conclude the proof, we first write the expected distance as follows:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\mid\mathcal{H}_{t-1}}\right]$ $=\int_{0}^{+\infty}\mathbb{P}\mathopen{}\mathclose{{}\left(\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>y\,\mid\,\mathcal{H}_{t-1}}\right)\,\mathrm{d}y,$
Since both $w^{\star}_{t}(\theta)$ and $\tilde{w}_{t}(\theta)$ belong to the simplex, we can restrict the integral to:
$=\int_{0}^{2}\mathbb{P}\mathopen{}\mathclose{{}\left(\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>y\,\mid\,\mathcal{H}_{t-1}}\right)\,\mathrm{d}y$ (21)
$=\int_{0}^{\alpha_{t}}\mathbb{P}\mathopen{}\mathclose{{}\left(\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>y\,\mid\,\mathcal{H}_{t-1}}\right)\,\mathrm{d}y$ (22)
$+\int_{\alpha_{t}}^{2}\mathbb{P}\mathopen{}\mathclose{{}\left(\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>y\,\mid\,\mathcal{H}_{t-1}}\right)\,\mathrm{d}y$
$\leqslant\alpha_{t}+\int_{\alpha_{t}}^{2}\mathbb{P}\mathopen{}\mathclose{{}\left(\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>y\,\mid\,\mathcal{H}_{t-1}}\right)\,\mathrm{d}y$ (23)
We now apply the change of variable $y=2\alpha_{t}\ln(d)\varepsilon^{-1}$ to obtain:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\mid\mathcal{H}_{t-1}}\right]\\ =\alpha_{t}+2\alpha_{t}\ln(d)\int_{(\alpha_{t}\ln(d))}^{(2\ln(d))}\mathbb{P}\mathopen{}\mathclose{{}\left(\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>2\frac{\alpha_{t}\ln(d)}{\varepsilon}\,\mid\,\mathcal{H}_{t-1}}\right)\frac{d\varepsilon}{\varepsilon^{2}}.$ (24)
Moreover, for any $\varepsilon>0$ we have by H 2 and Lemma 4 that:
$1-C_{0}\varepsilon^{\beta}$ $\leqslant\mathbb{P}\mathopen{}\mathclose{{}\left(\inf_{j\neq I_{t}}\{u_{j}(\theta,X_{t})-u_{I_{t}}(\theta,X_{t})\}\geqslant\varepsilon\,\mid\,\mathcal{H}_{t-1}}\right)$
$\leqslant\mathbb{P}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}\leqslant 2\alpha_{t}\ln(d)\varepsilon^{-1}\,\mid\,\mathcal{H}_{t-1}}\right)\;,$
so it follows that:
$\mathbb{P}_{t}\mathopen{}\mathclose{{}\left(\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>2\frac{\alpha_{t}\ln(d)}{\varepsilon}\mid\mathcal{H}_{t-1}}\right)$ $=1-\mathbb{P}_{t}\mathopen{}\mathclose{{}\left(\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\leqslant 2\frac{\alpha_{t}\ln(d)}{\varepsilon}\mid\mathcal{H}_{t-1}}\right)$
$\leqslant C_{0}\varepsilon^{\beta}\;,$ (25)
Hence, plugging (25) in (24) gives:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\,\mid\,\mathcal{H}_{t-1}}\right]$ $\leqslant\alpha_{t}+2\alpha_{t}\ln(d)C_{0}\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon\;.$
We now proceed to bound $\varphi(\beta)=\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon$ for $\beta\in(0,1)$. With $a=\alpha_{t}\ln(d)$ and $b=2\ln(d)$, we have:
$\varphi(\beta)=(1-\beta)^{-1}(a^{\beta-1}-b^{\beta-1})=(1-\beta)^{-1}[e^{-(1-\beta)\ln(a)}-e^{-(1-\beta)\ln(b)}]\;,$
[…]
$\varphi(\beta)$ $\leqslant(1-\beta)^{-1}[1-(1-\beta)\ln(a)+(1-\beta)^{2}\ln^{2}(a)-1+(1-\beta)\ln(b)]$
$=\ln\mathopen{}\mathclose{{}\left(\frac{b}{a}}\right)+(1-\beta)\ln^{2}(a)\;.$
Substituting in the values of $a$ and $b$ yields the desired result. ∎
Lemma 5.
Assume that for any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}-$Lipschitz almost-surely and that the sequence of steps $(\eta_{t})_{t\geqslant 1}$ is non-increasing. Denote $\tilde{\mathfrak{R}_{T}}=\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})$. Then it holds almost surely that:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}_{T}}\right]\leqslant\frac{1}{2\eta_{T}}\mathopen{}\mathclose{{}\left(D_{\Theta}^{2}+2D_{\Theta}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|}\right)+\sum_{t=1}^{T}\frac{\eta_{t}}{2}K_{t}^{2}\;,$
where the expectation is taken over the sequence $(u_{1},\ldots,u_{T})$ in Algorithm 1.
Proof.
In what follows, we write $\mathbb{E}_{u_{t}}$ the expectation under the distribution $\text{Unif}([0,1])$ of $u_{t}$, and $\mathbb{E}_{u_{1:T}}$ the expectation under the joint distribution $\text{Unif}([0,1])^{\otimes T}$ of $(u_{1},\ldots,u_{T})$. For any $t\in[T]$, we have:
[…]
$=\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left\langle\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\nabla\tilde{f}_{t}(\vartheta_{t}+u_{t}(\theta_{t}-\vartheta_{t}))}\right],\theta_{t}-\vartheta_{t}}\right\rangle$
$=\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left\langle\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\tilde{\nabla}_{t}(u_{t})}\right],\theta_{t}-\vartheta_{t}}\right\rangle\;,$ (26)
where ~$\tilde{\nabla}(u_{t})=\nabla\tilde{f}_{t}(\vartheta_{t}+u_{t}(\theta_{t}-\vartheta_{t})$ as in Algorithm 1. To control (26) we note that by definition of $\theta_{t+1}$ in Algorithm 1,
$\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}}\right]$ $=\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\Pi_{\Theta}[\theta_{t}-\eta_{t}\tilde{\nabla}_{t}(u_{t})]-\vartheta_{t}}\right\|^{2}}\right]$
$\leqslant\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}-\eta_{t}\tilde{\nabla}_{t}(u_{t})}\right\|^{2}}\right]$
$\leqslant\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}}\right\|^{2}-2\eta_{t}\mathopen{}\mathclose{{}\left\langle\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\tilde{\nabla}_{t}(u_{t})}\right],\theta_{t}-\vartheta_{t}}\right\rangle+\eta_{t}^{2}K_{t}^{2}$ (27)
where we have used that ~$\mathopen{}\mathclose{{}\left\|\tilde{\nabla}_{t}(u_{t})}\right\|^{2}\leqslant K_{t}^{2}$ because $\tilde{f}_{t}$ is $K_{t}-$Lipschitz. Re-arranging (27) and taking the expectation over $(u_{1},\ldots,u_{T})$ yields:
$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\tilde{\nabla}_{t}(u_{t}),\theta_{t}-\vartheta_{t}}\right\rangle}\right]$ $\leqslant\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\frac{1}{2\eta_{t}}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}}\right\|^{2}-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}}\right)+\frac{\eta_{t}K_{t}^{2}}{2}}\right]\;,$
Now notice that, for any $t\in\{1,\ldots,T-1\}$,
$\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|^{2}$ $=\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}+\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|^{2}-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}$
$=\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}$
$+(\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|+\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|)(\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|)$
$\leqslant\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}+2D_{\Theta}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|\;,$ (29)
where we have used H 1-(i) and the reversed triangular inequality in the last line. It follows that:
$-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}$ $\leqslant-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|^{2}+2D_{\Theta}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|$ (30)
Plugging (30) into (F.1) then gives:
$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\tilde{\nabla}_{t}(u_{t}),\theta_{t}-\vartheta_{t}}\right\rangle}\right]$ $\leqslant\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\frac{1}{2\eta_{t}}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}}\right\|^{2}-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|^{2}}\right)}\right.$
$\quad\mathopen{}\mathclose{{}\left.+\frac{1}{\eta_{t}}D_{\theta}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|+\frac{\eta_{t}K_{t}^{2}}{2}}\right]\;.$
Thus, summing over $T$ and plugging the resulting sum in (26) gives:
$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\sum_{t=1}^{T}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}\right]$ $\leqslant\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\sum_{t=1}^{T}\frac{\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}}\right\|^{2}}{2}\mathopen{}\mathclose{{}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}}\right)}\right.$
$\quad\mathopen{}\mathclose{{}\left.+\sum_{t=1}^{T}\frac{1}{\eta_{t}}D_{\theta}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|+\sum_{t=1}^{T}\frac{\eta_{t}K_{t}^{2}}{2}}\right]\;.$
$\leqslant\frac{D_{\Theta}^{2}}{2}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}}\right)+\frac{D_{\theta}}{\eta_{T}}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|+\sum_{t=1}^{T}\frac{\eta_{t}K_{t}^{2}}{2}\;.$
where we used that for all $t$, $\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}}\right\|\leqslant D_{\Theta}$ by H 1-(i) and that for all $t\in[T],\,\eta_{t-1}\geqslant\eta_{t}\geqslant\eta_{T}$. Then, we have by telescoping the first sum:
$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\sum_{t=1}^{T}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}\right]$ $\leqslant\frac{D_{\Theta}^{2}+2D_{\Theta}}{2\eta_{T}}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|+\sum_{t=1}^{T}\frac{\eta_{t}K_{t}^{2}}{2}.$
See 1
Proof.
In this proof, for the sake of conciseness, we denote by $\mathbb{E}_{t}$ the conditional expectation $\mathbb{E}\mathopen{}\mathclose{{}\left[\cdot\mid\mathcal{H}_{t-1}}\right]$.
Observe that the dynamic regret can be decomposed as follows:
$\mathfrak{R}_{T}$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}\right\rangle}\right]+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle}\right]$ (31)
$+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\vartheta_{t})}\right\rangle}\right]-\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle}\right]$ (32)
$+\sum_{t\in[T]}\mathopen{}\mathclose{{}\left[\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle}\right]-\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}\right\rangle}\right]}\right]$ (33)
Taking the $\sup$ over $\Theta$, defining by $\mathrm{Reg}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle}\right]$ in (31), using H 3 in (32), and noticing that $\inf_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)]-\inf_{\theta\in\Theta}\mathbb{E}_{t}[\tilde{f}_{t}(\theta)]\leqslant\sup_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)-\tilde{f}_{t}(\theta)]$ in (33) leads to:
$\mathfrak{R}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\rangle}\right]+\xi T$
Using the fact that $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{\infty}$ is dual to $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{1}$:
$\mathfrak{R}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]+\xi T$
Since for any $t\in[T]$, $\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{2}\leqslant D_{\mathcal{Z}}$ almost surely by H 1-(ii),
$\mathfrak{R}_{T}$ $\leqslant\mathrm{Reg}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]+\xi T\;$ (34)
First, notice that, by Lemma 3, for all $t$:
$\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]$ $\leqslant\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)\;.$ (35)
Second, taking the expectation (denoted by $\mathbb{E}$) over $(u_{t})_{t\in[T]}$ and $X_{1},\cdots X_{T}$ on both sides gives:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\mathrm{Reg}_{T}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)+\xi T\;.$ (36)
To bound the first term in the right-hand side of Equation 36. Remark that, by the definition of conditional expectation:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathrm{Reg}_{T}}\right]=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}\right].$
One recognises the definition of $\tilde{\mathfrak{R}_{T}}$ of Lemma 5. Thus, by this lemma, we know that for any $t\in[T]$, because $\tilde{f}_{t}$ is $K_{t}-$Lipschitz with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$ almost surely by Lemma 2, we have:
$\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}\right]\leqslant\frac{1}{2\eta_{T}}(D_{\Theta}^{2}+D_{\Theta}P_{T})+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}^{2}G^{2}\eta_{t}}{32\alpha_{t}^{2}}\;.$ (37)
Plugging Equation 37 into Equation 36 and then dividing by $T>0$ on both sides gives the desired result.
See 1
Proof.
We know by Proposition 1 that for any decreasing sequence of learning rates $(\eta_{1},\ldots,\eta_{T})\in\mathbb{R}_{+}^{T}$ and any regularization coefficients $(\alpha_{1},\ldots,\alpha_{T})\in\mathbb{R}_{+}^{T}$, the average regret of Algorithm 1 is bounded as follows:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{1+P_{T}}{\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{\alpha_{t}^{2}}+(1+\ln d)\sum_{t\in[T]}\alpha_{t}+\xi T}\right]}\right)\;.$
With $\alpha_{t}\propto t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto\alpha_{t}t^{-1/2}(1+P_{t})^{1/2}\propto t^{-3/4}(1+P_{t})^{3/4}$, the sequence $(\eta_{t})_{t\in[T]}$ is non-increasing because so is $(t^{-1}(1+P_{t}))_{t\in[T]}$ by assumption. Moreover, the first term satisfies:
$\frac{1+P_{T}}{\eta_{T}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\frac{T^{3/4}}{(1+P_{T})^{3/4}}\mathopen{}\mathclose{{}\left(1+P_{T}}\right)}\right)\quad\text{so}\quad\frac{1+P_{T}}{T\eta_{T}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(\frac{1+P_{T}}{T}}\right)^{1/4}}\right)\;.$
For the second term,
$\sum_{t=1}^{T}\frac{\eta_{t}}{\alpha_{t}^{2}}=\sum_{t=1}^{T}\frac{\sqrt{1+P_{t}}}{\sqrt{t}\alpha_{t}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left(\frac{1+P_{t}}{t}}\right)^{1/4}}\right)\;..$
Moreover, since $1+P_{t}\leqslant 1+P_{T}$ for any $t\in[T]$, we have:
$\sum_{t=1}^{T}\frac{\eta_{t}}{\alpha_{t}^{2}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(1+P_{T}}\right)^{1/4}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left(\frac{1}{t}}\right)^{1/4}}\right)=\mathcal{O}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(1+P_{T}}\right)^{1/4}T^{3/4}}\right)\;,$
where we used in the last line that $\sum_{t=1}^{T}t^{-\gamma}=\mathcal{O}\mathopen{}\mathclose{{}\left(T^{1-\gamma}}\right)$ for $\gamma\in(0,1)$. Dividing by $T$ again gives a rate of $\mathcal{O}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(\frac{1+P_{T}}{T}}\right)^{1/4}}\right)$ for the second term. Finally, we obtain by the same reasoning that
$\sum_{t\in[T]}\alpha_{t}\leqslant(1+P_{T})^{1/4}\sum_{t\in[T]}t^{-1/4}=\mathcal{O}((1+P_{T})^{1/4}T^{3/4})\;,$
and dividing by $T>0$ yields the desired rate. ∎
F.2 Proofs of Section 4.2.
Lemma 6.
Assume H 1, $\mathcal{W}=\mathcal{W}_{1}$ and $\mathcal{R}=\mathcal{R}_{1}$. Then for any $t\in[T]$,$\tilde{f}_{t}$ is differentiable and for any $\theta\in\Theta$, $\nabla\tilde{f}_{t}(\theta)=\nabla\tilde{w}_{t}(\theta)\bar{g}_{t}(X_{t})$ where:
$\nabla\tilde{w}_{t}(\theta)=-\alpha_{t}^{-1}\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}(b_{i}-A_{i}^{\mathsf{\scriptscriptstyle T}}\tilde{w}_{t}(\theta))^{-2}A_{i}A_{i}^{\mathsf{\scriptscriptstyle T}}}\right)\nabla_{\theta}g(\theta,X_{t})\;.$
Proof.
Let $\theta\in\mathbb{R}$ and $t\in[T]$. With
$h_{t}:(w,\theta)\mapsto\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),w}\right\rangle+\alpha_{t}\mathcal{R}_{1}(w)\;,$
we have $h_{t}(w)\rightarrow+\infty$ as $w\rightarrow\text{bdry}(\mathcal{W}_{1})$. Since $h_{t}(\tilde{w}_{t}(\theta),\theta)=\min_{w\in\mathcal{W}_{1}}h_{t}(w,\theta)$, we deduce that $\tilde{w}_{t}(\theta)\in\text{int}(\mathcal{W}_{1})$. It follows from the first order condition and the implicit function theorem [21, Theorem 2] that $\tilde{w}_{t}:\theta\mapsto\tilde{w}_{t}(\theta)$ is differentiable, and
$\nabla\tilde{w}_{t}(\theta)$ $=-(\nabla^{2}_{ww}h_{t}(\tilde{w}_{t}(\theta),\theta))^{-1}\nabla^{2}_{\theta w}h_{t}(\tilde{w}_{t}(\theta),\theta)$
$=-\frac{1}{\alpha_{t}}[\nabla^{2}\mathcal{R}_{1}(\tilde{w}_{t}(\theta))]^{-1}\nabla_{\theta}g(\theta,X_{t})\;.$ (38)
Since $\mathcal{R}_{1}(w)=-\sum_{i=1}^{n}\ln(b_{i}-A_{i}w)$, simple computations give:
$\nabla^{2}\mathcal{R}_{1}(\tilde{w}_{t}(\theta))=\sum_{i=1}^{n}(b_{i}-A_{i}^{\mathsf{\scriptscriptstyle T}}w_{t}(\theta))^{-2}A_{i}A_{i}^{\mathsf{\scriptscriptstyle T}}\;,$ (39)
Moreover, since $\text{rank}(AA^{\mathsf{\scriptscriptstyle T}})=d$ by assumption, $\nabla^{2}\mathcal{R}_{1}(\tilde{w}_{t}(\theta))$ is invertible. Plugging (39) in (38) gives the result. ∎
Lemma 7.
Assume H 1 and $\mathcal{R}=\mathcal{R}_{1}$. For any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}$-Lipschitz almost-surely, with
$K_{t}=\frac{GD_{\mathcal{Z}}}{\alpha_{t}}\frac{\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}}{\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}\;.$
Proof.
   Let $(\theta,\theta^{\prime})\in\Theta^{2}$. We proved in (17) that:
$\mathopen{}\mathclose{{}\left|\tilde{f}_{t}(\theta)-\tilde{f}_{t}(\theta^{\prime})}\right|\leqslant\,D_{\mathcal{Z}}\int_{t=0}^{1}\mathopen{}\mathclose{{}\left\|\nabla\tilde{w}_{t}(\theta+t(\theta^{\prime}-\theta))}\right\|_{\mathrm{op}}\mathopen{}\mathclose{{}\left\|\theta-\theta^{\prime}}\right\|\,\mathrm{d}t\;.$
Denoting $\zeta=\theta+t(\theta^{\prime}-\theta)$, by (38), since $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{\mathrm{op}}$ is sub-multiplicative:
$\mathopen{}\mathclose{{}\left\|\nabla\tilde{w}_{t}(\zeta)}\right\|_{\mathrm{op}}$ $=-\alpha_{t}^{-1}\mathopen{}\mathclose{{}\left\|[\nabla^{2}\mathcal{R}_{1}(\tilde{w}_{t}(\zeta)]^{-1}\nabla_{\theta}g(\theta,X_{t})}\right\|_{\mathrm{op}}$
[…]
Moreover, we have:
$\mathopen{}\mathclose{{}\left\|[\nabla^{2}\mathcal{R}_{1}(\zeta)]^{-1}}\right\|_{\mathrm{op}}$ $=\mathopen{}\mathclose{{}\left\|\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}(b_{i}-A_{i}^{\mathsf{\scriptscriptstyle T}}\tilde{w}_{t}(\zeta))^{-2}A_{i}A_{i}^{\mathsf{\scriptscriptstyle T}}}\right)^{-1}}\right\|_{\mathrm{op}}$
$\leqslant\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}^{2}\mathopen{}\mathclose{{}\left\|(AA^{\mathsf{\scriptscriptstyle T}})^{-1}}\right\|_{\mathrm{op}}\;,$
Since $AA^{\mathsf{\scriptscriptstyle T}}$ is symmetric, $\mathopen{}\mathclose{{}\left\|(AA^{\mathsf{\scriptscriptstyle T}})^{-1}}\right\|_{\mathrm{op}}=\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})^{-1}$, where $\lambda_{\min}(M)$ is the lowest eigenvalue of $M$. We therefore obtain by (17) and the previous inequalities that almost-surely:
$\mathopen{}\mathclose{{}\left|\tilde{f}_{t}(\theta)-\tilde{f}_{t}(\theta^{\prime})}\right|$ $\leqslant\frac{GD_{\mathcal{Z}}}{\alpha_{t}}\frac{\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}^{2}}{\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}\int_{t=0}^{1}\mathopen{}\mathclose{{}\left\|\theta-\theta^{\prime}}\right\|\,\mathrm{d}t$
$=\frac{GD_{\mathcal{Z}}}{\alpha_{t}}\frac{\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}^{2}}{\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}\mathopen{}\mathclose{{}\left\|\theta-\theta^{\prime}}\right\|\;.$
Lemma 8.
Assume H 2, $\mathcal{W}=\mathcal{W}_{1}$ and $\mathcal{R}=\mathcal{R}_{1}$. Then for any $t\in[T]$ and $\theta\in\Theta$,
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}\right\|_{1}\mid\mathcal{H}_{t-1}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{}\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)\;.$
Proof.
Let $\theta\in\Theta$ and $t\in[T]$. Since $\mathcal{W}_{1}=\mathrm{Conv}\mathopen{}\mathclose{{}\left(v_{1},\ldots,v_{K}}\right)$, there exists $(\lambda_{t,I_{t}}(\theta),\ldots,\lambda_{t,I_{t}}(\theta))\in[0,1]^{K}$ such that
$\tilde{w}_{t}(\theta)=\sum_{i=1}^{K}\lambda_{t,i}(\theta)v_{i}\quad\text{and}\quad\sum_{i=1}^{K}\lambda_{t,i}(\theta)=1\;.$
We recall that $u_{j}(\theta,X_{t})=\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),v_{j}}\right\rangle$ denotes the value of the objective function on the vertex $v_{j}\in\mathcal{W}$, and $u_{I_{t}}(\theta,X_{t})=\min_{j\in[K]}u_{j}(\theta,X_{t})$ so $v_{I_{t}}=w^{\star}_{t}(\theta)$. We start by proving the following lemma:
Lemma 9.
If there exists $\varepsilon>0$ such that $u_{I_{t}}+\varepsilon\leqslant u_{j}\,$ for any $j\in\{1,\ldots,K\}\setminus\{I_{t}\}$, then,
$\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}\leqslant\frac{2n\alpha_{t}}{\varepsilon}\max_{k\in[K]}\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{1}\;.$
Proof.
On the one hand, $\tilde{w}_{t}(\theta)$ is by definition the solution to the problem:
$\tilde{w}_{t}(\theta)=\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),w}\right\rangle-\alpha_{t}\sum_{i=1}^{d}\ln(b_{i}-A_{i}w)\;,$ (40)
which is uniquely determined by strong convexity of the objective. We know by [15, page 566] that:
$\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),\tilde{w}_{t}(\theta)}\right\rangle-\inf_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),w}\right\rangle\leqslant n\alpha_{t}\;,$
that is
$\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),\tilde{w}_{t}(\theta)}\right\rangle\leqslant u_{I_{t}}+n\alpha_{t}\;.$ (41)
On the other hand, by assumption we know that for any $j\neq I_{t}$, $u_{j}\geqslant u_{I_{t}}+\varepsilon$ so:
$\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),\tilde{w}_{t}(\theta)}\right\rangle$ $=\mathopen{}\mathclose{{}\left\langle g(\theta,X_{t}),\sum_{j=1}^{K}\lambda_{t,j}(\theta)v_{j}}\right\rangle$ (42)
$=\lambda_{t,I_{t}}(\theta)u_{I_{t}}+\sum_{j\neq I_{t}}\lambda_{t,j}(\theta)u_{j}$
$\geqslant\lambda_{t,I_{t}}(\theta)u_{I_{t}}+\sum_{j\neq I_{t}}\lambda_{t,j}(\theta)(u_{I_{t}}+\varepsilon)$
$=\lambda_{t,I_{t}}(\theta)u_{I_{t}}+(u_{I_{t}}+\varepsilon)(1-\lambda_{t,I_{t}}(\theta))\;.$ (43)
Combining (41) and (43) yields:
$\lambda_{t,I_{t}}(\theta)\geqslant 1-\frac{n\alpha_{t}}{\varepsilon}\;,$ (44)
and it follows from (44) that:
$\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}=\mathopen{}\mathclose{{}\left\|v_{I_{t}}-\tilde{w}_{t}(\theta)}\right\|_{1}$ $=\mathopen{}\mathclose{{}\left\|\sum_{j=1}^{K}(\mathbbm{1}_{\{j=I_{t}\}}-\lambda_{t,j}(\theta))v_{j}}\right\|_{1}$
$\leqslant[(1-\lambda_{t,I_{t}}(\theta))+\sum_{j\neq I_{t}}\lambda_{t,j}(\theta)]\max_{k\in[K]}\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{1}$
$\leqslant\frac{2n\alpha_{t}}{\varepsilon}\max_{k\in[K]}\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{1}\;.$
The result follows from combining H 2 and Lemma 9 according to the same lines of computation as in the proof of Lemma 3. ∎
See 2
Proof.
In this proof, for the sake of conciseness, we use the notation $\mathbb{E}_{t}$ to denote $\mathbb{E}\mathopen{}\mathclose{{}\left[\cdot\mid\mathcal{H}_{t-1}}\right]$. We showed in (34) of the proof of Proposition 1 that if we define $\mathrm{Reg}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle}\right]$, then we have:
$\mathfrak{R}_{T}$ $\leqslant\mathrm{Reg}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]+\xi T.$
First, by Lemma 8, we know have for any $t$:
$\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{}\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)\;.$
Thus combining the two last equations and taking the expectation (denoted by $\mathbb{E}$) over $(u_{t})_{t\in[T]}$ and $X_{1},\cdots X_{T}$ on both sides gives:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\mathrm{Reg}_{T}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}\mathopen{}\mathclose{{}\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)+\xi T\;.$ (45)
Just as in the proof of Proposition 1, we can bound the first term in Equation 45 by noting that the definition of conditional expectation implies:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathrm{Reg}_{T}}\right]=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}\right].$
One recognizes the definition of $\tilde{\mathfrak{R}_{T}}$ of Lemma 5. Since $\tilde{f}_{t}$ is $K_{t}-$Lipschitz with $K_{t}=\frac{GD_{\mathcal{Z}}}{\alpha_{t}}\frac{\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}}{\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}$ almost surely for any $t$ by Lemma 7, we deduce from Lemma 5 that:
$\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}\right]\leqslant\frac{D_{\Theta}^{2}-D_{\Theta}P_{T}}{2\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{}\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}^{2}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}\right)^{2}\;.$ (46)
Plugging Equation 46 into Equation 45 and then dividing by $T>0$ on both sides gives the desired result. ∎
See 2
Proof.
We know by Proposition 2 that the regret of Algorithm 1 satisfies the following asymptotic bound:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{1+P_{T}}{\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{\alpha_{t}^{2}}+n\sum_{t\in[T]}\alpha_{t}+\xi T}\right]}\right)\;.$
The rest of the proof proceeds exactly as the one of Corollary 1. ∎
F.3 Proofs of Appendix B.
See 4
Proof.
For the sake of conciseness we use the noation $\mathbb{E}_{t}$ to denote $\mathbb{E}\mathopen{}\mathclose{{}\left[\cdot\mid\mathcal{H}_{t-1}}\right]$ We have, for any $\theta$, the following decomposition of the static regret
$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}\right\rangle}\right]$ (47)
$+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}\right\rangle}\right]-\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle}\right]$ (48)
$+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle}\right]-\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}\right\rangle}\right]$ (49)
Then, taking the $\sup$ over $\Theta$ in Equations 47 and 49, and defining
$\mathrm{Reg}^{\mathrm{S}}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}\right\rangle}\right]-\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle}\right],$
in Equation 48 yields for any $\theta$:
$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\rangle}\right]$
Using the fact that $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{\infty}$ is dual to $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{1}$:
$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]$
Since for any $t\in[T]$, $\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{2}\leqslant D_{\mathcal{Z}}$ by H 1-(ii),
$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}\;$
First, notice that, by Lemma 3, and because for any $t,\alpha_{t}=\alpha$ we have for all $t$:
$\mathbb{E}_{X_{1},\cdots,X_{T}}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]$ $\leqslant 2D_{\mathcal{Z}}\alpha\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}\right)}\right)\;.$ (50)
Second, taking the expectation (denoted by $\mathbb{E}$) over $(\sigma_{t})_{t\in[T]}$ and $X_{1},\cdots X_{T}$ on both sides gives:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}^{\mathrm{S}}_{T}(\theta)}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\mathrm{Reg}^{\mathrm{S}}_{T}}\right]+2D_{\mathcal{Z}}\alpha T\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}\right)}\right).$ (51)
To bound the first term in the right-hand-side of Equation 51. by the definition of conditional expectation, we have:
$\mathrm{Reg}^{\mathrm{S}}_{T}$ $=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}\right\rangle-\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle}\right]$
Then,
$\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}^{\mathrm{S}}_{T}}\right]$ $\leqslant\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}\right].$
One recognises (up to a factor $T$) the left-hand side of Proposition 3 on the loss sequence $(\tilde{f}_{t})_{t\in[T]}$. Furthermore, we can use this proposition as, given our choice of $\mathcal{R}$, for any $t\in[T]$, $\tilde{f}_{t}$ is $L-$Lipschitz with $L=5D_{\mathcal{Z}}G(4\alpha)^{-1}$ almost surely by Lemma 2. We then have:
$\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{}\left[\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}\right]\leqslant\mathcal{O}\mathopen{}\mathclose{{}\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi m\frac{1}{\alpha}}\right)\;.$ (52)
Dividing Equation 51 by $T$, and plugging Equation 52 into it concludes the proof. ∎
F.4 Proofs of Appendix D.
Proof of Lemma 1.
The proof starts exactly in the same way as the one of Lemma 3, where proved in (24) that:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\mid\mathcal{H}_{t-1}}\right]\\ =\alpha_{t}+2\alpha_{t}\ln(d)\int_{(\alpha_{t}\ln(d))}^{(2\ln(d))}\mathbb{P}\mathopen{}\mathclose{{}\left(\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>2\frac{\alpha_{t}\ln(d)}{\varepsilon}\,\mid\,\mathcal{H}_{t-1}}\right)\frac{d\varepsilon}{\varepsilon^{2}}.$ (53)
Moreover, we have for any $\varepsilon>0$:
$1-C_{0}\varepsilon^{\beta}$ $\leqslant\mathbb{P}\mathopen{}\mathclose{{}\left(\forall\theta\in\Theta,\;\inf_{j\neq I_{t}}u_{j}(\theta,X_{t})-u_{I_{t}}(\theta,X_{t})\geqslant\varepsilon\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right)$
$\leqslant\mathbb{P}\mathopen{}\mathclose{{}\left(\forall\theta\in\Theta,\;\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}{\mathcal{H}_{t-1}}\leqslant 2\alpha_{t}\ln(d)\varepsilon^{-1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right)$
$\leqslant\mathbb{P}\mathopen{}\mathclose{{}\left(\;\sup_{\theta\in\Theta}\;\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}\leqslant 2\alpha_{t}\ln(d)\varepsilon^{-1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right)$
where the first inequality stems from H 4, and the second one from Lemma 4. It follows that:
$\mathbb{P}\mathopen{}\mathclose{{}\left(\sup_{\theta\in\Theta}\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>2\frac{\alpha_{t}\ln(d)}{\varepsilon}\mid\mathcal{H}_{t-1}}\right)$ $=1-\mathbb{P}\mathopen{}\mathclose{{}\left(\sup_{\theta\in\Theta}\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\leqslant 2\frac{\alpha_{t}\ln(d)}{\varepsilon}\mid\mathcal{H}_{t-1}}\right)$
$\leqslant C_{0}\varepsilon^{\beta}\;,$ (54)
Hence, plugging (54) in (24) gives:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right]$ $\leqslant\alpha_{t}+2\alpha_{t}\ln(d)C_{0}\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon\;.$
And the result is established by bounding $\varphi(\beta)=\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon$ in the same way as in the proof of Lemma 3. ∎
Proof of Proposition 5.
Observe that the pseudo dynamic regret can be decomposed as follows:
$\mathfrak{R}_{T}^{\mathrm{d}}$ $=\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}\right\rangle+\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle$ (55)
$+\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\vartheta_{t})}\right\rangle-\inf_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle$ (56)
$+\sum_{t\in[T]}\mathopen{}\mathclose{{}\left[\inf_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle-\inf_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}\right\rangle}\right]$ (57)
Taking the $\sup$ over $\Theta$ and defining by ~$\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}:=\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle$ in (55), using H 3 in (56), and noticing that $\inf_{\theta\in\Theta}f_{t}(\theta)-\inf_{\theta\in\Theta}\tilde{f}_{t}(\theta)\leqslant\sup_{\theta\in\Theta}f_{t}(\theta)-\tilde{f}_{t}(\theta)$ in (57) leads to:
$\mathfrak{R}_{T}^{\mathrm{d}}$$\leqslant\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\rangle+\xi T$
Using the fact that $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{\infty}$ is dual to $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{1}$:
$\mathfrak{R}_{T}^{\mathrm{d}}$$\leqslant\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}+\xi T$
Observing that $\lVert\bar{g}_{t}(X_{t})\rVert_{\infty}\leqslant\lVert\bar{g}_{t}(X_{t})\rVert_{2}\leqslant D_{\mathcal{Z}}$ almost surely by H 1-(ii), and taking the expectation over $X_{1},\ldots,X_{T}$ and $u_{1},\ldots,u_{T}$ then gives:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right]$ $\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}}\right]+2D_{\mathcal{Z}}\sum_{t\in[T]}\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]+\xi T\;$ (58)
We now proceed to bound the two first terms in (58). First, notice that by Lemma 1:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]$ $\leqslant\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)\;.$ (59)
Second, noticing that $\tilde{f}_{t}$ is $K_{t}-$Lipschitz with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$ almost surely by Lemma 2, we have by Lemma 5:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}}\right]\leqslant\frac{1}{2\eta_{T}}(D_{\Theta}^{2}+D_{\Theta}P_{T})+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}^{2}G^{2}\eta_{t}}{32\alpha_{t}^{2}}\;.$ (60)
Plugging (59) and (60) in (58) and dividing by $T$ on both sides gives the desired result. ∎
F.5 Proofs of Appendix E
Proof of Proposition 6.
The regret simply decomposes as follows:
$\mathfrak{R}_{T}$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\underline{w}_{t}(\theta_{t})}\right\rangle}\right]-\sum_{t\in[T]}\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}\right\rangle}\right]$
$=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\underline{w}_{t}(\theta_{t})-w^{\star}_{t}(\theta_{t})}\right\rangle}\right]$
$+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})}\right\rangle}\right]-\sum_{t\in[T]}\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}\right\rangle}\right]$
$\leqslant\kappa+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})}\right\rangle}\right]-\sum_{t\in[T]}\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}\right\rangle}\right]\;,$
and the proof continues as in Proposition 2. ∎
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: [Yes] Yes, the main claims in the abstract and introduction accurately reflect the paper’s contributions and scope. They clearly summarize the key methods, findings, and the intended impact, aligning well with the content presented in the body of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: All the assumptions underlying the analysis are highlighted and clearly discussed (in particular H 2, which has a dedicated discussion appendix in Appendix D). The scope of our results is also extensively discussed, see the comments after Proposition 1, Proposition 2, Corollary 1, Corollary 2. Likewise, the experimental setting is clearly described.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We checked the correctness of all our proofs. The assumptions are systematically recalled in the propositions and lemmas, see the statements of Proposition 1, Proposition 2, Corollary 1, Corollary 2.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The experimental setting is extensively described in discussed in Appendix C.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All our experiments are easily reproducible, as our experimental setting is precisely described in Appendix C.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, see Appendix C.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes] .
  34. Justification: see the plots in Section 5.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Our experiments are particularly light and can be run on any personal computer in reasonable time.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We made sure that the research conducted in the paper conforms in every aspcet with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss in the introduction (Section 1) the potential impact of online Decision-Focused learning in several fields of Operation Research, including logistics, healthcare operation or supply chain management.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: There is no need for such safeguards as we only use simple, synthetic data for our experiments.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We explicitly cite all the papers upon which the current work builds.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our algorithm is thoroughly explained in the main text of this paper (see Section 3).
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not have any crowdsourcing experiment or research with human subjects in this paper.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Same answer as before.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: No LLMs were used in this study.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 Instead of merely optimizing for predictive accuracy, DFL trains models to directly minimize the loss associated with downstream decisions.
-This end-to-end strategy holds promise for tackling complex combinatorial problems; however, existing studies focus solely on scenarios where a fixed batch of data is available and the objective function does not change over time.
+However, existing studies focus solely on scenarios where a fixed batch of data is available and the objective function does not change over time.
 We instead investigate DFL in dynamic environments where the objective function and data distribution evolve over time.
-This setting is challenging because the objective function has zero or undefined gradients—which prevents the use of standard first-order optimization methods—and is generally non-convex.
-To address these difficulties, we (i) regularize the objective to make it differentiable and (ii) make use of the optimism principle, based on a near-optimal oracle along with an appropriate perturbation.
-This leads to a practical online algorithm for which we establish bounds on the expected dynamic regret, both when the decision space is a simplex and when it is a general bounded convex polytope.
-1 CMAP, CNRS, École polytechnique
-2 Inria Saclay, Université Paris Saclay, LMO
-3 Inria Paris, Ecole Normale Supérieure, PSL
+This setting is challenging for online learning because the objective function has zero or undefined gradients—which prevents the use of standard first-order optimization methods—and is generally non-convex.
+To address these difficulties, we (i) regularize the objective to make it differentiable and (ii) use perturbation techniques along with a near-optimal oracle to overcome non-convexity.
+Combining those techniques yields two original online algorithms tailored for DFL, for which we establish respectively static and dynamic regret bounds.
+These are the first provable guarantees for the online decision-focused problem.
+Finally, we showcase the effectiveness of our algorithms on a knapsack experiment, where they outperform two standard benchmarks.
 1 Introduction.
-Many real-world decision problems involve uncertainty, and a common approach to handling it is through the predict-then-optimize framework bertsimas2020predictive .
+Many real-world decision problems involve uncertainty, and a common approach to handling it is through the predict-then-optimize framework [10].
 First, a prediction model is trained on historical data; then, its output is fed into an optimization problem to guide decision-making.
-This natural strategy has been successfully applied in many operation research (OR) problems, ranging from supply chain management acimovic2015making ; fisher2016optimal ; ban2019big ; bertsimas2020predictive and revenue management farias2013nonparametric ; ferreira2016analytics ; cohen2017impact ; chen2022statistical to healthcare operation bertsimas2013fairness ; aswani2019data ; gupta2020maximizing ; rath2017integrated , see mivsic2020data for an extensive review.
+This natural strategy has been successfully applied in many operation research (OR) problems, ranging from supply chain management [2, 28, 7, 10] and revenue management [26, 27, 18, 16] to healthcare operation [11, 5, 34, 61], see Mivsić and Perakis [56] for an extensive review.
 It is clear that this approach would yield optimal decisions if the predictions were perfectly accurate.
 However, in practice, prediction errors are inevitable—and even small inaccuracies can propagate through the optimization process, potentially leading to poor decisions.
-To address this limitation, an approach known as decision-focused learning mandi2024decision —also referred to as smart predict-then-optimize elmachtoub2022smart or integrated learning-optimization sadana2025survey —has emerged.
+To address this limitation, an approach known as decision-focused learning [54]—also referred to as smart predict-then-optimize [25] or integrated learning-optimization [62]—has emerged.
 Instead of optimizing for prediction accuracy alone, this method trains the predictive model to directly minimize the downstream decision loss.
 By aligning the learning objective with the decision-making goal, it produces models that are more robust to prediction errors in practical applications.
-While decision-focused learning yields strong empirical performance donti2017task ; verma2022case ; verma2023restless ; wang2023scalable , theoretical development has so far been limited to the batch setting, where models are trained on pre-collected, independently and identically distributed (i.i.d.) data wilder2018melding ; mandi2022decision ; shah2022decision ; schutte2024robust .
-This assumption breaks down in many real-world scenarios involving dynamic environments cheung2019learning ; padakandla2020reinforcement and shifting data distributions lu2018learning ; quinonero2022dataset .
-Online learning cesa2006prediction ; hazan2023introductiononlineconvexoptimization provides a general way to cope with such non-stationarity of data-generating processes.
+While decision-focused learning yields strong empirical performance [23, 71, 72, 73], theoretical development has so far been limited to the batch setting, where models are trained on pre-collected, independently and identically distributed (i.i.d.) data [74, 53, 64, 63].
+This assumption breaks down in many real-world scenarios involving dynamic environments [17, 57] and shifting data distributions [50, 60].
+Online learning [15, 37] provides a general way to cope with such non-stationarity of data-generating processes.
 This framework considers a learner who makes decisions sequentially, at each round leveraging data collected from previous rounds to inform its next decision.
@@
 This presents a significant technical challenge, as the inherent difficulties of decision-focused learning—such as the non-differentiability of the objective function or the lack of convexity of the losses due to the bi-level nature of the problem—compound those already present in online learning.
-The algorithm that we develop addresses these non-differentiability and non-convexity challenges by leveraging a regularization-based approach combined with perturbation-driven optimism.
-More precisely, our contributions are as follows:
+Our contributions are as follows:
   1. (i)
@@
   2. (ii)
-    We present a novel algorithm—Decision-Focused Online Gradient Descent (DF-OGD) for solving this problem.
-Our approach relies on two main ingredients: (a) regularizing the inner problem described above to make the resulting decision differentiable, and (b) applying an original gradient-based method to the regularized outer problem, leveraging optimistic online learning rakhlin2013online together with perturbation techniques suggala2020online to handle the non-convexity of the problem.
+    We present two algorithms to tackle this problem, Decision-Focused Follow-the-Perturbed-Leader (DF-FTPL) and Decision-Focused Online Gradient Descent (DF-OGD).
+While both rely on regularizing the inner problem to make the resulting decision differentiable, they differ on the way they update the parameters of the prediction model.
+DF-FTPL uses the FTPL approach [43], and DF-OGD leverages a variant of Online Gradient Descent [81].
+We establish sublinear convergence guarantees for both procedures, in the form of a static regret bound for the former and a dynamic regret bound for the latter.
+To our knowledge, these are the first provable guarantees for the online decision-focused learning problem.
   3. (iii)
-    Based on a margin assumption originating from the classification and bandit literature, we provide convergence guarantees in the form of dynamic regret bounds (defined in Section 2) for DF-OGD, when the decision space is either the simplex or a general convex polytope.
-  4. (iv)
-    Finally, we compare the empirical performance of DF-OGD to that of a classical prediction-focused approach in an online knapsack experiment inspired by mandi2024decision .
-In our simulation, our method outperforms the standard PFL approach in terms of downstream decision loss, in spite of degraded prediction loss.
+    Finally, we assess the performance of our algorithms on a knapsack experiment inspired by [54].
+Our simulations demonstrate that our approach outperforms the online version of two popular baselines, namely prediction-focused learning and Smart-Predict-then-Optimize, in both static and dynamic environments.
 Additional related work.
-Our work addresses the decision-focused learning problem, pioneered by elmachtoub2022smart .
-Several approaches have been proposed in the batch setting, among which differentiating the associated KKT conditions gould2016differentiating ; amos2017optnet ; donti2017task ; wilder2018melding ; mandi2020interior , smoothing via random perturbation berthet2020learning , building surrogate losses via duality elmachtoub2022smart or directional gradients huang2024decision and relying on pairwise ranking techniques mandi2022decision , see mandi2024decision ; sadana2025survey for additional references.
+Our work addresses the decision-focused learning problem, pioneered by [25].
+Several approaches have been proposed in the batch setting, among which differentiating the associated KKT conditions [33, 4, 23, 74, 52], smoothing via random perturbation [9], building surrogate losses via duality [25] or directional gradients [42] and relying on pairwise ranking techniques [53], see [54, 62] for additional references.
 However, the extension of decision-focused learning to the online setting remains unexplored.
-While several recent studies address online bi-level optimization shen2023online ; tarzanagh2024online ; lin2023non , none of them are applicable to decision-focused learning as they rely on restrictive smoothness assumptions, that are incompatible with the structure of decision-focused problems.
+While several recent studies address online bi-level optimization [65, 69, 49], none of them are applicable to decision-focused learning as they rely on restrictive smoothness assumptions, that are incompatible with the structure of decision-focused problems.
 In particular, the decision-focused objective is typically non-convex and features gradients that are either zero or undefined, owing to its underlying linear structure.
 This motivates the use of different online methods, specifically tailored to address these problems.
-On the one hand, lack of differentiability is usually tackled through zero-th order methods heliou2020gradient ; frezat2023gradient , sub-gradient duchi2011adaptive , proximal dixit2019online or smoothing approaches abernethy2014online .
-On the other hand, methods to address non-convexity in online learning often rely on the existence of a near optimal oracle kalai2005efficient ; agarwal2019learning ; suggala2020online ; xu2024online or additional smoothness conditions lesage2020second ; ghai2022non .
-All those results are derived for static regret zinkevich2003online ; hazan2023introductiononlineconvexoptimization which compares the learner efficiency to the best static strategy (see Appendix A).
-An alternative approach to obtain guarantees in the online non-convex setting is to weaken static regret to local regret hazan2017efficient ; aydore2019dynamic ; zhuang2020no ; hallak2021regret , which is the sum of the objective gradient norms evaluated in the iterates over time.
+On the one hand, lack of differentiability is usually tackled through zero-th order methods [39, 29], sub-gradient [24], proximal [22] or smoothing approaches [1].
+On the other hand, methods to address non-convexity in online learning often rely on the existence of a near optimal oracle [45, 3, 67, 75] or additional smoothness conditions [46, 32].
+All those results are derived for static regret [81] which compares the learned predictors to the best static strategy and is defined in Section˜2.
+A more challenging criterion, also described in Section˜2 is dynamic regret, introduced in [81] and later developed in [35, 12, 79, 78] among others, which takes into account the evolution of the environment.
+An alternative approach to obtain guarantees in the online non-convex setting is to weaken static regret to local regret [38, 6, 80, 36], which is the sum of the objective gradient norms evaluated in the iterates over time.
 Minimizing local regret corresponds to encouraging convergence toward stationary points.
 However as we shall see, this approach is not sensible in our framework, as objective gradients are either zero or undefined.
-In this work, we focus on dynamic regret, defined in Section 2.
-Originally introduced in zinkevich2003online and later developed in hall2013dynamical ; besbes2015non ; zhao2020dyn ; zhao2021improved among others, this regret takes into account the evolution of the environment and is well suited to our framework.
-We also provide an analysis for the more classic static regret notion based on techniques from suggala2020online , see Appendix B.
 Outline.
-In Section 2, we introduce the online decision-focused problem, our regret notion and our assumptions.
-In Section 3, we present the DF-OGD algorithm, before analyzing its dynamic regret performance in the simplex and convex polytope cases in Section 4.1 and Section 4.2 respectively.
-Finally, we present our experiment in Section 5.
+In Section˜2, we introduce the online decision-focused problem, our notions of regret and our assumptions.
+In Section˜3, we present the DF-FTPL and DF-OGD algorithms, before deriving bounds on the static regret of the former and the dynamic regret of the latter.
+Finally, we present our experiment in Section˜4.
 Notation.
-For a differentiable map $g:\mathbb{R}^{m}\rightarrow\mathcal{Y}$ such that $\mathcal{Y}\subseteq\mathbb{R}$, $\nabla g(x)$ denotes the gradient of $g$ at $x\in\mathbb{R}^{m}$ and $\nabla^{2}g(x)$ denotes the Hessian of $g$.
-In the case where $\mathcal{Y}\subseteq\mathbb{R}^{d}$, $\nabla g(x)$ denotes the Jacobian of $g$.
+For a differentiable map $\varphi:\mathbb{R}^{m}\rightarrow\mathcal{Y}$ such that $\mathcal{Y}\subseteq\mathbb{R}$, $\nabla\varphi(x)$ denotes the gradient of $g$ at $x\in\mathbb{R}^{m}$ and $\nabla^{2}\varphi(x)$ denotes the Hessian of $g$.
+In the case where $\mathcal{Y}\subseteq\mathbb{R}^{d}$, $\nabla\varphi(x)$ denotes the Jacobian of $\varphi$.
 For two vectors $(v,w)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $v\succcurlyeq w$ means that $v_{i}\geqslant w_{i}$ for any $i\in[d]$, and $\mathopen{}\mathclose{{\left\langle v,w}}\right\rangle=v^{\mathsf{\scriptscriptstyle T}}w$ refers to the standard Euclidian inner product.
@@
 $\min_{w\in\mathcal{W}}\ \langle\bar{g}_{t}(X_{t}),w\rangle\;,$ (1)
-where $X_{t}\in\mathcal{X}$ are covariates (possibly random), $\mathcal{W}=\mathrm{Conv}(v_{1},\ldots,v_{K})$ is a bounded convex polytope of $\mathbb{R}^{d}$ with non-empty interior and vertices $\{v_{i}\}_{i=1}^{K}$, and $\bar{g}_{t}:\mathcal{X}\to\mathcal{Z}$, $\mathcal{Z}\subset\mathbb{R}^{d}$ is a cost function, which for instance satisfies $\bar{g}_{t}(X_{t})=\mathbb{E}[Z_{t}\mid X_{t}]$ for some hidden state $Z_{t}\in\mathcal{Z}\subset\mathbb{R}^{d}$.
+where $X_{t}\in\mathcal{X}$ are covariates (possibly random) and $\mathcal{W}=\mathrm{Conv}(v_{1},\ldots,v_{K})$ is a bounded convex polytope of $\mathbb{R}^{d}$ with non-empty interior and vertices $\{v_{i}\}_{i=1}^{K}$.
+This feasible set appears naturally in many problems such as shortest-path [30], portfolio selection [47] or mixed strategy design in games [68].
+In (1), $\bar{g}_{t}:\mathcal{X}\to\mathbb{R}^{d}$, is a cost function, which satisfies for instance $\bar{g}_{t}(X_{t})=\mathbb{E}[Z_{t}\mid X_{t}]$ for some hidden state $Z_{t}\in\mathbb{R}^{d}$.
 Importantly, $X_{t}$ is revealed at the beginning of the round, but not $\bar{g}_{t}(X_{t})$ which is only available at the end of the round.
-While $\bar{g}_{t}(X_{t})$ is unknown to the decision-maker at the decision time, they have access to a family of models $g:\Theta\times\mathcal{X}\rightarrow\mathcal{Z}$, parameterized by $\Theta\subset\mathbb{R}^{m}$, to predict it.
-Then, the decision-maker seeks to jointly fit $\theta$ and solve the resulting decision problem.
-More precisely, the decision-maker considers a joint process $(\theta_{t},w_{t})_{t\in[T]}$, starting from $\theta_{1}\in\Theta$ and a history $\mathcal{H}_{0}=\emptyset$, and defined by the following recursion.
-At round $t\in[T]$, the decision-maker takes the best action given the collected history $\mathcal{H}_{t-1}$, the current estimate $\theta_{t}$ and a new feature $X_{t}$.
-Formally, they make the decision $w_{t}=w^{\star}_{t}(\theta_{t})$ where for any $\theta\in\Theta$,
-$w_{t}^{\star}(\theta)\in\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\langle g(\theta,X_{t}),w\rangle\;.$ (2)
+While $\bar{g}_{t}(X_{t})$ is unknown to the decision-maker at the decision time, they have access to a family of models $g:\Theta\times\mathcal{X}\rightarrow\mathbb{R}^{d}$, parameterized by $\Theta\subset\mathbb{R}^{m}$, to predict it.
+Then, the general form of the decision-making dynamics we consider can be described as follows: at each round $t\in[T]$, for a horizon $T\in\mathbb{N}$, given the current parameter $\theta_{t}$,
+  1. 1.
+    The decision-maker observes $X_{t}$ and compute its prediction as $g(\theta_{t},X_{t})$.
+  2. 2.
+    Then, they take an action minimizing the resulting predicted cost
+    $w_{t}=w^{\star}_{t}(\theta_{t})\in\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),w}}\right\rangle$ (2)
+  3. 3.
+    Finally, they observe $\bar{g}_{t}(X_{t})$ and update $\theta_{t+1}$ for the next round.
+Formally, the decision-maker considers a joint process $(\theta_{t},w_{t})_{t\in[T]}$, starting from $\theta_{1}\in\Theta$ and a history $\mathcal{H}_{0}=\emptyset$, and defined by the following recursion.
+At round $t\in[T]$, the decision-maker takes the best action given the collected history $\mathcal{H}_{t-1}$, the current estimate $\theta_{t}$ and a new feature $X_{t}$ as in (2).
 Then, they observe $\bar{g}_{t}(X_{t})$ and update their history $\mathcal{H}_{t}=\mathcal{H}_{t-1}\cup\{(X_{t},\bar{g}_{t}(X_{t}),\theta_{t},w^{\star}_{t}(\theta_{t}))\}$.
@@
 Online decision-focused learning.
-From an online perspective, prediction-focused learning consists in selecting at each $t\in[T]$ an algorithm $\mathsf{Alg}_{t}$ to estimate $\operatorname*{\arg\!\min}_{\theta\in\Theta}\mathrm{R}_{t}(\theta)$, where $\mathrm{R}_{t}$ is a risk function based on the historical observations $\mathcal{H}_{t-1}$, typically chosen as an empirical risk of the form $\mathrm{R}_{t}(\theta)=\sum_{i=1}^{t}\ell(\bar{g}_{i}(X_{i}),g(\theta,X_{i}))$, for some loss function $\ell$ (e.g.
-cross-entropy or squared error).
+From an online perspective, prediction-focused learning consists in selecting at each $t\in[T]$ an algorithm $\mathsf{Alg}_{t}$ to estimate $\operatorname*{\arg\!\min}_{\theta\in\Theta}\mathrm{R}_{t}(\theta)$, where $\mathrm{R}_{t}$ is a statistical risk based on the historical observations $\mathcal{H}_{t-1}$, typically chosen as the empirical risk $\mathrm{R}_{t}(\theta)=\sum_{i=1}^{t}\ell(\bar{g}_{i}(X_{i}),g(\theta,X_{i}))$, for some loss function $\ell$ (e.g., cross-entropy or squared error).
 Alternatively, we consider in this paper, the decision-focused learning approach, which selects parameters $\theta\in\Theta$ by directly minimizing the downstream objective instead of a risk function.
 Specifically, the decision-maker chooses $\mathsf{Alg}_{t}$ to solve $\operatorname*{\arg\!\min}_{\theta\in\Theta}\langle\bar{g}_{t}(X_{t}),w_{t}^{\star}(\theta)\rangle$ where $w_{t}^{\star}$ is defined in (2).
-From the previous formulation, we see that the decision-focused learning formulation corresponds to a bilevel optimization problem colson2007overview ; sinha2017review ; ji2021bilevel .
+From the previous formulation, we see that the decision-focused learning formulation corresponds to a bilevel optimization problem [19, 66, 44].
 We emphasize that no stationary assumptions are made on the process $\{X_{t},\bar{g}_{t}(X_{t})\}_{t\in[T]}$, allowing an adversarial selection of the features $X_{t}$ and cost functions $\bar{g}_{t}$.
-To fully accommodate this flexibility, we assess the optimality of $\{\theta_{t}\}_{t\in[T]}$ with a dynamic regret.
-It compares the learner’s performance against a sequence of best decisions that may vary over time.
-More precisely, let
+To fully accommodate this flexibility, we assess the optimality of $\{\theta_{t}\}_{t\in[T]}$ with two notions of regret.
+Let
 $f_{t}:\theta\in\Theta\mapsto\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}}\right\rangle\;,$ (3)
 denotes the loss incurred when taking an action based on the prediction parameter $\theta_{t}\in\Theta$.
-In line with zinkevich2003online ; haddouche2024optimisticallytemperedonlinelearning , we define the following dynamic regret:
-$\mathfrak{R}_{T}=\sum_{t\in[T]}F_{t}(\theta_{t})-\sum_{t\in[T]}\inf_{\theta\in\Theta}F_{t}(\theta)\;,\quad\text{where}\quad F_{t}:\theta\mapsto\mathbb{E}\mathopen{}\mathclose{{}\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}\right]\;.$ (4)
-We also consider a static regret and derive associated bounds in Appendix B. We make the following mild regularity assumptions for the rest of the analysis.
+Inspired from [81], we consider the notions of static and dynamic regret to measure the effectiveness of a learning strategy $\mathopen{}\mathclose{{\left\{\theta_{t}}}\right\}_{t\in[T]}$, which are respectively:
+$\mathfrak{R}^{s}_{T}=\sum_{t\in[T]}f_{t}(\theta_{t})-\inf_{\theta\in\Theta}\sum_{t\in[T]}\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)}}\right]\quad\text{and}\quad\mathfrak{R}^{d}_{T}=\sum_{t\in[T]}F_{t}(\theta_{t})-\sum_{t\in[T]}\inf_{\theta\in\Theta}F_{t}(\theta),$ (4)
+where $\quad F_{t}:\theta\mapsto\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}}\right]$.
+Note that our static regret compare to the best fixed strategy, with respect to the averaged losses.
+This notion makes sense here as we aim to control those regrets in expectation over the randomness of the process.
+In our dynamic regret, we consider the expectation of the loss given past information to prevent the dynamic comparator from overfitting the current realization of $X_{t}$.
+We make the following mild regularity assumptions for the rest of the analysis.
 H 1.
-(i) $\Theta$ is a compact and convex set with diameter $D_{\Theta}<\infty$.
-(ii) For any $\theta\in\Theta$ and $t\in[T]$, $\nabla_{\theta}g(\theta,X_{t})$ is continuously differentiable and $\mathopen{}\mathclose{{}\left\|\nabla_{\theta}g(\theta,X_{t})}\right\|_{\mathrm{op}}\leqslant G<\infty$ almost-surely.
+(i) $\Theta\subset\mathbb{R}^{m}$ is a compact and convex set with diameter $D_{\Theta}<\infty$.
+(ii) For any $\theta\in\Theta$ and $t\in[T]$, $\nabla_{\theta}g(\theta,X_{t})\in\mathbb{R}^{m\times d}$ is continuously differentiable and $\mathopen{}\mathclose{{\left\|\nabla_{\theta}g(\theta,X_{t})}}\right\|_{\mathrm{op}}\leqslant G<\infty$ almost-surely.
 (iii) For any $t\in[T]$, $\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|\leqslant D_{\mathcal{Z}}<\infty$ almost-surely.
-It is common in online learning to assume boundedness of the parameter space, model gradient and prediction space boyd2004convex ; bishop2006pattern .
+It is common in online learning to assume boundedness of the parameter space, model gradient and prediction space [14, 13].
 Note that in the well-specified setting, i.e., $\bar{g}_{t}=g(\bar{\theta}_{t},\cdot)$ for some $\bar{\theta}_{t}$, these two former conditions automatically imply the latter.
-We emphasize however that we do not assume this realizability condition, in contrast to most of the literature on decision-focused learning bennouna2024addressing .
+We emphasize however that we do not assume this realizability condition, in contrast to most of the literature on decision-focused learning [8].
 Without further assumptions on the sequence of costs $\bar{g}_{t}$ and models $g$, the problem can still be made arbitrarily hard.
-Therefore, we make an assumption coming from the classification mammen1999smooth ; tsybakov2004optimal and bandit literature zeevi2009woodroofe ; perchet2013multi about the margins of the cost function.
+Therefore, we make an assumption coming from the classification [51, 70] and bandit literature [76, 59] about the margins of the cost function.
 Recall that we denote by $\{v_{i}\}_{i=1}^{K}$ the vertices of $\mathcal{W}$.
-We define for any $x\in\mathcal{X}$, $u_{i}(\theta,x)=\mathopen{}\mathclose{{}\left\langle g(\theta,x),v_{i}}\right\rangle$ for $i\in[d]$, and assume that
+We define for any $x\in\mathcal{X}$, $u_{i}(\theta,x)=\mathopen{}\mathclose{{\left\langle g(\theta,x),v_{i}}}\right\rangle$ for $i\in[d]$.
 H 2.
-There exists $C_{0}>0$ and $\beta\in\mathopen{}\mathclose{{}\left[0,1}\right]$, such that almost surely for any $t\in[T]$, $\theta\in\Theta$ and $\varepsilon\in\mathopen{}\mathclose{{}\left[0,1}\right]$,
+There exist $C_{0}>0$ and $\beta\in\mathopen{}\mathclose{{\left[0,1}}\right]$, such that almost surely for any $t\in[T]$, $\theta\in\Theta$ and $\varepsilon\in\mathopen{}\mathclose{{\left[0,1}}\right]$,
 $\mathbb{P}\mathopen{}\mathclose{{\left(\inf_{j\neq I_{t}(\theta)}\{u_{j}(\theta,X_{t})-u_{I_{t}(\theta)}(\theta,X_{t})\}\geqslant\varepsilon\,\mid\mathcal{H}_{t-1}}}\right)\geqslant 1-C_{0}\varepsilon^{\beta}\;,$
 where $I_{t}(\theta)\in\operatorname*{\arg\!\min}_{i\in[K]}u_{i}(\theta,X_{t})$.
-H 2 controls how difficult it is to solve problem (2) since it determines the objective gap between the optimal vertex $v_{I_{t}(\theta)}\in\mathcal{W}$ and the other vertices.
+H˜2 controls how difficult it is to solve problem (2) since it determines the objective gap between the optimal vertex $v_{I_{t}(\theta)}\in\mathcal{W}$ and the other vertices.
 In other words, it quantifies how identifiable is the optimal vertex, $\varepsilon=0$ meaning it is not distinguishable from the others.
-Remark that H 2 is only related to the distribution of the features $\{X_{t}\}_{t\in[T]}$.
-If it is satisfied for $\beta\geqslant 1$, then it is automatically satisfied for $\beta=1$ and therefore we do not loose any generality restraining $\beta$ to $\mathopen{}\mathclose{{}\left[0,1}\right]$.
-This assumption is critical in our analysis, as it allows to bound the expected distance between the actual optimal decision $w^{\star}_{t}(\theta)$ and the regularized approximation $\tilde{w}_{t}(\theta)$ introduced in (5) below.
-3 An online algorithm for decision-focused learning.
-A natural approach to minimize the regret (4) would be to use an Online Gradient Descent (OGD) scheme zinkevich2003online for the algorithms $\{\mathsf{Alg}_{t}\}_{t\in[T]}$, which is standard in online learning.
-However, since in (2), $w^{\star}_{t}(\theta)$ minimizes a linear function over the convex polytope $\mathcal{W}$, $\theta\mapsto w_{t}^{\star}(\theta)$ ranges in the finite set of vertices $\{v_{1},\ldots,v_{K}\}$, thus $\nabla w^{\star}_{t}(\theta)$ is either zero or undefined.
-To address this issue, following wilder2018melding , we propose to add a regularizer $\mathcal{R}$ to the objective function in (2). Accordingly, we define for any $\theta\in\Theta$,
+If it is satisfied for $\beta\geqslant 1$, then it is automatically satisfied for $\beta=1$ thus we do not loose any generality restraining $\beta$ to $\mathopen{}\mathclose{{\left[0,1}}\right]$.
+This assumption is critical in our analysis, as it allows bounding the expected distance between the actual optimal decision $w^{\star}_{t}(\theta)$ and the regularized approximation $\tilde{w}_{t}(\theta)$ introduced in (5) below.
+We expect H˜2 to hold in a wide variety of classical statistical settings.
+For example, in Appendix˜A we show that it is satisfied when $g:(\theta,X)\mapsto X\theta$ where $X$ has i.i.d standard Gaussian columns.
+3 Online algorithms for decision-focused learning .
+On the need of regularization to get differentiation.
+From an online learning perspective, a natural approach to minimize the two regrets defined in (4) would be to update $\theta_{t+1}$ at each round $t$ using a variant of either the Follow-The-Leader algorithm [FTL, 45] or Online Gradient Descent [OGD, 81], applied to the objective function $f_{t}$ defined in (3), thus specifying the choice of $\mathsf{Alg}_{t}$.
+However, these algorithms are designed for single-level optimization problems and require access to (sub-)gradients of the objective.
+In our setting, this requirement is problematic because the function $f_{t}$ does not admit an informative gradient: by construction, computing the gradient of $f_{t}$ involves differentiating the mapping $\theta\mapsto w_{t}^{\star}(\theta)$, which is generally not (sub-)differentiable.
+Indeed, $w^{\star}_{t}$ minimizes a linear function over the convex polytope $\mathcal{W}$, thus $\theta\mapsto w_{t}^{\star}(\theta)$ ranges in the finite set of vertices $\{v_{1},\ldots,v_{K}\}$.
+To address this issue, following [74], we propose to add a regularizer $\mathcal{R}$ to the objective function in (2). Accordingly, we define for any $\theta\in\Theta$,
 $\tilde{w}_{t}(\theta)\in\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\{\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),w}}\right\rangle+\alpha_{t}\mathcal{R}(w)\}\;,$ (5)
 which is a regularized approximation of $w^{\star}_{t}$.
-When this surrogate is continuously differentiable—which holds for our choices of $\mathcal{R}$ in Section 4.1 and Section 4.2—we can use $\nabla\tilde{w}_{t}(\theta)$ in our algorithmic routine in place of $\nabla w^{\star}_{t}(\theta)$.
+When this surrogate is continuously differentiable—which holds for our choices of $\mathcal{R}$ in Section˜3—we can use $\nabla\tilde{w}_{t}(\theta)$ in our algorithmic routine.
 This approach amounts to minimizing the surrogate
 $\tilde{f}_{t}:\theta\mapsto\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle\;,$ (6)
-instead of $f_{t}$, which admits gradients of the form
-$\nabla\tilde{f}_{t}(\theta)=\nabla\tilde{w}_{t}(\theta)^{\mathsf{\scriptscriptstyle T}}\bar{g}_{t}(X_{t})\;.$
-Note that $\tilde{f}_{t}$ is not necessarily convex (even when $\theta\mapsto g(\theta,X)$ is linear) because it involves the parametrized minimizer $\tilde{w}_{t}(\theta)$.
-This prevents us from directly using known online convex optimization algorithms hazan2023introductiononlineconvexoptimization .
-However, a recent line of research kalai2005efficient ; agarwal2019learning ; suggala2020online ; xu2024online has developed online algorithms in the non-convex case, provided the losses are Lipschitz-continuous—which is the case with our choices of regularizers in Section 4.1 and Section 4.2 (see Lemmas 2 and 7).
+instead of $f_{t}$, which admits gradients of the form $\nabla\tilde{f}_{t}(\theta)=\nabla\tilde{w}_{t}(\theta)^{\mathsf{\scriptscriptstyle T}}\bar{g}_{t}(X_{t})$.
+Note that there is a natural trade-off in the choice of the regularization parameter $\alpha_{t}$.
+On the one hand, choosing a large $\alpha_{t}$ makes the function $\tilde{w}_{t}(\theta)$ smoother.
+On the other hand, the larger $\alpha_{t}$ is, the more $\tilde{w}_{t}$ deviates from the true function $w^{\star}_{t}$ that we aim to approximate.
+As we will see later, it is possible to balance these two extremes by carefully tuning the parameter $\alpha_{t}$.
+On the choice of regularization on a general polytope.
+In what follows, we write $\mathcal{W}$ as the intersection of $n$ half-spaces, that is $\mathcal{W}=\{w\in\mathbb{R}^{d}:\>A^{\mathsf{\scriptscriptstyle T}}w-b\preccurlyeq 0\}$ where $n>0$ is the number of faces, $A\in\mathbb{R}^{d\times n}$ and $b\in\mathbb{R}^{n}$.
+We assume that $\mathcal{W}$ is not degenerated, i.e.
+$AA^{\mathsf{\scriptscriptstyle T}}\in\mathbb{R}^{d\times d}$ is full-rank.
+It remains to determine what regularizer $\mathcal{R}$ to choose in (6).
+We recall that our aim is to obtain a $\tilde{w}_{t}(\theta)$ in (5) which is differentiable for any $\theta\in\Theta$.
+A possible strategy is to choose $\mathcal{R}$ so $\tilde{w}_{t}$ remains in the strict interior of $\mathcal{W}$.
+In this case, $\tilde{w}_{t}(\theta)$ is differentiable in a neighborhood of any $\theta\in\Theta$ and admits a close-form Jacobian $\nabla\tilde{w}_{t}(\theta)$ by the implicit function theorem.
+A natural choice to force $\tilde{w}_{t}$ to remains in the interior of $\mathcal{W}$ is the corresponding log-barrier function:
+$\mathcal{R}:w\mapsto-\sum_{i=1}^{n}\ln(b_{i}-A_{i}^{\mathsf{\scriptscriptstyle T}}w)\;,$ (7)
+where $A_{i}\in\mathbb{R}^{d}$ is the $i$-th column of $A$. With this choice of regularization, we show in Lemma˜4 that $\nabla\tilde{w}_{t}$ has an explicit formulation, allowing its use in practice.
+Remark 1.
+We remark that in the special case where $\mathcal{W}=\{w\in\mathbb{R}^{d}:w\succcurlyeq 0\;\text{,}\;{\bf 1}^{\mathsf{\scriptscriptstyle T}}w=1\}$ is the simplex of $\mathbb{R}^{d}$, an alternative choice for $\mathcal{R}$ in (5) is the negative entropy $\mathcal{R}_{0}:w\mapsto\sum_{i\in[d]}w_{i}\ln(w_{i})$.
+In this case, $\tilde{w}_{t}(\theta)$ in (5) reduces to the softmax mapping, which is differentiable and admits a close-form Jacobian.
+In Appendix˜C, we theoretically study the performances of our algorithms in this special case.
+Approximate oracles to handle non-convexity.
+The bi-level structure of the problem yields unexpected properties.
+In particular, even when the regularizer $\mathcal{R}$ is strongly-convex in (26)—which the case with both log-barriers and negative entropy—and the model $g$ is simple (such as $g(\theta,X)=X\theta$), $\tilde{f}_{t}:\theta\mapsto\mathopen{}\mathclose{{\left\langle\tilde{w}_{t}(\theta),\bar{g}_{t}(X_{t})}}\right\rangle$ may not be convex, but is lipschitz (see Lemmas˜5 and 1).
+This prevents us from directly using known online convex optimization algorithms [37].
+However, a recent line of research [3, 67, 75] has developed online algorithms in the non-convex case, provided the losses are Lipschitz-continuous—which is the case with log-barriers or negative entropy (see Lemmas˜1 and 5).
 These studies combine near-optimal oracles with perturbation techniques to establish sublinear bounds on the expected regret.
-In line with this literature, we also assume to have access to a near-optimal oracle.
-In the following assumption, we define for any $t\in[T]$, $\tilde{F}_{t}:\theta\mapsto\mathbb{E}[\tilde{f}_{t}(\theta)\,|\,\mathcal{H}_{t-1}]$.
-H 3.
-There exists $\xi>0$ such that for any $t\in[T]$, upon observing $\mathcal{H}_{t}$, we can compute $\vartheta_{t}\in\Theta$ satisfying almost surely:
-$\tilde{F}_{t}(\vartheta_{t})\leqslant\inf_{\theta\in\Theta}\tilde{F}_{t}(\theta)+\xi\;.$ (7)
-One way to design such oracles is to apply Stochastic Gradient Descent amari1993backpropagation to the differentiable function $\tilde{f}_{t}$.
-In this case, $\xi\geqslant 0$ captures both the error made by the offline minimization method generating $\vartheta_{t}$, and the gap between the local optimum to which it converges and the global optimum.
-In particular, $\xi$ vanishes for benign loss landscapes.
-Finally, note that in the case where $\tilde{w}_{t}(\theta)$ can only be approximated, then $\xi$ can also encompass the resulting error.
-Leveraging the regularization and perturbation approaches previously mentioned, we now present our method, referred to as DF-OGD, in Algorithm 1.
+In line with this literature, we also assume to have access to an approximate offline optimization oracle, a notion that appeared in a slightly modified form in Suggala and Netrapalli [67].
+Definition 1.
+An $\xi$-approximate offline optimization oracle (or approximate oracle), adapted to a class $\mathcal{C}$ is a mapping $\mathbf{O}_{\xi}$ taking a function $f\in\mathcal{C}$ and outputting $\vartheta=\mathbf{O}_{\xi}(f)\in\Theta$ satisfying:
+$f(\vartheta)\leqslant\inf_{\theta\in\Theta}f(\theta)+\xi\;.$ (8)
+The notion of an approximate oracle formalizes that, in non-convex settings, we cannot rely on subroutines that provably attain a global minimizer of $\tilde{f}_{t}$ at time $t$ (as offline gradient descent would in convex problems).
+Instead, we must settle for local minimizers, whose quality is characterized by a parameter $\xi$—which vanishes in favorable loss landscapes.
+As a concrete example, consider the stochastic gradient descent (SGD) algorithm as $\mathbf{O}_{\xi}$.
+A large body of work has shown that, even when $\Theta$ is a class of deep neural networks, SGD can converge to local minimizers [31, 55, 58, 20], justifying that $\mathbf{O}_{\xi}$ is indeed an approximate oracle.
+A more detailed discussion is provided in Appendix˜B.
+We conclude this remark by mentioning that it is possible to conduct a non-convex analysis without an approximate oracle by considering the weaker notion of local regret defined as $\sum_{t\in[T]}\mathopen{}\mathclose{{\left\|\nabla F_{t}(\theta_{t})}}\right\|$; see [38, 6, 80, 36].
+However, such a definition is not meaningful in the DFL framework, where gradients are zero or undefined due to the structure of the problem.
+The combined use of approximate oracles and regularization allows us to derive original online algorithms.
+In Section˜3.1, we focus on a variant of the FTL algorithm, enjoying a static regret bound, while in Section˜3.2 we propose a version of OGD enjoying a dynamic regret bound, both results holding in the non-convex case.
+3.1 Decision-Focused Follow The Perturbed Leader
+Our first algorithm is displayed in Algorithm˜1, where $\mathrm{Exp}(\eta)$ refers to an exponential distribution with parameter $\eta>0$.
+It is inspired from the classic Follow-the-Leader approach [45], which consists in making a decision minimizing the sum of objective functions observed so far.
+When losses are non-convex, it is common to inject random noise at each step to regularize the total cost function, resulting in an approach known as Follow-the-Perturbed-Leader [43, 67].
+We employ this strategy as $\mathsf{Alg}_{t}$ to update $\theta_{t+1}$ in Algorithm˜1.
+Note that the oracle in line 6 of Algorithm 1 crucially works with the regularized losses $\tilde{f}_{1},\ldots,\tilde{f}_{t}$.
+This allows to use in practice gradient-based methods to obtain an approximate minimizer as per our discussion in Section˜3, since these surrogate losses are differentiable and Lipschitz.
+Algorithm 1 Decision-Focused Follow The Perturbed Leader DF-FTPL
+1:Input: horizon $T>0$, initialization $\theta_{1}\in\Theta$, $\xi$-approximate oracle $\mathbf{O}_{\xi}$ and history $\mathcal{H}_{0}=\emptyset$.
+2:for each $t\in\{1,\ldots,T\}$ do
+3:  Observe $X_{t}\in\mathcal{X}$ and play $w^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),w}}\right\rangle\;.$
+4:  Observe $\bar{g}_{t}(X_{t})\in\mathcal{Z}$ and update the history $\mathcal{H}_{t}$.
+5:  Draw $\sigma_{t}\in\mathbb{R}^{d}$ such that for all $j\in[d]$, the $j$-th component $\sigma_{j,t}\sim\mathrm{Exp}(\eta)$.
+6:  Update
+$\theta_{t+1}=\mathbf{O}_{\xi}\mathopen{}\mathclose{{\left(\sum_{i=1}^{t}\tilde{f}_{i}-\mathopen{}\mathclose{{\left\langle\sigma_{t},\cdot}}\right\rangle}}\right)$
+7:end for
+Note that Algorithm˜1 requires determining $w^{\star}$, which can be efficiently achieved by classic numerical solvers in reasonable settings.
+However, we show in Appendix˜F that only being able to determine $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ such that $\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),\underline{w}_{t}(\theta_{t})}}\right\rangle-\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),w^{\star}_{t}(\theta_{t})}}\right\rangle\leqslant\kappa$ for some $\kappa>0$ only shifts the regret bounds of the next section by $\kappa$.
+We provide a theoretical guarantee on the convergence of Algorithm˜1 in the form of the following static regret bound.
+Theorem 1.
+Assume H˜1, H˜2 and having access to an $\xi$-approximate oracle $\mathbf{O}_{\xi}$ adapted to $\{\sum_{i=1}^{t}\tilde{f}_{i}-\mathopen{}\mathclose{{\left\langle\sigma_{t},\cdot}}\right\rangle\}_{t\in[T]}$.
+Let $\{\theta_{t}\}_{t\in[T]}$ be the output of DF-FTPL (Algorithm˜1) instantiated with learning step $\eta>0$ and regularization coefficients $\alpha_{t}=\alpha>0$ for any $t$.
+Then:
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi+\alpha n}}\right)\;,$
+where $\mathbb{E}$ denotes the expectation on both data and the intrinsic randomness of DF-FTPL and $\tilde{\mathcal{O}}$ contains polynomial dependency in $\ln(1/\alpha),\ln(\ln(d))$.
+Furthermore, taking $\eta\propto m^{1/4}T^{-3/4}n^{-1/2}$ and $\alpha\propto m^{3/4}n^{1/2}T^{-1/4}$ yields:
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(m^{3/4}\sqrt{n}T^{-1/4}+\xi}}\right).$
+The proof of Theorem˜1 can be found in Section˜G.2.
+In particular, Theorem˜1 shows that DF-FTPL enjoys an average regret bound decaying in $T^{-1/4}$ as long as $\xi=\mathcal{O}(T^{-1/4})$.
+While it features a polynomial dependency $m^{3/4}$ on the dimension of $\Theta$, it only depends on the dimension of the decision space $\mathcal{W}$ through a $\ln\ln(d)$ term.
+This makes Algorithm˜1 a particularly competitive approach when the decision space is of high dimension.
+It is informative to compare our guarantee to existing bounds in the literature.
+[67], who also study non-convex online learning, achieves a rate of $\mathcal{O}(T^{-1/2})$ as long as their offline oracle satisfies $\xi=\mathcal{O}(T^{-1/2})$, at the cost of a degraded $m^{3/2}$ dependency on the dimension.
+Their faster rate in $T$ comes from the fact that they tackle a simpler, single-level problem as compared to our bi-level, non differentiable setting.
+Indeed, we need to regularize the inner problem to overcome non-differentiability as discussed in Section˜3.
+This introduces an additional trade-off on the regularization strength $\alpha$ on top of the usual trade-off in the learning rate $\eta$, which is reflected in our rate.
+3.2 Decision-Focused Online Gradient Descent
+While DF-FTPL (Algorithm˜1) benefits from a converging static regret bound, the techniques involved in [67] are not enough to reach dynamic regret guarantees, which is particularly relevant in highly non-stationary environments where the optimal decisions may change significantly from one round to another.
+To this end, we go beyond the Follow-the-Leader approach and propose an original algorithm based on the celebrated Online Gradient Descent [81].
+This procedure, which we call DF-OGD, is presented in Algorithm˜2.
+In words, at each time $t\in[T]$, Algorithm˜2 first makes a decision $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ based on the current parameters $\theta_{t}\in\Theta$.
+Then, it observes $\bar{g}_{t}(X_{t})\in\mathcal{Z}$, determines the surrogate objective $\tilde{f}_{t}$ and computes one of its near minimizer via the offline oracle from Definition˜1.
+Next, to handle non-convexity, it determines the gradient $\nabla\tilde{w}_{t}(u_{t})$, evaluated in a point $u_{t}$ drawn uniformly at random in $\theta_{t},\vartheta_{t}]$.
+From there, it updates $\theta_{t+1}$ through the classical gradient step scheme of OGD.
+[…]
+Moreover, note that Algorithm˜1 is instanciated with a single pair of parameters $(\alpha,\eta)$ whereas Algorithm˜2 uses on a sequence $(\alpha_{t},\eta_{t})_{t\in[T]}$.
+This additional flexibility is crucial for the algorithm to adapt to the variation of the problem so as to maintain a low dynamic regret.
 Algorithm 2 Decision-Focused Online Gradient Descent (DF-OGD)
@@
 4:  Observe $\bar{g}_{t}(X_{t})\in\mathcal{Z}$ and update the history $\mathcal{H}_{t}$.
-5:     Observe $\vartheta_{t}\in\Theta$ satisfying (7).
-6:     Draw $\delta_{t}\sim\text{Unif}([0,1])$, compute $u_{t}=\vartheta_{t}+\delta_{t}(\theta_{t}-\vartheta_{t})$ and ~$\tilde{\nabla}_{t}(u_{t})=\nabla\tilde{w}_{t}(u_{t})^{\mathsf{\scriptscriptstyle T}}\bar{g}_{t}(X_{t})\;.$
+5:  Get $\vartheta_{t}=\mathbf{O}_{\xi}(\tilde{f}_{t})$.
+6:  Draw $\delta_{t}\sim\text{Unif}([0,1])$, compute $u_{t}=\vartheta_{t}+\delta_{t}(\theta_{t}-\vartheta_{t})$ and $\tilde{\nabla}_{t}(u_{t})=\nabla\tilde{w}_{t}(u_{t})^{\mathsf{\scriptscriptstyle T}}\bar{g}_{t}(X_{t})\;.$
 7:  Update $\theta_{t+1}=\Pi_{\Theta}(\theta_{t}-\eta_{t}\tilde{\nabla}_{t}(u_{t}))\;.$
 8:end for
-At each time $t\in[T]$, Algorithm 1 first makes a decision $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ based on the current parameters $\theta_{t}\in\Theta$.
-Then, it observes $\bar{g}_{t}(X_{t})\in\mathcal{Z}$, determines the surrogate objective $\tilde{f}_{t}$ and computes the oracle $\vartheta_{t}$ from H 3.
-Next, it determines the gradient $\nabla\tilde{w}_{t}(u_{t})$, evaluated in a point $u_{t}$ drawn uniformly at random in $[\theta_{t},\vartheta_{t}]$.
-From there, it updates $\theta_{t+1}$ through the classical gradient step scheme of OGD, with the noticeable difference that the gradient is evaluated in $u_{t}$ instead of $\theta_{t}$.
-Algorithm 1 requires determining $w^{\star}$, which can be efficiently achieved by classic numerical solvers in reasonable settings.
-However, we show in Appendix E that only being able to determine $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ such that $\mathopen{}\mathclose{{}\left\langle g(\theta_{t},X_{t}),\underline{w}_{t}(\theta_{t})}\right\rangle-\mathopen{}\mathclose{{}\left\langle g(\theta_{t},X_{t}),w^{\star}_{t}(\theta_{t})}\right\rangle\leqslant\kappa$ for some $\kappa>0$ only shifts the regret bounds of the next section by $\kappa$.
-Note that while DF-OGD relies on OGD, it is possible to use other algorithmic approaches such as Follow the Perturbed Leader (FTPL) suggala2020online , see DF-FTPL in Appendix B for which we provide a static regret analysis.
-4 A theoretical study of convergence
-We now provide convergence guarantees for Algorithm 1 in the form of expected dynamic regret bounds in two settings.
-First, we deal with the case where $\mathcal{W}$ is the simplex.
-Second, we consider the general setting where $\mathcal{W}$ is any bounded convex polytope with non-empty interior.
-4.1 Online decision-focused learning on the simplex
-We first focus on the case where $\mathcal{W}$ is the simplex of $\mathbb{R}^{d}$, that is
-[…]
-This setting encompasses various decision-making scenarios such as portfolio selection li2014online or mixed strategy design in multiplayer games syrgkanis2015fast .
-In this case, we choose the negative entropy
+We provide a convergence guarantee for Algorithm˜2 with the following dynamic regret bound.
+Theorem 2.
+Assume H˜1, H˜2, access to a $\xi$-approximate oracle adapted to $\{\tilde{f}_{t}\}_{t\in[T]}$.
+Let $\{\theta_{t}\}_{t\in[T]}$ be the output of DF-OGD (Algorithm˜2) instantiated with the non-increasing sequence $\mathopen{}\mathclose{{\left(\eta_{t}}}\right)_{t\in[T]}$ and regularization coefficients $(\alpha_{t})_{t\in[T]}$.
+Then:
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\mathbb{E}\mathopen{}\mathclose{{\left[\frac{1+P_{T}}{T\eta_{T}}+\frac{1}{T}\sum_{t\in[T]}\frac{\eta_{t}}{\alpha_{t}^{2}}+n\alpha_{t}}}\right]+\xi}}\right)\;.$
+where $P_{T}=\sum_{t=1}^{T-1}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|$, $\mathbb{E}$ denotes the expectation on both data and the intrinsic randomness of DF-OGD and $\tilde{\mathcal{O}}$ contains polynomial dependency in $\ln(1/\alpha),\ln(\ln(d))$.
+Furthermore, assume $(t^{-1}(1+P_{t}))_{t\geqslant 1}$ is non-increasing almost surely with $P_{t}=\sum_{s=1}^{t}\mathopen{}\mathclose{{\left\|\vartheta_{s+1}-\vartheta_{s}}}\right\|$.
+Then, using $\alpha_{t}\propto n^{-1/2}t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto n^{-1/2}t^{-3/4}(1+P_{t})^{3/4}$ for any $t\in[T]$ leads to :
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\mathbb{E}\mathopen{}\mathclose{{\left[\sqrt{n}(1+P_{T})^{1/4}T^{-1/4}}}\right]+\xi}}\right)\;.$
+The complete statement of the main bound, including constants is provided in Section˜G.3.
+Our dynamic regret bound naturally depends on $P_{T}$, which captures the problem’s variability by measuring the cumulative distance between approximate minimizers over time.
+When $\xi=\mathcal{O}((1+P_{T})^{1/4}T^{-1/4})$, the average dynamic regret decreases at the rate $\mathcal{O}((1+P_{T})^{1/4}T^{-1/4})$.
+Notably, the bound is independent of the dimension of $\Theta$, and depends only mildly on the dimension of $\mathcal{W}$ through a $\ln\ln(d)$ factor.
+This constitutes a key strength of the optimistic strategy underlying DF-OGD, making it particularly well-suited for high-dimensional settings.
+It is instructive to compare our guarantee with those established in recent studies on dynamic regret.
+For instance, [77] derive a $\mathcal{O}(\sqrt{T(P_{T}+1)})$ bound.
+However, their setting is not directly comparable to ours, as they consider a simpler single-level problem with differentiable and convex objectives.
+In contrast, our framework involves non-convex, non-differentiable losses due to the bi-level nature of decision-focused learning.
+More recently, [41] obtained a $\mathcal{O}((1+P^{\infty}_{T})^{1/3}T^{-1/3})$ bound, where $P^{\infty}_{T}:=\sum_{t=1}^{T}\|f_{t+1}-f_{t}\|_{\infty}$.
+Yet, this rate is achieved under substantially more favorable conditions: a single-level problem with losses that are strongly convex or Lipschitz, an additional assumption of “quasi-stationary” and a more challenging path involving the full landscapes of $f_{t}$s.
+By comparison, our losses are neither Lipschitz (indeed, they may even be discontinuous due to the linearity of the lower-level problem) nor required to be stationary over time—the only restriction being H˜2 to hold.
+We view the ability of our algorithm to achieve efficient convergence despite these demanding conditions as a core contribution of our work.
+4 Experiments
+In this section, we compare the performances of our algorithms DF-FTPL and DF-OGD to two important benchmarks, namely prediction-focused learning and SPO [25].
+Setting.
+Our experimental setup is inspired by the knapsack example from Mandi et al.
+[54].
+More precisely, we consider a decision maker who must pick at each $t\in[T]$ an object $v_{t}\in\mathcal{V}$ among $K$ items denoted $\mathcal{V}=\{1,2,\ldots,K\}$ with respective costs $\bar{g}_{t}(X)=(\bar{g}_{t,1}(X),\ldots,\bar{g}_{t,K}(X))\in[0,1]^{K}$ depending on some covariates $X\in\mathcal{X}$.
+At the beginning of each period, the decision-maker only observe covariates $X_{t}\in\mathcal{X}\subset\mathbb{R}^{p}$, and have at their disposal a parametric model $g:\Theta\times\mathcal{X}\rightarrow[0,1]^{K}$ to predict $\bar{g}_{t}$.
+Given their current parameter $\theta_{t}\in\Theta$, they predict item costs as $g(\theta_{t},X_{t})=(g_{1}(\theta_{t},X_{t}),\ldots,g_{K}(\theta_{t},X_{t}))$ and pick an item $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{i\in[K]}g_{i}(\theta_{t},X_{t})$.
+This setting is depicted in Figure˜2 for $K=2$ in Appendix˜E.
+After having made their decision, the decision-maker observes the true item costs $\bar{g}_{t}(X_{t})$, and update $\theta_{t+1}$ for the next round based on this feedback.
+Note that this setting can directly be mapped in the simplex example of Remark˜1, since $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{w\in\mathcal{W}_{0}}\mathopen{}\mathclose{{\left\langle w,g(\theta_{t},X_{t})}}\right\rangle$, $\mathcal{W}_{0}$ being the simplex of $\mathbb{R}^{K}$.
+For this reason, we instantiate our algorithms DF-OGD and DF-FTPL with the negative entropy regularizer, following Remark˜1.
+Synthetic data.
+We instantiate the previous problem with the following synthetic data.
+For any $t\in[T]$, we draw $X_{t}\in\mathbb{R}^{K\times p}$ with correlated rows, and generate a cost vector $c_{t}(X_{t})\in\mathbb{R}^{k}$ as :
+$c_{t}(X_{t})=A\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon_{t}\;,$ (9)
+where $A>0$, $\varepsilon_{t}\sim\mathrm{N}(0,I_{K})$ is a Gaussian noise and $\theta^{\star}_{t}\in\mathbb{R}^{p}$ is a parameter which satisfies $\theta^{\star}_{t}=\nicefrac{{1}}{{2}}\,\theta^{\star}+\nicefrac{{1}}{{2}}\,\zeta_{t}\,,\quad\text{where}\quad\zeta_{t}\sim\mathrm{N}(0,I_{p})$ for some $\theta^{\star}\in\mathbb{R}^{p}$.
+This is a challenging data generating process, since $\bar{g}_{t}(X_{t})=\sin^{4}((2X_{t}^{\mathsf{\scriptscriptstyle T}}\theta^{\star}_{t})^{-1})$ is non-stationary and highly non-linear, and features are correlated.
+Equation˜9 is discussed more in detail in Appendix˜E.
+To predict $c_{t}$ from $X_{t}$, we assume that the decision-maker has access to a class of linear predictors of the form $g:(\theta,X)\mapsto X\theta$.
+Benchmarks.
+In this setting, we compare the performances of DF-FTPL (Algorithm˜1) and DF-OGD (Algorithm˜2) to two benchmarks.
+First, we implement Prediction-Focused Online Gradient Descent (PF-OGD).
+This strategy consists in training in an online manner the model $g$ at each timestep so it minimizes the statistical loss $\ell^{\mathrm{mse}}:(v,X\theta)\mapsto\mathopen{}\mathclose{{\left\|v-X\theta}}\right\|^{2}$, irrespective of the downstream decision problem.
+Then, decisions are greedily made based on the predictions of the model.
+This approach is formally described in Algorithm˜3 in Appendix˜E.
+Second, we compare our algorithms to an online version of the Smart Predict-then-Optimize (online SPO, see Algorithm˜4) approach from [25].
+This algorithm introduces a differentiable and convex loss which upper-bounds the actual decision-focused loss.
+Given its effectiveness and versatility, it is considered as a very important benchmark in the literature.
+Results.
+In Figure˜1, we plot on the left-hand panel the average cumulated cost $t\mapsto t^{-1}\sum_{s=1}^{t}\bar{g}_{s,v_{s}}(X_{s})$ incurred by DF-OGD, DF-FTPL, PF-OGD and online SPO over 10 runs of $T=5000$ timesteps, as well as the associated 95% confidence intervals.
+On the right-hand panel, we plot the average Mean Square Error (MSE) resulting from the sequence of prediction parameters $(\theta_{t})_{t\in[T]}$.
+It appears that both DF-FTPL and DF-OGD outperform PF-OGD and online SPO from a decision point of view, while incurring a higher MSE.
+This is in line with decision-focused approach, which cares about decision cost rather than statistical loss.
+This experiment shows that (i) DFL outperforms PFL when models are clearly mispecified and (ii) our algorithms also outperforms the celebrated online SPO.
+Figure 1: Average cumulated cost (left) and prediction errors (right) of DF-OGD, PF-OGD, DF-FTPL and online SPO.
+We also mention the presence of additional numerical experiments in Appendix˜E, to study how model misspecification affects the relative performances of DFL and PFL.
+5 Conclusion.
+Decision-focused learning offers a promising way to integrate prediction into decision making.
+We extend its analysis from the batch to the online setting, enabling non-stationary data and varying objectives.
+Our algorithms DF-FTPL and DF-OGD comes with provable guarantees, and empirically outperform prediction-focused learning and online SPO in our experiment.
+We believe that this work can be extended in several ways.
+First, we believe that alternative smoothing techniques could be successfully employed, such as Moreau-Yosida transform and proximal operators.
+Second, it would be valuable to investigate less adversarial environments, such as those involving i.i.d.
+data, to explore whether stronger theoretical guarantees could be obtained and whether novel algorithmic designs might emerge.
+Finally, implementing our method in more ambitious experimental settings would be of great interest from a practical perspective.
+Appendix A H˜2 is satisfied in the Gaussian linear case.
+In this section, we prove that H˜2 holds true when $\mathcal{W}$ is the simplex of $\mathbb{R}^{d}$, $g$ is a linear model and $X$ has columns distributed according to a Gaussian distribution.
+In what follows, we denote by $X_{j}\in\mathbb{R}^{d}$ the $j$-th column of $X$.
+Proposition 1.
+Assume that $g:(\theta,X)\in\Theta\times\mathcal{X}\mapsto X\theta$ where $\Theta=\{\theta\in\mathbb{R}^{m}:\>\mathopen{}\mathclose{{\left\|\theta}}\right\|_{2}=1\}$ and $X=(X_{1}\,|\,\ldots\,|\,X_{m})\in\mathbb{R}^{d\times m}$ where $X_{j}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(\bm{0},\bm{I})$.
+Then, H˜2 is true for any $\theta\in\Theta$ with
+$C_{0}=\frac{d(d-1)}{2\sqrt{\pi}}\quad\text{and}\quad\beta=1\;.$
+Proof.
+Let $\theta\in\Theta$.
+In what follows, we denote $i=I(\theta)$ where $I(\theta)$ is defined as in H˜2.
+Since $\mathcal{W}$ is the simplex of $\mathbb{R}^{d}$, $v_{j}$ is the $j$-th vector of the canonical basis of $\mathbb{R}^{d}$.
+For any $\varepsilon>0$ we have:
+$\mathbb{P}\mathopen{}\mathclose{{\left(\inf_{j\neq i}{\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle X\theta,v_{j}}}\right\rangle-\mathopen{}\mathclose{{\left\langle X\theta,v_{i}}}\right\rangle}}\right|}\,\leqslant\,\varepsilon}}\right)$ $\leqslant\mathbb{P}\mathopen{}\mathclose{{\left(\bigcup_{j\neq i}\,\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle X\theta,v_{j}}}\right\rangle-\mathopen{}\mathclose{{\left\langle X\theta,v_{i}}}\right\rangle}}\right|\,\leqslant\,\varepsilon}}\right\}}}\right)$
+$\leqslant\sum_{j\neq i}\mathbb{P}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle\theta,X^{\mathsf{\scriptscriptstyle T}}(v_{j}-v_{i})}}\right\rangle}}\right|\,\leqslant\,\varepsilon}}\right)$
+$=\sum_{j\neq i}\mathbb{P}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle\theta,X_{j}-X_{i}}}\right\rangle}}\right|\leqslant\varepsilon}}\right)\;.$ (10)
+Since for any $j\neq i$, $\mathopen{}\mathclose{{\left\langle\theta,X_{j}-X_{i}}}\right\rangle\sim\mathcal{N}(0,2\mathopen{}\mathclose{{\left\|\theta}}\right\|_{2}^{2})$ with density $f_{j}$, we have:
+$\mathbb{P}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle\theta,X_{j}-X_{i}}}\right\rangle}}\right|\,\leqslant\,\varepsilon}}\right)$ $=\int_{-\varepsilon}^{\varepsilon}f_{j}(x)\,\mathrm{d}x=\mathopen{}\mathclose{{\left(2\sqrt{\pi}\mathopen{}\mathclose{{\left\|\theta}}\right\|_{2}}}\right)^{-1}\int_{-\varepsilon}^{\varepsilon}\exp\mathopen{}\mathclose{{\left(-\frac{x^{2}}{4\mathopen{}\mathclose{{\left\|\theta}}\right\|_{2}^{2}}}}\right)\,\mathrm{d}x$
+and since the integrand reaches its maximum in $x=0$,
+$\leqslant\varepsilon\mathopen{}\mathclose{{\left(\sqrt{\pi}\mathopen{}\mathclose{{\left\|\theta}}\right\|_{2}}}\right)^{-1}\;.$ (11)
+Then, plugging (11) in (10) and summing over all pairs $(i,j)$ such that $j\neq i$ yields the desired result. ∎
+Note that the $d^{2}$ factor in $C_{0}$ comes from the use of an union bound, and could be reduced with a refined analysis.
+Appendix B Extended discussion on the relevance of approximate oracles.
+The notion of an approximate oracle reflects the fact that, in non-convex settings, we cannot rely on subroutines that provably reach a global minimizer of $\tilde{f}_{t}$ at time $t$ (as offline gradient descent would in convex problems).
+Instead, we must settle for local minimizers, whose quality is governed by a parameter $\xi$—which vanishes in favorable loss landscapes.
+As a concrete example, consider $\mathbf{O}$ as the stochastic gradient descent (SGD) algorithm.
+Even in the context of deep networks, it is plausible that $\mathbf{O}$ converges to a local minimizer.
+Indeed, a large body of work has analyzed SGD in non-convex settings, establishing convergence to stationary points either in expectation [31] or almost surely [55, 58, 20].
+Such stationary points may correspond to local/global minima or saddle points.
+Recent studies further show that SGD avoids saddle points.
+For instance, Mertikopoulos et al.
+[55] proved that the trajectories of SGD almost surely avoid all strict saddle manifolds—i.e., sets of critical points $x$ where the Hessian has at least one negative eigenvalue.
+These manifolds include connected families of non-isolated saddle points, a phenomenon common in the loss landscapes of overparametrized neural networks [48].
+Beyond SGD, similar guarantees extend to more general methods.
+In particular, the stochastic Riemannian Robbins–Monro method (a broad template encompassing various algorithms) has been shown to converge almost surely to local or global minima [40].
+Appendix C Supplementary results on the simplex.
+We now provide convergence guarantees for Algorithms˜1 and 2 when $\mathcal{W}$ is the simplex. To do so we choose another regularizer than done in the main document.
+C.1 Additional framework
+We focus on the case where $\mathcal{W}$ is the simplex of $\mathbb{R}^{d}$, that is
+$\mathcal{W}_{0}=\mathopen{}\mathclose{{\left\{w\in\mathbb{R}^{d}:w\succcurlyeq 0,\;{\bf 1}^{\mathsf{\scriptscriptstyle T}}w=1}}\right\}\;.$
+This setting encompasses various decision-making scenarios such as portfolio selection [47] or mixed strategy design in multiplayer games [68]. In this case, we choose the negative entropy
 $\mathcal{R}_{0}:w\mapsto\sum_{i\in[d]}w_{i}\ln(w_{i})\;,$
@@
 The minimizer $\tilde{w}_{t}(\theta)\in\mathcal{W}_{0}$ in (5) with $\mathcal{R}=\mathcal{R}_{0}$ can easily be shown to be the softmax mapping, that is it satisfies for any $i\in[d]$:
-$\tilde{w}_{t,i}(\theta)=\frac{\exp(-\alpha_{t}^{-1}g_{i}(\theta,X_{t}))}{\sum_{k\in[d]}\exp(-\alpha_{t}^{-1}g_{k}(\theta,X_{t}))}\;.$ (8)
+$\tilde{w}_{t,i}(\theta)=\frac{\exp(-\alpha^{-1}g_{i}(\theta,X_{t}))}{\sum_{k\in[d]}\exp(-\alpha^{-1}g_{k}(\theta,X_{t}))}\;.$
 It is clear from this expression that $\tilde{w}_{t}$ is differentiable, and that for any $\theta\in\Theta$:
-$\nabla\tilde{w}_{t}(\theta)=-\frac{1}{\alpha_{t}}\mathopen{}\mathclose{{}\left[\mathrm{diag}[\tilde{w}_{t}(\theta)]-\tilde{w}_{t}(\theta)\tilde{w}_{t}(\theta)^{\mathsf{\scriptscriptstyle T}}}\right]\nabla_{\theta}g(\theta,X_{t})\;.$ (9)
-The regularized minimizer (8) and its Jacobian (9) can readily be used in Algorithm 1. In this case, the regret (4) incurred by Algorithm 1 can be bounded as follows.
+$\nabla\tilde{w}_{t}(\theta)=-\frac{1}{\alpha}\mathopen{}\mathclose{{\left[\mathrm{diag}[\tilde{w}_{t}(\theta)]-\tilde{w}_{t}(\theta)\tilde{w}_{t}(\theta)^{\mathsf{\scriptscriptstyle T}}}}\right]\nabla_{\theta}g(\theta,X_{t})\;.$ (12)
+C.2 Results
+DF-FTPL
 Proposition 2.
-Assume H 1, H 2, H 3, $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and that Algorithm 1 is instantiated with a non-increasing sequence $\mathopen{}\mathclose{{}\left(\eta_{t}}\right)_{t\in[T]}$.
-Then, denoting $P_{T}=\sum_{t=2}^{T}\mathopen{}\mathclose{{}\left\|\vartheta_{t}-\vartheta_{t-1}}\right\|$,
-$T^{-1}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}G\eta_{t}}{32T\alpha_{t}^{2}}$
-$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2\ln(d)C_{0}\{\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d)\}]+\xi\;.$
-The term $P_{T}$–which ultimately depends on the sequence $(X_{1},\bar{g}_{1}(X_{1})),\ldots,(X_{t},\bar{g}_{T}(X_{T}))$– controls the variation of the problem through time.
-Proposition 1 makes clear that keeping the dynamic regret (4) low becomes difficult when the revealed cost vectors vary drastically over time.
-Our guarantee does not depend on the dimension $m$ of $\Theta$, a notable strength stemming from the optimistic approach.
-This is sharp contrast with non-optimistic algorithms, such as DF-FTPL in Appendix B, which suffers a $m^{2}$ factor in its associated static regret bound (Proposition 4).
-In addition to the usual trade-off associated with the learning rate $\eta_{t}$, Proposition 1 highlights a new trade-off introduced by the regularization strength $\alpha_{t}$.
-As $\alpha_{t}$ increases, the second term decreases as the mapping $\tilde{f}_{t}$ becomes smoother in the sense that its Lipschitz constant decreases (see Lemma 2).
-However, the third term grows larger as $\tilde{w}_{t}(\theta)$ becomes a poorer approximation of $w^{\star}_{t}(\theta)$.
-It is therefore natural to wonder how to pick the sequences $(\eta_{t})_{t\in[T]}$ and $(\alpha_{t})_{t\in[T]}$ to address these trade-offs.
-In the following corollary, we define for any $t\in[T]$,
-$P_{t}=\begin{cases}\sum_{s=1}^{t}\mathopen{}\mathclose{{}\left\|\vartheta_{s}-\vartheta_{s-1}}\right\|&\text{if }t\geqslant 2\;,\\ 0&\text{otherwise}\end{cases}\;.$
-We prove that under a mild assumption regarding the growth of $(P_{t})_{t\in[T]}$, we can pick $(\alpha_{t})_{t\in[T]}$ and $(\eta_{t})_{t\in[T]}$ in a fully adaptative way to obtain an explicit regret bound.
-Corollary 1.
-Assume H 1,H 2, H 3, $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and $(t^{-1}(1+P_{t}))_{t\in[T]}$ is non-increasing almost surely.
-Then, using Algorithm 1 with $\alpha_{t}\propto t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto t^{-3/4}(1+P_{t})^{3/4}$ for any $t\in[T]$ gives:
-$\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]}{T}=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\ln(d)\mathopen{}\mathclose{{}\left(\frac{1+\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{T}}\right)^{1/4}+\xi}\right)\;,$
-where $\tilde{\mathcal{O}}$ contains polynomial dependencies in $\ln(1/\alpha),\ln(\ln(d))$.
-Note that the resulting expected regret bound only depends logarithmically on the dimension of the decision space.
-We stress that the assumption about $(t^{-1}(1+P_{t}))_{t\in[T]}$ being non-increasing can be weakened to the existence of a non-increasing sequence $(r_{t})_{t\in[T]}$ such that $t^{-1}P_{t}\leqslant r_{t}$ for any $t\in[T]$ without affecting the result.
-This assumption is merely technical, and we believe that Corollary 1 would hold without it up to slight modifications.
-Finally, under the setting of Corollary 1, we remark that $\eta_{t}$ and $\alpha_{t}$ are computable on the fly as they only require the knowledge of $P_{t}$, which depends itself on observed variables $(\vartheta_{s})_{s\in[T]}$.
-This differs, for instance, from the dynamic regret bound of OGD in (hazan2023introductiononlineconvexoptimization, , Chapter 10) which requires the knowledge of $P_{T}$ before training to reach their optimal rate.
-We conclude by mentioning that in Appendix D, we consider a stronger version of H 2 which allows us to derive a counterpart of Proposition 1 and Corollary 1 for a stronger notion of dynamic regret.
-4.2 Generic polytope
-In Section 4.1, the very specific structure of the simplex $\mathcal{W}_{0}$ hinted us to choose $\mathcal{R}$ to be the negative entropy.
-However many problems, such as shortest-path, min-cost-flow lin2014quickest or unit commitment problems saravanan2013solution , require optimizing on a generic polytope.
-In what follows, we go back to the generic case where $\mathcal{W}=\mathcal{W}_{1}$ is a bounded, convex polytope with non-empty interior.
-We can write $\mathcal{W}_{1}$ as the intersection of $n$ half-spaces, denoting by $n>0$ its number of faces, that is
-$\mathcal{W}_{1}=\{w\in\mathbb{R}^{d}:\>A^{\mathsf{\scriptscriptstyle T}}w-b\preccurlyeq 0\}\;,$
-where $A\in\mathbb{R}^{d\times n},b\in\mathbb{R}^{n}$.
-In what follows, we assume that $\mathcal{W}_{1}$ is not degenerated, i.e.
-$AA^{\mathsf{\scriptscriptstyle T}}\in\mathbb{R}^{d\times d}$ is full-rank.
-To tackle this more general structure, our strategy is to use a regularizer $\mathcal{R}$ such that $\tilde{w}_{t}$ always lies in the strict interior of $\mathcal{W}_{1}$, in which case the Jacobian map $\theta\mapsto\nabla\tilde{w}_{t}(\theta)$ can be determined by the first order condition and the implicit function theorem.
-To this end, we propose to use log-barriers as the regularizer in (5), that is
-$\mathcal{R}_{1}:w\mapsto-\sum_{i=1}^{n}\ln(b_{i}-A_{i}^{\mathsf{\scriptscriptstyle T}}w)\;,$ (10)
-where $A_{i}\in\mathbb{R}^{d}$ is the $i$-th column of $A$. With this choice of regularization, we show in Lemma 6 of Section F.2 that:
-$\nabla\tilde{w}_{t}(\theta)=-\alpha_{t}^{-1}\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}(A_{i}^{\mathsf{\scriptscriptstyle T}}\tilde{w}_{t}(\theta)-b_{i})^{-2}A_{i}A_{i}^{\mathsf{\scriptscriptstyle T}}}\right)^{-1}\nabla_{\theta}g(\theta,X_{t})\;,$ (11)
-which can be readily used in Algorithm 1.
-Note that (11) involves inverting a sum of rank-one matrices—a common operation in interior point methods—which can be performed efficiently; see jiang2020faster .
-With this choice, we can establish the following dynamic regret bound.
+Assume H˜1, H˜2 and having access to an $\xi$-approximate optimization oracle $\mathbf{O}_{\xi}$ adapted to $\mathopen{}\mathclose{{\left\{\sum_{i=1}^{t}\tilde{f}_{i}-\mathopen{}\mathclose{{\left\langle\sigma_{t},\cdot}}\right\rangle}}\right\}_{t\in[T]}$.
+Fix $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$.
+Let $\{\theta_{t}\}_{t\in[T]}$ be the output of DF-FTPL (Algorithm˜1) instantiated with learning step $\eta>0$ and regularization coefficients $\alpha_{t}=\alpha>0$ for any $t$.
+Then:
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $\leqslant\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi+\alpha\ln(d)}}\right)\;,$
+where $\mathbb{E}$ denotes the expectation on both data and the intrinsic randomness of DF-FTPL and $\tilde{\mathcal{O}}$ contains polynomial dependency in $\ln(1/\alpha),\ln(\ln(d))$.
+The proof of this result is postponed to Section˜G.4.
+DF-OGD
 Proposition 3.
-Assume H 1, H 2, H 3, $\mathcal{W}=\mathcal{W}_{1}$, $AA^{\mathsf{\scriptscriptstyle T}}$ is full-rank, $\mathcal{R}=\mathcal{R}_{1}$ and Algorithm 1 is instantiated with a non-increasing sequence $\mathopen{}\mathclose{{}\left(\eta_{t}}\right)_{t\in[T]}$.
-Then,
-$T^{-1}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{}\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}^{2}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}\right)^{2}$
-$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2n\max_{k\in[K]}\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{1}C_{0}(\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d))]+\xi\;,$
-where $P_{T}$ is defined as in Proposition 1.
-Proposition 2 closely resembles to Proposition 1, but with an additional linear dependence on the number of faces $n>0$.
-This can be seen as the cost of operating over a more complex admissible space than the simplex.
-As before, it is possible to choose sequences $(\alpha_{t})_{t\in[T]}$ and $(\eta_{t})_{t\in[T]}$ to balance the trade-off discussed in Section 4.1 in a fully adaptative way.
-Corollary 2.
-Assume H 1, H 2, H 3, $\mathcal{W}=\mathcal{W}_{1}$, $\mathcal{R}=\mathcal{R}_{1}$ and $(t^{-1}(1+P_{t}))_{t\in[T]}$ is non-increasing almost surely.
-Then, using Algorithm 1 with $\alpha_{t}\propto t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto t^{-3/4}(1+P_{t})^{3/4}$ for any $t\in[T]$ leads to :
-$\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]}{T}=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(n\frac{1+\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{T}^{1/4}+\xi}\right)\;.$
-Note that since we assume $AA^{\mathsf{\scriptscriptstyle T}}$ to be full rank, we have in particular that $n\geqslant d$, which is a deteriorated dependence compared to the $\ln(d)$ factor of Corollary 1.
-5 Experiments
-00111122223333444455550011112222333344445555Value of Item 1Value of Item 2
-Figure 1: Figure 2 from mandi2024decision .
-Setting.
-[…]
-After having made their decision, the decision-maker observe the true item costs $\bar{g}_{t}(X_{t})=\mathbb{E}[v_{t}\,|\,X_{t}]$, and update $\theta_{t+1}$ for the next round based on this feedback.
-Note that this setting can directly be mapped in the simplex example of Section 4.2, since $a^{\star}_{t}(\theta_{t})=w^{\star}_{t}(\theta_{t})$ with $w^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{w\in\mathcal{W}_{0}}\mathopen{}\mathclose{{}\left\langle w,g(\theta_{t},X_{t})}\right\rangle$, $\mathcal{W}_{0}$ being the simplex of $\mathbb{R}^{K}$.
-[…]
-Synthetic data.
-We instantiate the previous knapsack problem with the following synthetic data.
-For any $t\in[T]$, we draw $X_{t}\in\mathbb{R}^{K\times p}$ with correlated rows, and generate $v_{t}\in[0,1]^{k}$ as follows:
-$v_{t}=\alpha\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon_{t}\;,$ (12)
-where $\alpha>0$, $\varepsilon_{t}\sim\mathrm{N}(0,I_{K})$ is a Gaussian noise and $\theta^{\star}_{t}\in\mathbb{R}^{p}$ is a parameter which follows an autoregressive recursion:
-$\theta^{\star}_{t}=\frac{1}{2}\theta^{\star}_{t-1}+\frac{1}{2}\zeta_{t}\,,\quad\text{where}\quad\zeta_{t}\sim\mathrm{N}(0,I_{p})\;.$
-This is a challenging data generating process, since $\bar{g}_{t}(X_{t})=\sin^{4}((2X_{t}^{\mathsf{\scriptscriptstyle T}}\theta^{\star}_{t})^{-1})$ is non-stationary and highly non-linear, and features are correlated.
-Equation 12 is discussed more in detail in Appendix C.
-To predict $\hat{v}_{t}$ from $X_{t}$, we assume that the decision-maker has access to a class of linear predictors of the form $g:(\theta,X)\mapsto X\theta$.
-Comparison between DFL and PFL.
-In this framework, we compare DF-OGD  (Algorithm 1) to a Prediction-Focused Online Gradient Descent (PF-OGD) described in Algorithm 2, which minimizes the loss $\ell^{\mathrm{mse}}:(\theta,v,X)\mapsto\mathopen{}\mathclose{{}\left\|v-X\theta}\right\|^{2}$.
-Algorithm 2 Prediction-Focused Online Gradient Descent (PF-OGD).
-1:  Input: horizon $T>0$, initialization $\theta_{1}\in\Theta$.
-2:  for each $t\in\{1,\ldots,T\}$ do
-3:     Observe $X_{t}$, predict $\hat{v}_{t}=g(\theta_{t},X_{t})$
-4:     Play $a^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{i\in[K]}\hat{v}_{t}(i)\;,$ observe $\bar{g}_{t}(X_{t})$.
-5:     Update $\theta_{t+1}=\Pi_{\Theta}(\theta_{t}-\eta\nabla\ell^{\mathrm{mse}}(\theta_{t},\bar{g}_{t}(X_{t}),X_{t}))\;.$
-6:  end for
-In Figure 2, we plot on the left-hand panel the average cumulated cost $t\mapsto t^{-1}\sum_{i=1}^{t}v_{t}(a^{\star}_{t}(\theta_{t}))$ incurred by DF-OGD and PF-OGD over 70 runs of $T=2000$ timesteps, as well as the associated 95% confidence intervals.
-On the right-hand panel, we plot the average Mean Square Error resulting from the sequence of prediction parameters $(\theta_{t})_{t\in[T]}$.
-Interestingly, DF-OGD produces parameters that induce better decisions as indicated by the lower average cumulated loss, but degraded predictions as suggested by the constantly higher MSE.
-This is in line with the intuition according to which the decision-focused approach sacrifices accuracy for the sake of better downstream decisions.
-Figure 2: Average cumulated cost (left) and prediction errors (right) of DF-OGD and PF-OGD.
-We conclude this section by mentioning that we perform additional numerical experiments to study how the relative performances of DFL and PFL are affected by model mispecification in Appendix C.
-This experiment empirically supports the idea that DFL outperforms PFL when models are clearly mispecified.
-6 Conclusion.
-Decision-focused learning is a promising approach for better leveraging prediction in real-life decision making processes.
-While the theoretical analysis of decision-focused learning is currently restricted to the batch setting, we extend it to the online framework.
-This allows us to consider non-stationary data processes and varying objective functions.
-We propose an algorithm based on regularization, perturbation and optimism which only requires picking a regularizer along with a sequence of regularization coefficients and learning rates.
-In particular, we show how to pick these hyper-parameters in two canonical setups, namely when the decision space is the simplex, and when it is a general bounded, convex polytope with non-empty interior.
-In both cases, we provide theoretical guarantees for our algorithm, in the form of upper bounds on the expected average dynamic regret.
-Finally, we compare the empirical performances of our approach to a more standard prediction-focused learning procedure in a knapsack experiment with a mispecified prediction model.
-In this experiment, our method outperforms the PFL one, which supports the idea that DFL leads to decision that are more robust to model mispecification.
-We believe that this work can be extended in several ways.
-First, it would be interesting to study whether meaningful guarantees could be obtained under weaker assumptions than ours.
-Second, we believe that alternative smoothing techniques could be successfully employed, such as Moreau-Yosida transform and proximal operator, or random perturbation and Fenchel conjugate techniques berthet2020learning .
-Analyzing the theoretical performance of these methods would be of significant interest.
-Third, it would be valuable to investigate less adversarial environments, such as those involving i.i.d.
-data, to explore whether stronger theoretical guarantees could be obtained and whether novel algorithmic designs might emerge.
-Finally, implementing our method in more ambitious experimental settings would be of great interest from a practical point of view.
-Appendix A A brief recap on online learning
-Online learning (OL) is a paradigm in which data is processed sequentially, either because the practitionner does not collect all data prior to analysis or because the dataset dynamically evolves through time, or simply because handling batch of data is numerically too demanding.
-From the seminal work of [77], which proposed an online version of the celebrated gradient descent algorithm, OL has been intensively studied (see [36] for an overview).
-The classical performance criterion of an online learning algorithm is the static regret.
-Given a sequence of loss functions $(\ell_{t}:\mathcal{K}\rightarrow\mathbb{R})_{t\geqslant 1}$, the static regret compares the efficiency of a sequence of decisions $\hat{\mu}=(\hat{\mu}_{t})_{t\geqslant 1}$ to the best fixed strategy for $T>0$:
-$\operatorname{Regret}_{T}^{\mathrm{S}}(\hat{\mu})=\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\inf_{\mu_{0}\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mu_{0}),$ (13)
-Classical upper bounds on static regret involve a sub-linear rate. For instance, [77] proposed a $\mathcal{O}(\sqrt{T})$ bound for Online Gradient Descent (OGD) which is valid for convex losses.
-Static regret may not be sufficient to assert the efficiency of an online algorithm as the class of static strategies is limited compared to all possible strategies.
-Hence the notion of dynamic regret introduced by [77] and further developed by [34].
-For any sequence $\hat{\mu}$ of predictors and any sequence $\mu$ of dynamic strategies, the dynamic regret is given by
-$\operatorname{Regret}_{T}(\hat{\mu},\mu)=\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\sum_{t=1}^{T}\ell_{t}(\mu_{t})\,,\qquad T\geq 1\,.$
-Dynamic regret has attracted a lot of interest recently, especially when the comparator sequence is set to $(\inf_{\mu\in\mathcal{K}}\ell_{t}(\mu))_{t\geqslant 1}$ (worst-case dynamic regret, as in [13, 74]) but also for any comparator sequence (universal dynamic regret, as in [75]).
-Note that dynamic regret is a stronger notion that static regret, since it compares the performance of $\hat{\mu}$ to the best dynamic strategy ($\inf_{\mu\in\mathcal{K}}\ell_{t}(\mu)$), while static regret relies on the best static strategy $(\inf_{\mu\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mu))$.
-Appendix B An online decision-focused algorithm with static regret guarantees
-In our work, we mainly focused on upper bounds of the dynamic regret, given its versatility and its ability to accommodate non-stationary environments.
-However, static regret as defined in (13) is a more common measure of performance.
-In this subsection, we show that it is possible to extend our analysis to static regret based on techniques from [64].
-More precisely, we propose in what follows a decision-focused version of the Follow-The-Perturbed-Leader (FTPL) algorithm for nonconvex losses.
-We first introduce some necessary background in Section B.1, and apply it to Decision-Focused learning in Section B.2.
-B.1 Background on FTPL.
+Assume H˜1, H˜2, and having access to an $\xi$-approximate optimization oracle $\mathbf{O}_{\xi}$ adapted to $\mathopen{}\mathclose{{\left\{\tilde{f}_{i}}}\right\}_{t\in[T]}$.
+Fix $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$.
+Let $\{\theta_{t}\}_{t\in[T]}$ be the output of DF-OGD (Algorithm˜2) instantiated with the non-increasing sequence $\mathopen{}\mathclose{{\left(\eta_{t}}}\right)_{t\in[T]}$ and regularization coefficients $(\alpha_{t})_{t\in[T]}$.
+Then:
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{\left[P_{T}}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}G\eta_{t}}{32T\alpha_{t}^{2}}$
+$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2\ln(d)C_{0}\{\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d)\}]+\xi\;,$
+where $P_{T}=\sum_{t=1}^{T-1}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|$.
+The proof of this result is postponed to Section˜G.5.
+Note that the principal gain, compared to the general polytope case, is to avoid the dependency in $n$ the number of faces and enjoys a nice $\ln(d)$ dependency in the dimension of the decision space.
+Appendix D Background on the FTPL algorithm
 Approximate Optimization Oracle.
@@
 $\ell\mathopen{}\mathclose{{\left(\mu^{*}}}\right)-\mathopen{}\mathclose{{\left\langle\sigma,\mu^{*}}}\right\rangle\leqslant\inf_{\mu\in\mathcal{K}}[\ell(\mu)-\langle\sigma,\mu\rangle]+\mathopen{}\mathclose{{\left(\xi+\chi\|\sigma\|_{1}}}\right)\;.$ (13)
-We denote the output of such an optimization oracle by $\mu\star=\mathcal{O}_{\xi,\chi}(\ell-\langle\sigma,\cdot\rangle)$.
-Note the property (14) is very close from the assumption H 3 we made in Section 3 on the expert sequence $(\vartheta_{t})_{t\in[T]}$ (in this case $\chi=0$ as no $\sigma$ is involved).
+We denote the output of such an optimization oracle by $\mu^{*}=\mathcal{O}_{\xi,\chi}(\ell-\langle\sigma,\cdot\rangle)$.
+Note the notion of oracle described in (13) is very close from the Definition˜1 we made in Section˜3 on the expert sequence $(\vartheta_{t})_{t\in[T]}$ (in this case $\chi=0$ as no $\sigma$ is involved).
+Remark 2.
+Note that, in our work, we made the choice to fix $\chi=0$. This stronger assumption is due to the will of having a unifying framework encompassing both the setup for DF-FTPL and DF-OGD.
 Follow The Perturbed Leader.
@@
 $\hat{\mu}_{t+1}=\mathcal{O}_{\xi,\chi}\mathopen{}\mathclose{{\left(\sum_{i=1}^{t-1}\ell_{i}-\mathopen{}\mathclose{{\left\langle\sigma_{t},\cdot}}\right\rangle}}\right)$ (14)
-where $\sigma_{t}\in\mathbb{R}^{d}$ is a random perturbation such that $\sigma_{t,j}$, the j-th ccoordinate of $\sigma_{t}$, is sampled from $\operatorname{Exp}(\eta)$, the exponential distribution with parameter $\eta$.
+where $\sigma_{t}\in\mathbb{R}^{d}$ is a random perturbation such that $\sigma_{t,j}$, the j-th coordinate of $\sigma_{t}$, is sampled from $\operatorname{Exp}(\eta)$, the exponential distribution with parameter $\eta$.
 Following this update route, the addition of an exponential noise allowing to handle the non-convexity of the losses, they reach the following static regret bound.
@@
 Suppose that $\ell_{t}$ is $L$-Lipschitz w.r.t L1 norm, for all $t\in[T]$.
-For any fixed $\eta$, FTPL (Equation 15) with access to a $(\xi,\chi)$-approximate" optimization oracle satisfies the following static regret bound:
-$\frac{1}{T}\operatorname{Regret}_{T}^{\mathrm{S}}(\hat{\mu})\leqslant\mathcal{O}\mathopen{}\mathclose{{}\left(\eta m^{2}DL^{2}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi mL}\right),$
-where $\operatorname{Regret}_{T}^{\mathrm{S}}(\hat{\mu})$ is defined as in (13), and $\mathbb{E}$ denotes the expectation over $\sigma_{1}\cdots\sigma_{T}$.
-B.2 Application to Decision-Focused Learning.
-We now show how to adapt the techniques and results from Section B.1 to our online decision-focused problem.
-First, we instantiate the static regret defined (13) in our framework.
-For any $T>0$, loss sequence $(f_{t})_{t\in[T]}$ as defined in (3), and algorithms $(\mathsf{Alg}_{t})_{t\in[T]}$ outputting $(\theta_{t})_{t\in[T]}$, we define for any $\theta\in\Theta$:
-$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)=\sum_{t\in[T]}F_{t}(\theta_{t})-\sum_{t\in[T]}F_{t}(\theta)\;.$ (16)
-where $F_{T}:\theta\rightarrow\mathbb{E}\mathopen{}\mathclose{{}\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}\right]$ is the conditional expectation of the loss $f_{t}$ with respect to the history $\mathcal{H}_{t-1}$.
-Note that Proposition 3 holds for any sequence of lipschitz losses, and that the surrogate losses ($\tilde{f}_{t}$) are indeed Lipschitz for our choice of regularizers in Section 4.1 and Section 4.2 (see Lemmas 2 and 7).
-from this observation, we derive a non-optimistic online algorithm following the same pattern as DF-OGD (Algorithm 1), but which leverages FTPL to enjoy the guarantee proposition 3.
-Algorithm 3 Decision-Focused Follow The Perturbed Leader DF-FTPL
-1:  Input: horizon $T>0$, initialization $\theta_{1}\in\Theta$.
-2:  for each $t\in\{1,\ldots,T\}$ do
-3:     Observe $X_{t}\in\mathcal{X}$ and play $w^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\mathopen{}\mathclose{{}\left\langle g(\theta_{t},X_{t}),w}\right\rangle\;.$
-4:     Observe $\bar{g}_{t}(X_{t})\in\mathcal{Z}$ and update the history $\mathcal{H}_{t}$.
-5:     Draw $\sigma_{t}\in\mathbb{R}^{d}$ such that for all $j\in[d]$, $\sigma_{j,t}\sim\mathrm{Exp}(\eta)$.
-6:     Update
-$\theta_{t+1}=\mathcal{O}_{\xi,\chi}\mathopen{}\mathclose{{}\left(\sum_{i=1}^{t}\tilde{f}_{i}-\mathopen{}\mathclose{{}\left\langle\sigma_{t},\cdot}\right\rangle}\right)$
-7:  end for
-In Algorithm 3, recall that $\tilde{f}_{i}$ are the surrogate losses defined in Equation 5.
-At each time $t\in[T]$, Algorithm 3 first makes a decision $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ based on the current parameters $\theta_{t}\in\Theta$.
-Then, it observes $\bar{g}_{t}(X_{t})\in\mathcal{Z}$, determines the surrogate objective $\tilde{f}_{t}$.
-Next, it updates $\theta_{t+1}$ following the FTPL algorithm applied to the surrogate losses $(\tilde{f}_{t})_{t\in[T]}$.
-In order to obtain static regret bounds, we follow a similar route than in Section 4.1: we work on the simplex $\mathcal{W}_{0}$ and use the negative entropy $\mathcal{R}_{0}$ as regularizer to define $(\tilde{f}_{t})_{t\in[T]}$.
-We refer to Section 4.1 for a precise definition of these notions.
-We then have the following result.
-Proposition 4.
-Assume H 1, H 2 and having access to an ($\xi,\chi$ )-approximate optimization oracle $\mathcal{O}$.
-Fix for any $t$, $\alpha_{t}=\alpha$ in (6), $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and that DF-FTPL (Algorithm 3) is instantiated with $\eta>0$.
-Then, for any $\theta\in\Theta$:
-$T^{-1}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{S}}(\theta)}\right]$ $\leqslant\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi m\frac{1}{\alpha}+\alpha\ln(d)}\right)\;,$
-where $\mathbb{E}$ denotes the expectation on both data and the intrinsic randomness of DF-FTPL and $\tilde{\mathcal{O}}$ contains polynomial dependency in $\ln(1/\alpha),\ln(\ln(d))$.
-Thus, it is possible to have a decision-focused algorithm with static regret guarantees.
-A notable strength of Algorithm 3 compared to Algorithm 1 is that it does not require any optimism to run.
-However, this comes at a major cost: a dependency in the dimension $m^{2}$ of $\Theta$ which is not involved in Proposition 1 thanks to the optimistic approach.
-The proof of Proposition 4 is deferred to Section F.3.
-Appendix C Additional material on our numerical experiment.
-In this section, we provide more detail about the experimental setup in Section 5 and perform additional numerical simulations.
+For any fixed $\eta$, FTPL (Equation˜14) with access to a $(\xi,\chi)$-approximate" optimization oracle satisfies the following static regret bound:
+$\frac{1}{T}\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\inf_{\mu}\sum_{t=1}^{T}\ell_{t}(\mu)\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}DL^{2}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi mL}}\right),$
+where $\mathbb{E}$ denotes the expectation over $\sigma_{1}\cdots\sigma_{T}$.
+Appendix E Additional experimental material
+The motivating example from [54].
+[…]
+0112233445501122334455Value of Item 1Value of Item 2
+Figure 2: Figure 2 from [54].
 Details on the experimental setup.
+In this paragraph, we provide more detail about the experimental setup in Section˜4 and perform additional numerical simulations.
 We start by explaining more precisely how the data used in our experiment are drawn.
@@
 Second, we generate $v_{t}\in\mathbb{R}^{K}$ as follows:
-$v_{t}=\min\mathopen{}\mathclose{{}\left(\max\mathopen{}\mathclose{{}\left(\tilde{v}_{t},0}\right),1}\right)\quad\text{with}\quad\tilde{v}_{t}=\alpha\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon\;.$
-In the above equation, $\alpha>0$ is a constant, $\theta_{t}^{\star}$ follows the following autoregressive recursion:
-$\theta^{\star}_{t}=\frac{1}{2}\theta^{\star}_{t-1}+\frac{1}{2}\zeta_{t}\,,\quad\text{where}\quad\zeta_{t}\sim\mathrm{N}(0,I_{K})\;,$
+$c_{t}=\min\mathopen{}\mathclose{{\left(\max\mathopen{}\mathclose{{\left(\tilde{c}_{t},0}}\right),1}}\right)\quad\text{with}\quad\tilde{c}_{t}=A\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon\;.$
+In the above equation, $A>0$ is a constant, $\theta_{t}^{\star}$ satisfies:
+$\theta^{\star}_{t}=\frac{1}{2}\theta^{\star}+\frac{1}{2}\zeta_{t}\,,\quad\text{where}\quad\zeta_{t}\sim\mathrm{N}(0,I_{K})\quad\text{and}\quad\theta^{\star}\in\mathbb{R}^{K}\;,$
 and $\varepsilon_{t}\sim\mathrm{N}(0,I_{K})$ is Gaussian noise.
@@
 We consider $K=5$ items and $p=10$ features.
-The horizon is set to $T=2000$.
-For each plot, we run $N=70$ times DF-OGD and PF-OGD.
+The horizon is set to $T=5000$.
+For each plot, we run $N=10$ times DF-OGD and PF-OGD.
 The reported error bars are 95% Gaussian confidence intervals (from the CLT).
-In (12), $\alpha=45$ and the correlation coefficient for $X_{t}$ is $\rho=0.8$.
-Our algorithms are instantiated with the following parameters (optimized through a grid search procedure).
-First, DF-OGD runs with $(\alpha_{t})$ and $(\eta_{t})$ as in Corollary 1.
-The oracle from H 2 is obtained through a SGD algorithm performing $t_{\vartheta}=10$ steps at each iteration with a learning rate set to $\eta_{\vartheta}=0.01$, while being initialized at $\vartheta_{t-1}$.
-On the other hand, PF-OGD runs with a learning rate $\eta=0.01$.
-Our code mostly builds on pytorch and numpy.
+In (9), $A=45$ and the correlation coefficient for $X_{t}$ is $\rho=0.8$.
+Our algorithms are instantiated with the following parameters.
+First, DF-OGD runs with $(\alpha_{t})$ and $(\eta_{t})$ as suggested by our theoretical analysis.
+The oracle from H˜2 is obtained through a SGD algorithm performing $t_{\vartheta}=10$ steps at each iteration with a learning rate set to $\eta_{\vartheta}=0.01$, while being initialized at $\vartheta_{t-1}$.
+On the other hand, PF-OGD runs with a learning rate $\eta=10$ determined by grid search.
+Our code is implemented with pytorch.
+Benchmark algorithms.
+Algorithm 3 Prediction-Focused Online Gradient Descent (PF-OGD).
+1:Input: horizon $T>0$, initialization $\theta_{1}\in\Theta$.
+2:for each $t\in\{1,\ldots,T\}$ do
+3:  Observe $X_{t}$, predict $c_{t}=g(\theta_{t},X_{t})$
+4:  Play $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{i\in[K]}c_{t,i}\;,$ observe $\bar{g}_{t}(X_{t})$.
+5:  Update $\theta_{t+1}=\Pi_{\Theta}(\theta_{t}-\eta\nabla\ell^{\mathrm{mse}}(\bar{g}_{t}(X_{t}),X_{t}\theta_{t}))\;.$
+6:end for
 Deviation from realizability.
-As argued in Section 1, prediction-focused learning would be optimal if the prediction model made no mistake, since in this case problem (2) would perfectly align with (1).
+As argued in Section˜1, prediction-focused learning would be optimal if the prediction model made no mistake, since in this case problem (2) would perfectly align with (1).
 As a consequence, one would expect prediction-focused learning to perform very well in the realizable setting, that is when for any $t\in[T]$, there exists $\bar{\theta}_{t}\in\Theta$ such that $\bar{g}_{t}(X_{t})=g(\bar{\theta}_{t},X_{t})$.
@@
 $v^{(\gamma)}_{t}=(1-\gamma)X_{t}\theta^{\star}_{t}+\gamma\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon_{t}\;.$
-In Figure 3, we make $\gamma$ vary from 0 to 1.
+In Figure˜3, we make $\gamma$ vary from 0 to 1.
 At each value, we plot the average cumulated cost gap between DF-OGD and PF-OGD, averaged over $50$ runs with horizon $T=1000$.
@@
 Figure 3: Average cumulated cost gap PFL - DFL as a function of $\gamma$.
-Appendix D A stronger version of H 2 for a stronger notion of regret
-In this section, we present a stronger version of H 2, displayed in H 4. We show that under this new assumption, our guarantees still hold for a stronger notion of regret.
-Here, we focus on the simplex setting of Section 4.1: we take $\mathcal{W}=\mathcal{W}_{0}$ the simplex and $\mathcal{R}=\mathcal{R}_{0}$ the negative entropy.
-This section show that at the cost of slightly strengthening this assumption, we may obtain regret bounds for DF-OGD with a stronger version of regret, which we define below.
-A novel dynamic regret.
-We define the pseudo dynamic regret as, for any $t$ any loss sequence $(f_{t})_{t\in[T]}$,
-$\mathfrak{R}_{T}^{\mathrm{d}}:=\sum_{t\in[T]}f_{t}(\theta_{t})-\sum_{t\in[T]}\inf_{\theta\in\Theta}f_{t}(\theta).$
-This notion is exactly what appeared in [77].
-Note that since $\mathbb{E}\mathopen{}\mathclose{{}\left[\inf_{\theta\in\Theta}h(\theta,X)}\right]\leqslant\inf_{\theta\in\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[h(\theta,X)}\right]$, we immediately obtain that:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right].$
-In this sense, obtaining a guarantee on $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right]$ is stronger than on $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$.
-Strengthening H 2.
-We propose the following assumption:
-H 4.
-There exist $C_{0}>0$ and $\beta\in\mathopen{}\mathclose{{}\left[0,1}\right]$, such that for any $t\in[T]$ and $\varepsilon\in\mathopen{}\mathclose{{}\left[0,1}\right]$,
-$\mathbb{P}\mathopen{}\mathclose{{}\left(\forall\theta\in\Theta,\>\inf_{j\neq I_{t}(\theta)\;}u_{j}(\theta,X_{t})-u_{I_{t}(\theta)}(\theta,X_{t})\geqslant\varepsilon\,\mid\,\mathcal{H}_{t-1}}\right)\geqslant 1-C_{0}\varepsilon^{\beta}\;,$
-where $I_{t}(\theta)\in\operatorname*{\arg\!\min}_{i\in[K]}u_{i}(\theta,X_{t})$.
-This reinforcement of H 2 comes from the fact that we took $\forall\theta\in\Theta$ inside the probability and not outside. Based on this assumption we adapt Lemma 3 to get the following lemma.
-Lemma 1.
-Assume H 4, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $\theta\in\Theta$ and $t\in[T]$,
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}\right\|_{1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)\;.$
-Lemma 1 shows that assuming H 4 instead of H 2 transform the control of the supremum of an expectation of a supremum (Lemma 3) into one on the expectation of a supremum.
-Exploiting Lemma 1, we prove a novel regret bound involving the deterministic dynamic regret.
-Proposition 5.
-Assume H 1, H 4, H 3, $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and that Algorithm 1 is instantiated with a non-increasing sequence $\mathopen{}\mathclose{{}\left(\eta_{t}}\right)_{t\in[T]}$.
-Then, denoting $P_{T}=\sum_{t=1}^{T-1}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|$,
-$\frac{1}{T}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}G\eta_{t}}{32T\alpha_{t}^{2}}$
-$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2\ln(d)C_{0}\{\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d)\}]+\xi\;,$
-where $\mathbb{E}$ denotes the expectation on both data and the intrinsic randomness of DF-OGD.
-From Proposition 5 we can recover the same convergence rate than Corollary 1 and also perform a similar study than in Section 4.2 to have results for a general polytope.
-The tradeoff highlighted is that strengthening H 2 into H 4 is the price to pay to control the stronger notion of regret to $\mathfrak{R}_{T}^{\mathrm{d}}$.
-The proofs of Lemma 1 and Proposition 5 are deferred to appendix Section F.4.
-Appendix E Regret bound with an approximate optimal solution.
+Algorithm 4 Online Smart Predict-then-Optimize Online Gradient Descent (Online SPO).
+1:Input: horizon $T>0$, initialization $\theta_{1}\in\Theta$.
+2:for each $t\in\{1,\dots,T\}$ do
+3:  Observe $X_{t}$, predict $\hat{c}_{t}=g(\theta_{t},X_{t})$.
+4:  Play $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{v\in\mathcal{\mathcal{V}}}\hat{c}_{t}^{\mathsf{\scriptscriptstyle T}}v$, observe $c_{t}$.
+5:  Compute $v^{\star}(\hat{c}_{t})=\operatorname*{\arg\!\min}_{v\in\mathcal{\mathcal{V}}}\hat{c}_{t}^{\mathsf{\scriptscriptstyle T}}v$ and $v^{\star}(c_{t})=\operatorname*{\arg\!\min}_{v\in\mathcal{\mathcal{V}}}c_{t}^{\mathsf{\scriptscriptstyle T}}v$.
+6:  Define SPO+ surrogate:
+$\ell^{\mathrm{SPO+}}(\hat{c}_{t},c_{t})=2c_{t}^{\mathsf{\scriptscriptstyle T}}v^{\star}(\hat{c}_{t})-c_{t}^{\mathsf{\scriptscriptstyle T}}v^{\star}(c_{t})-\hat{c}_{t}^{\mathsf{\scriptscriptstyle T}}v^{\star}(\hat{c}_{t}).$
+7:  Update $\theta_{t+1}=\Pi_{\Theta}\big(\theta_{t}-\eta\,\nabla_{\theta}\ell^{\mathrm{SPO+}}(\hat{c}_{t},c_{t})\big)$, where the gradient is computed via chain rule through $\hat{c}_{t}=g(\theta_{t},X_{t})$.
+8:end for
+Appendix F Regret bound with an approximate optimal solution.
 In this section, we show that when we cannot access $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ but only $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ such that $\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),\underline{w}_{t}(\theta_{t})-w^{\star}_{t}(\theta_{t})}}\right\rangle\leqslant\kappa$ for $\kappa>0$, our regret bounds are only shifted by $\kappa$.
@@
 Proposition 5.
-Assume the assumptions from Proposition 2 and that at each step, $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ is replaced by $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ which satisfies
+Assume the assumptions from Theorem˜2 and that at each step, $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ is replaced by $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ which satisfies
 $\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),\underline{w}_{t}(\theta_{t})-w^{\star}_{t}(\theta_{t})}}\right\rangle\leqslant\kappa\;,$
 where $\kappa>0$. Then,
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$ $\leqslant\frac{D_{\Theta}^{2}-D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{}\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}\right)^{2}$
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]$ $\leqslant\frac{D_{\Theta}^{2}-D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{\left[P_{T}}}\right]}{2\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{\left\|b-Aw}}\right\|_{\infty}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}}\right)^{2}$
 $+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}(\ln(1\alpha_{t}^{-1})+(1-\beta)^{2}\ln^{2}(\alpha_{t}\ln d))]+\xi+\kappa\;.$
-The proof Proposition 6 is deferred to Section F.5.
-Appendix F Proofs.
-F.1 Proofs of Section 4.1.
+The proof Proposition˜5 is deferred to Section˜G.6.
+Appendix G Proofs
+G.1 Proofs of preliminary lemmas.
+G.1.1 Simplex
 Lemma 1.
-Assume H 1, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$.
+Assume H˜1, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$.
 Then for any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}$-Lipchitz almost surely, with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$.
@@
 Lemma 2.
-Assume H 2, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $\theta\in\Theta$ and $t\in[T]$, for any $\theta\in\theta$
+Assume H˜2, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $\theta\in\Theta$ and $t\in[T]$, for any $\theta\in\theta$
 $\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)}}\right)\;.$
@@
 Proof.
-In this proof, $I_{t}(\theta)$ from H 2 is denoted $I_{t}$ to lighten notation. We have by assumption:
+In this proof, $I_{t}(\theta)$ from H˜2 is denoted $I_{t}$ to lighten notation. We have by assumption:
 $\langle g(\theta,X_{t}),\tilde{w}_{t}(\theta)\rangle$ $=\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),\sum_{j=1}^{d}\tilde{w}_{t,j}(\theta)\,e_{j}}}\right\rangle$
@@
 $\mathbb{E}\mathopen{}\mathclose{{\left[\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\mid\mathcal{H}_{t-1}}}\right]\\ =\alpha_{t}+2\alpha_{t}\ln(d)\int_{(\alpha_{t}\ln(d))}^{(2\ln(d))}\mathbb{P}\mathopen{}\mathclose{{\left(\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>2\frac{\alpha_{t}\ln(d)}{\varepsilon}\,\mid\,\mathcal{H}_{t-1}}}\right)\frac{d\varepsilon}{\varepsilon^{2}}.$ (22)
-Moreover, for any $\varepsilon>0$ we have by H 2 and Lemma 4 that:
+Moreover, for any $\varepsilon>0$ we have by H˜2 and Lemma˜3 that:
 $1-C_{0}\varepsilon^{\beta}$ $\leqslant\mathbb{P}\mathopen{}\mathclose{{\left(\inf_{j\neq I_{t}}\{u_{j}(\theta,X_{t})-u_{I_{t}}(\theta,X_{t})\}\geqslant\varepsilon\,\mid\,\mathcal{H}_{t-1}}}\right)$
@@
 $\varphi(\beta)=(1-\beta)^{-1}(a^{\beta-1}-b^{\beta-1})=(1-\beta)^{-1}[e^{-(1-\beta)\ln(a)}-e^{-(1-\beta)\ln(b)}]\;,$
-[…]
+and since for any $x>0,\,1-x\leqslant e^{-x}\leqslant 1-x+x^{2}$:
 $\varphi(\beta)$ $\leqslant(1-\beta)^{-1}[1-(1-\beta)\ln(a)+(1-\beta)^{2}\ln^{2}(a)-1+(1-\beta)\ln(b)]$
@@
 Substituting in the values of $a$ and $b$ yields the desired result. ∎
+G.1.2 General polytope
+In this section, we denote by $\mathcal{W}_{1}$ the general polytope described in LABEL:eq:polytope and $\mathcal{R}_{1}$ the log-barrier regularization described in Equation˜7.
 Lemma 4.
-Assume that for any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}-$Lipschitz almost-surely and that the sequence of steps $(\eta_{t})_{t\geqslant 1}$ is non-increasing.
-Denote $\tilde{\mathfrak{R}_{T}}=\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})$.
-Then it holds almost surely that:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}_{T}}\right]\leqslant\frac{1}{2\eta_{T}}\mathopen{}\mathclose{{}\left(D_{\Theta}^{2}+2D_{\Theta}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|}\right)+\sum_{t=1}^{T}\frac{\eta_{t}}{2}K_{t}^{2}\;,$
-where the expectation is taken over the sequence $(u_{1},\ldots,u_{T})$ in Algorithm 1.
-Proof.
-In what follows, we write $\mathbb{E}_{u_{t}}$ the expectation under the distribution $\text{Unif}([0,1])$ of $u_{t}$, and $\mathbb{E}_{u_{1:T}}$ the expectation under the joint distribution $\text{Unif}([0,1])^{\otimes T}$ of $(u_{1},\ldots,u_{T})$.
-For any $t\in[T]$, we have:
-[…]
-$=\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left\langle\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\nabla\tilde{f}_{t}(\vartheta_{t}+u_{t}(\theta_{t}-\vartheta_{t}))}\right],\theta_{t}-\vartheta_{t}}\right\rangle$
-$=\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left\langle\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\tilde{\nabla}_{t}(u_{t})}\right],\theta_{t}-\vartheta_{t}}\right\rangle\;,$ (26)
-where ~$\tilde{\nabla}(u_{t})=\nabla\tilde{f}_{t}(\vartheta_{t}+u_{t}(\theta_{t}-\vartheta_{t})$ as in Algorithm 1. To control (26) we note that by definition of $\theta_{t+1}$ in Algorithm 1,
-$\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}}\right]$ $=\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\Pi_{\Theta}[\theta_{t}-\eta_{t}\tilde{\nabla}_{t}(u_{t})]-\vartheta_{t}}\right\|^{2}}\right]$
-$\leqslant\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}-\eta_{t}\tilde{\nabla}_{t}(u_{t})}\right\|^{2}}\right]$
-$\leqslant\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}}\right\|^{2}-2\eta_{t}\mathopen{}\mathclose{{}\left\langle\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{}\left[\tilde{\nabla}_{t}(u_{t})}\right],\theta_{t}-\vartheta_{t}}\right\rangle+\eta_{t}^{2}K_{t}^{2}$ (27)
-where we have used that ~$\mathopen{}\mathclose{{}\left\|\tilde{\nabla}_{t}(u_{t})}\right\|^{2}\leqslant K_{t}^{2}$ because $\tilde{f}_{t}$ is $K_{t}-$Lipschitz.
-Re-arranging (27) and taking the expectation over $(u_{1},\ldots,u_{T})$ yields:
-$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\tilde{\nabla}_{t}(u_{t}),\theta_{t}-\vartheta_{t}}\right\rangle}\right]$ $\leqslant\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\frac{1}{2\eta_{t}}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}}\right\|^{2}-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}}\right)+\frac{\eta_{t}K_{t}^{2}}{2}}\right]\;,$
-Now notice that, for any $t\in\{1,\ldots,T-1\}$,
-$\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|^{2}$ $=\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}+\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|^{2}-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}$
-$=\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}$
-$+(\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|+\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|)(\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|)$
-$\leqslant\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}+2D_{\Theta}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|\;,$ (29)
-where we have used H 1-(i) and the reversed triangular inequality in the last line. It follows that:
-$-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}$ $\leqslant-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|^{2}+2D_{\Theta}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|$ (30)
-Plugging (30) into (F.1) then gives:
-$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\tilde{\nabla}_{t}(u_{t}),\theta_{t}-\vartheta_{t}}\right\rangle}\right]$ $\leqslant\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\frac{1}{2\eta_{t}}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}}\right\|^{2}-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|^{2}}\right)}\right.$
-$\quad\mathopen{}\mathclose{{}\left.+\frac{1}{\eta_{t}}D_{\theta}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|+\frac{\eta_{t}K_{t}^{2}}{2}}\right]\;.$
-Thus, summing over $T$ and plugging the resulting sum in (26) gives:
-$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\sum_{t=1}^{T}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}\right]$ $\leqslant\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\sum_{t=1}^{T}\frac{\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}}\right\|^{2}}{2}\mathopen{}\mathclose{{}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}}\right)}\right.$
-$\quad\mathopen{}\mathclose{{}\left.+\sum_{t=1}^{T}\frac{1}{\eta_{t}}D_{\theta}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|+\sum_{t=1}^{T}\frac{\eta_{t}K_{t}^{2}}{2}}\right]\;.$
-$\leqslant\frac{D_{\Theta}^{2}}{2}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}}\right)+\frac{D_{\theta}}{\eta_{T}}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|+\sum_{t=1}^{T}\frac{\eta_{t}K_{t}^{2}}{2}\;.$
-where we used that for all $t$, $\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}}\right\|\leqslant D_{\Theta}$ by H 1-(i) and that for all $t\in[T],\,\eta_{t-1}\geqslant\eta_{t}\geqslant\eta_{T}$.
-Then, we have by telescoping the first sum:
-$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{}\left[\sum_{t=1}^{T}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}\right]$ $\leqslant\frac{D_{\Theta}^{2}+2D_{\Theta}}{2\eta_{T}}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|+\sum_{t=1}^{T}\frac{\eta_{t}K_{t}^{2}}{2}.$
-See 1
-Proof.
-In this proof, for the sake of conciseness, we denote by $\mathbb{E}_{t}$ the conditional expectation $\mathbb{E}\mathopen{}\mathclose{{}\left[\cdot\mid\mathcal{H}_{t-1}}\right]$.
-Observe that the dynamic regret can be decomposed as follows:
-$\mathfrak{R}_{T}$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}\right\rangle}\right]+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle}\right]$ (31)
-$+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\vartheta_{t})}\right\rangle}\right]-\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle}\right]$ (32)
-$+\sum_{t\in[T]}\mathopen{}\mathclose{{}\left[\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle}\right]-\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}\right\rangle}\right]}\right]$ (33)
-Taking the $\sup$ over $\Theta$, defining by $\mathrm{Reg}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle}\right]$ in (31), using H 3 in (32), and noticing that $\inf_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)]-\inf_{\theta\in\Theta}\mathbb{E}_{t}[\tilde{f}_{t}(\theta)]\leqslant\sup_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)-\tilde{f}_{t}(\theta)]$ in (33) leads to:
-$\mathfrak{R}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\rangle}\right]+\xi T$
-Using the fact that $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{\infty}$ is dual to $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{1}$:
-$\mathfrak{R}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]+\xi T$
-Since for any $t\in[T]$, $\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{2}\leqslant D_{\mathcal{Z}}$ almost surely by H 1-(ii),
-$\mathfrak{R}_{T}$ $\leqslant\mathrm{Reg}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]+\xi T\;$ (34)
-First, notice that, by Lemma 3, for all $t$:
-$\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]$ $\leqslant\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)\;.$ (35)
-Second, taking the expectation (denoted by $\mathbb{E}$) over $(u_{t})_{t\in[T]}$ and $X_{1},\cdots X_{T}$ on both sides gives:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\mathrm{Reg}_{T}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)+\xi T\;.$ (36)
-To bound the first term in the right-hand side of Equation 36. Remark that, by the definition of conditional expectation:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathrm{Reg}_{T}}\right]=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}\right].$
-One recognises the definition of $\tilde{\mathfrak{R}_{T}}$ of Lemma 5.
-Thus, by this lemma, we know that for any $t\in[T]$, because $\tilde{f}_{t}$ is $K_{t}-$Lipschitz with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$ almost surely by Lemma 2, we have:
-$\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}\right]\leqslant\frac{1}{2\eta_{T}}(D_{\Theta}^{2}+D_{\Theta}P_{T})+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}^{2}G^{2}\eta_{t}}{32\alpha_{t}^{2}}\;.$ (37)
-Plugging Equation 37 into Equation 36 and then dividing by $T>0$ on both sides gives the desired result.
-See 1
-Proof.
-We know by Proposition 1 that for any decreasing sequence of learning rates $(\eta_{1},\ldots,\eta_{T})\in\mathbb{R}_{+}^{T}$ and any regularization coefficients $(\alpha_{1},\ldots,\alpha_{T})\in\mathbb{R}_{+}^{T}$, the average regret of Algorithm 1 is bounded as follows:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{1+P_{T}}{\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{\alpha_{t}^{2}}+(1+\ln d)\sum_{t\in[T]}\alpha_{t}+\xi T}\right]}\right)\;.$
-With $\alpha_{t}\propto t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto\alpha_{t}t^{-1/2}(1+P_{t})^{1/2}\propto t^{-3/4}(1+P_{t})^{3/4}$, the sequence $(\eta_{t})_{t\in[T]}$ is non-increasing because so is $(t^{-1}(1+P_{t}))_{t\in[T]}$ by assumption.
-Moreover, the first term satisfies:
-$\frac{1+P_{T}}{\eta_{T}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\frac{T^{3/4}}{(1+P_{T})^{3/4}}\mathopen{}\mathclose{{}\left(1+P_{T}}\right)}\right)\quad\text{so}\quad\frac{1+P_{T}}{T\eta_{T}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(\frac{1+P_{T}}{T}}\right)^{1/4}}\right)\;.$
-For the second term,
-$\sum_{t=1}^{T}\frac{\eta_{t}}{\alpha_{t}^{2}}=\sum_{t=1}^{T}\frac{\sqrt{1+P_{t}}}{\sqrt{t}\alpha_{t}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left(\frac{1+P_{t}}{t}}\right)^{1/4}}\right)\;..$
-Moreover, since $1+P_{t}\leqslant 1+P_{T}$ for any $t\in[T]$, we have:
-$\sum_{t=1}^{T}\frac{\eta_{t}}{\alpha_{t}^{2}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(1+P_{T}}\right)^{1/4}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left(\frac{1}{t}}\right)^{1/4}}\right)=\mathcal{O}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(1+P_{T}}\right)^{1/4}T^{3/4}}\right)\;,$
-where we used in the last line that $\sum_{t=1}^{T}t^{-\gamma}=\mathcal{O}\mathopen{}\mathclose{{}\left(T^{1-\gamma}}\right)$ for $\gamma\in(0,1)$.
-Dividing by $T$ again gives a rate of $\mathcal{O}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(\frac{1+P_{T}}{T}}\right)^{1/4}}\right)$ for the second term.
-Finally, we obtain by the same reasoning that
-$\sum_{t\in[T]}\alpha_{t}\leqslant(1+P_{T})^{1/4}\sum_{t\in[T]}t^{-1/4}=\mathcal{O}((1+P_{T})^{1/4}T^{3/4})\;,$
-and dividing by $T>0$ yields the desired rate. ∎
-F.2 Proofs of Section 4.2.
-Lemma 6.
-Assume H 1, $\mathcal{W}=\mathcal{W}_{1}$ and $\mathcal{R}=\mathcal{R}_{1}$.
-Then for any $t\in[T]$,$\tilde{f}_{t}$ is differentiable and for any $\theta\in\Theta$, $\nabla\tilde{f}_{t}(\theta)=\nabla\tilde{w}_{t}(\theta)\bar{g}_{t}(X_{t})$ where:
+Assume H˜1, $\mathcal{W}=\mathcal{W}_{1}$ and $\mathcal{R}=\mathcal{R}_{1}$.
+Then for any $t\in[T]$$\tilde{f}_{t}$ is differentiable and for any $\theta\in\Theta$, $\nabla\tilde{f}_{t}(\theta)=\nabla\tilde{w}_{t}(\theta)\bar{g}_{t}(X_{t})$ where:
 $\nabla\tilde{w}_{t}(\theta)=-\alpha_{t}^{-1}\mathopen{}\mathclose{{\left(\sum_{i=1}^{n}(b_{i}-A_{i}^{\mathsf{\scriptscriptstyle T}}\tilde{w}_{t}(\theta))^{-2}A_{i}A_{i}^{\mathsf{\scriptscriptstyle T}}}}\right)\nabla_{\theta}g(\theta,X_{t})\;.$
@@
 Lemma 5.
-Assume H 1 and $\mathcal{R}=\mathcal{R}_{1}$. For any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}$-Lipschitz almost-surely, with
+Assume H˜1 and $\mathcal{R}=\mathcal{R}_{1}$. For any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}$-Lipschitz almost-surely, with
 $K_{t}=\frac{GD_{\mathcal{Z}}}{\alpha_{t}}\frac{\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{\left\|b-Aw}}\right\|_{\infty}}{\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}\;.$
@@
 Lemma 6.
-Assume H 2, $\mathcal{W}=\mathcal{W}_{1}$ and $\mathcal{R}=\mathcal{R}_{1}$. Then for any $t\in[T]$ and $\theta\in\Theta$,
+Assume H˜2, $\mathcal{W}=\mathcal{W}_{1}$ and $\mathcal{R}=\mathcal{R}_{1}$. Then for any $t\in[T]$ and $\theta\in\Theta$,
 $\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$
@@
 $\tilde{w}_{t}(\theta)=\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),w}}\right\rangle-\alpha_{t}\sum_{i=1}^{d}\ln(b_{i}-A_{i}w)\;,$ (26)
-which is uniquely determined by strong convexity of the objective. We know by [15, page 566] that:
+which is uniquely determined by strong convexity of the objective. We know by [14, page 566] that:
 $\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle-\inf_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),w}}\right\rangle\leqslant n\alpha_{t}\;,$
@@
 $\leqslant\frac{2n\alpha_{t}}{\varepsilon}\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\;.$
-The result follows from combining H 2 and Lemma 9 according to the same lines of computation as in the proof of Lemma 3. ∎
+The result follows from combining H˜2 and Lemma˜7 according to the same lines of computation as in the proof of Lemma˜2. ∎
+G.1.3 Other lemmas
+Lemma 8.
+Assume that for any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}$-Lipschitz almost-surely and that the sequence of steps $(\eta_{t})_{t\geqslant 1}$ is non-increasing.
+Denote $\tilde{\mathfrak{R}^{d}_{T}}=\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})$.
+Then it holds almost surely that:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\tilde{\mathfrak{R}^{d}}_{T}}}\right]\leqslant\frac{1}{2\eta_{T}}\mathopen{}\mathclose{{\left(D_{\Theta}^{2}+2D_{\Theta}\sum_{t=1}^{T}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|}}\right)+\sum_{t=1}^{T}\frac{\eta_{t}}{2}K_{t}^{2}\;,$
+where the expectation is taken over the sequence $(u_{1},\ldots,u_{T})$ in Algorithm˜2.
+Proof.
+In what follows, we write $\mathbb{E}_{u_{t}}$ the expectation under the distribution $\text{Unif}([0,1])$ of $u_{t}$, and $\mathbb{E}_{u_{1:T}}$ the expectation under the joint distribution $\text{Unif}([0,1])^{\otimes T}$ of $(u_{1},\ldots,u_{T})$.
+For any $t\in[T]$, we have:
+$\sum_{t=1}^{T}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})$ $=\sum_{t=1}\int_{u=0}^{1}\mathopen{}\mathclose{{\left\langle\nabla\tilde{f}_{t}(\vartheta_{t}+u(\theta_{t}-\vartheta_{t})),\theta_{t}-\vartheta_{t}}}\right\rangle du$
+$=\sum_{t=1}^{T}\mathopen{}\mathclose{{\left\langle\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{\left[\nabla\tilde{f}_{t}(\vartheta_{t}+u_{t}(\theta_{t}-\vartheta_{t}))}}\right],\theta_{t}-\vartheta_{t}}}\right\rangle$
+$=\sum_{t=1}^{T}\mathopen{}\mathclose{{\left\langle\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{\left[\tilde{\nabla}_{t}(u_{t})}}\right],\theta_{t}-\vartheta_{t}}}\right\rangle\;,$ (31)
+where $\tilde{\nabla}(u_{t})=\nabla\tilde{f}_{t}(\vartheta_{t}+u_{t}(\theta_{t}-\vartheta_{t})$ as in Algorithm˜2. To control (31) we note that by definition of $\theta_{t+1}$ in Algorithm˜2,
+$\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t}}}\right\|^{2}}}\right]$ $=\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\Pi_{\Theta}[\theta_{t}-\eta_{t}\tilde{\nabla}_{t}(u_{t})]-\vartheta_{t}}}\right\|^{2}}}\right]$
+$\leqslant\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\theta_{t}-\vartheta_{t}-\eta_{t}\tilde{\nabla}_{t}(u_{t})}}\right\|^{2}}}\right]$
+$\leqslant\mathopen{}\mathclose{{\left\|\theta_{t}-\vartheta_{t}}}\right\|^{2}-2\eta_{t}\mathopen{}\mathclose{{\left\langle\mathbb{E}_{u_{t}}\mathopen{}\mathclose{{\left[\tilde{\nabla}_{t}(u_{t})}}\right],\theta_{t}-\vartheta_{t}}}\right\rangle+\eta_{t}^{2}K_{t}^{2}$ (32)
+where we have used that $\mathopen{}\mathclose{{\left\|\tilde{\nabla}_{t}(u_{t})}}\right\|^{2}\leqslant K_{t}^{2}$ because $\tilde{f}_{t}$ is $K_{t}$-Lipschitz.
+Re-arranging (32) and taking the expectation over $(u_{1},\ldots,u_{T})$ yields:
+$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\tilde{\nabla}_{t}(u_{t}),\theta_{t}-\vartheta_{t}}}\right\rangle}}\right]$ $\leqslant\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{\left[\frac{1}{2\eta_{t}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\|\theta_{t}-\vartheta_{t}}}\right\|^{2}-\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t}}}\right\|^{2}}}\right)+\frac{\eta_{t}K_{t}^{2}}{2}}}\right]\;,$
+Now notice that, for any $t\in\{1,\ldots,T-1\}$,
+$\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t+1}}}\right\|^{2}$ $=\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t}}}\right\|^{2}+\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t+1}}}\right\|^{2}-\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t}}}\right\|^{2}$
+$=\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t}}}\right\|^{2}$
+$+(\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t+1}}}\right\|+\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t}}}\right\|)(\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t+1}}}\right\|-\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t}}}\right\|)$
+$\leqslant\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t}}}\right\|^{2}+2D_{\Theta}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|\;,$ (34)
+where we have used H˜1-(i) and the reversed triangular inequality in the last line. It follows that:
+$-\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t}}}\right\|^{2}$ $\leqslant-\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t+1}}}\right\|^{2}+2D_{\Theta}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|\$(35)
+Plugging (35) into (G.1.3) then gives:
+$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\tilde{\nabla}_{t}(u_{t}),\theta_{t}-\vartheta_{t}}}\right\rangle}}\right]$ $\leqslant\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{\left[\frac{1}{2\eta_{t}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\|\theta_{t}-\vartheta_{t}}}\right\|^{2}-\mathopen{}\mathclose{{\left\|\theta_{t+1}-\vartheta_{t+1}}}\right\|^{2}}}\right)}}\right$.
+$\quad\mathopen{}\mathclose{{\left.+\frac{1}{\eta_{t}}D_{\theta}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|+\frac{\eta_{t}K_{t}^{2}}{2}}}\right]\;.$
+Thus, summing over $T$ and plugging the resulting sum in (31) gives:
+$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{\left[\sum_{t=1}^{T}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}}\right]$ $\leqslant\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{\left[\sum_{t=1}^{T}\frac{\mathopen{}\mathclose{{\left\|\theta_{t}-\vartheta_{t}}}\right\|^{2}}{2}\mathopen{}\mathclose{{\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}}}\right)}}\right$.
+$\quad\mathopen{}\mathclose{{\left.+\sum_{t=1}^{T}\frac{1}{\eta_{t}}D_{\theta}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|+\sum_{t=1}^{T}\frac{\eta_{t}K_{t}^{2}}{2}}}\right]\;.$
+$\leqslant\frac{D_{\Theta}^{2}}{2}\sum_{t=1}^{T}\mathopen{}\mathclose{{\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}}}\right)+\frac{D_{\theta}}{\eta_{T}}\sum_{t=1}^{T}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|+\sum_{t=1}^{T}\frac{\eta_{t}K_{t}^{2}}{2}\;.$
+where we used that for all $t$, $\mathopen{}\mathclose{{\left\|\theta_{t}-\vartheta_{t}}}\right\|\leqslant D_{\Theta}$ by H˜1-(i) and that for all $t\in[T],\,\eta_{t-1}\geqslant\eta_{t}\geqslant\eta_{T}$.
+Then, we have by telescoping the first sum:
+$\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{\left[\sum_{t=1}^{T}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+2D_{\Theta}}{2\eta_{T}}\sum_{t=1}^{T}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|+\sum_{t=1}^{T}\frac{\eta_{t}K_{t}^{2}}{2}.$
+G.2 Proof of Theorem˜1.
+See 1
+Proof.
+For the sake of conciseness we use the notation $\mathbb{E}_{t}$ to denote $\mathbb{E}\mathopen{}\mathclose{{\left[\cdot\mid\mathcal{H}_{t-1}}}\right]$ In what follows, we consider, for any $\theta$, the intermediary regret: $\mathrm{R}_{T}^{s}(\theta):=\sum_{t=1}^{T}F_{t}(\theta_{t})-F_{t}(\theta)$, where $\quad F_{t}:\theta\mapsto\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}}\right]$.
+We have, for any $\theta$, the following decomposition of the static regret
+$\mathrm{R}_{T}^{s}(\theta)$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}}\right\rangle}}\right]$ (36)
+$+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}}\right\rangle}}\right]-\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$ (37)
+$+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]-\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}}\right\rangle}}\right]$ (38)
+Then, taking the $\sup$ over $\Theta$ in Equations˜36 and 38, and defining
+$\mathrm{Reg}^{\mathrm{S}}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}}\right\rangle}}\right]-\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right],$
+in Equation˜37 yields for any $\theta$:
+$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$
+[…]
+$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]$
+Since for any $t\in[T]$, $\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{2}\leqslant D_{\mathcal{Z}}$ by H˜1-(ii),
+$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\mathbb{E}_{t}{\sup_{\theta\in\Theta}\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}\;$
+First, notice that, by Lemma˜6, and because for any $t,\alpha_{t}=\alpha$ we have for all $t$:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$ (39)
+Second, taking the expectation (denoted by $\mathbb{E}$) over $(\sigma_{t})_{t\in[T]}$ and $X_{1},\cdots X_{T}$ on both sides gives:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]+2D_{\mathcal{Z}}\alpha T\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$ (40)
+To bound the first term in the right-hand-side of Equation˜40. by the definition of conditional expectation, we have:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]$ $=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}}\right\rangle-\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$
+Then,
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]$ $\leqslant\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}}\right].$
+One recognises (up to a factor $T$) the left-hand side of Proposition˜4 on the loss sequence $(\tilde{f}_{t})_{t\in[T]}$.
+Furthermore, we can use this proposition as, given our choice of $\mathcal{R}$, for any $t\in[T]$, $\tilde{f}_{t}$ is $L$-Lipschitz with $L=5D_{\mathcal{Z}}G(4\alpha)^{-1}$ almost surely by Lemma˜1 (with $\chi=0)$.
+We then have:
+$\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{\left[\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}}\right]\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi}}\right)\;.$ (41)
+Dividing Equation˜40 by $T$, and plugging Equation˜41 gives: for all $\theta\in\Theta$:
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]$ $\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi}}\right)$
+$+2D_{\mathcal{Z}}\alpha\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right),$
+$=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi+\alpha n}}\right).$
+Finally remark that, by the definition of the conditional expectation (thus of $F_{t}$), and because $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, we have for any $\theta$:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]$ $=\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{t=1}^{T}f_{t}(\theta_{t})-f_{t}(\theta)}}\right]$
+$=\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{t=1}^{T}f_{t}(\theta_{t})}}\right]-\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)}}\right]$
+Then taking the infimum over $\theta$ yields:
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $\leqslant\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi+\alpha n}}\right).$
+This concludes the proof. The second equation consists in simply plugging the proposed value of $\eta,\alpha$ in this bound. ∎
+G.3 Proof of Theorem˜2
 See 2
@@
 In this proof, for the sake of conciseness, we use the notation $\mathbb{E}_{t}$ to denote $\mathbb{E}\mathopen{}\mathclose{{\left[\cdot\mid\mathcal{H}_{t-1}}}\right]$.
-We showed in (34) of the proof of Proposition 1 that if we define $\mathrm{Reg}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle}\right]$, then we have:
-$\mathfrak{R}_{T}$ $\leqslant\mathrm{Reg}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]+\xi T.$
-First, by Lemma 8, we know have for any $t$:
+Observe that the dynamic regret can be decomposed as follows:
+$\mathfrak{R}^{d}_{T}$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}}\right\rangle}}\right]+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]$ (42)
+$+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]-\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$ (43)
+$+\sum_{t\in[T]}\mathopen{}\mathclose{{\left[\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]-\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}}\right\rangle}}\right]}}\right]$ (44)
+First, we remark that for any $t$, given $\vartheta_{t}=\mathbf{O}(\tilde{f}_{t})$, we control (43) by Jensen:
+$\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]-\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$ $\leqslant\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\tilde{f}_{t}(\vartheta_{t})-\inf_{\theta\in\Theta}\tilde{f}_{t}(\theta)}}\right]$
+$\leqslant\xi.$
+Then, taking the $\sup$ over $\Theta$ for each summand of the first sum of (42) (which is valid as $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable), defining by $\mathrm{Reg}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]$, and noticing for (44) that $\inf_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)]-\inf_{\theta\in\Theta}\mathbb{E}_{t}[\tilde{f}_{t}(\theta)]\leqslant\sup_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)-\tilde{f}_{t}(\theta)]$ leads to:
+$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right]+\xi T$
+[…]
+$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]+\xi T$
+Since for any $t\in[T]$, $\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{2}\leqslant D_{\mathcal{Z}}$ almost surely by H˜1-(ii),
+$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]+\xi T\;$ (45)
+First, by Lemma˜6, we know have for any $t$:
 $\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$
 Thus combining the two last equations and taking the expectation (denoted by $\mathbb{E}$) over $(u_{t})_{t\in[T]}$ and $X_{1},\cdots X_{T}$ on both sides gives:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\mathrm{Reg}_{T}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}\mathopen{}\mathclose{{}\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)+\xi T\;.$ (45)
-Just as in the proof of Proposition 1, we can bound the first term in Equation 45 by noting that the definition of conditional expectation implies:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}_{T}}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)+\xi T\;.$ (46)
+Now, to bound the first term in Equation˜46 note that the definition of conditional expectation implies:
 $\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}_{T}}}\right]=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}}\right].$
-One recognizes the definition of $\tilde{\mathfrak{R}_{T}}$ of Lemma 5.
-Since $\tilde{f}_{t}$ is $K_{t}-$Lipschitz with $K_{t}=\frac{GD_{\mathcal{Z}}}{\alpha_{t}}\frac{\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}}{\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}$ almost surely for any $t$ by Lemma 7, we deduce from Lemma 5 that:
+One recognizes the definition of $\tilde{\mathfrak{R}^{d}_{T}}$ of Lemma˜8.
+Since $\tilde{f}_{t}$ is $K_{t}$-Lipschitz with $K_{t}=\frac{GD_{\mathcal{Z}}}{\alpha_{t}}\frac{\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{\left\|b-Aw}}\right\|_{\infty}}{\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}$ almost surely for any $t$ by Lemma˜5, we deduce from Lemma˜8 that:
 $\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}}\right]\leqslant\frac{D_{\Theta}^{2}-D_{\Theta}P_{T}}{2\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{\left\|b-Aw}}\right\|_{\infty}^{2}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}}\right)^{2}\;.$ (47)
-Plugging Equation 46 into Equation 45 and then dividing by $T>0$ on both sides gives the desired result. ∎
+Plugging Equation˜47 into Equation˜46 and then dividing by $T>0$ on both sides gives:
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{\left[P_{T}}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{\left\|b-Aw}}\right\|_{\infty}^{2}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}}\right)^{2}$
+$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}(\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d))]+\xi\;.$ (48)
+Rewriting Section˜G.3 using $\tilde{\mathcal{O}}$ concludes the proof.
+Concerning the second bound, We know by Section˜G.3 that the regret of Algorithm˜2 satisfies the following bound:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\mathbb{E}\mathopen{}\mathclose{{\left[\frac{1+P_{T}}{\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{\alpha_{t}^{2}}+n\sum_{t\in[T]}\alpha_{t}+\xi T}}\right]}}\right)\;.$
+With $\alpha_{t}\propto n^{-1/2}t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto\alpha_{t}t^{-1/2}(1+P_{t})^{1/2}\propto n^{-1/2}t^{-3/4}(1+P_{t})^{3/4}$, the sequence $(\eta_{t})_{t\in[T]}$ is non-increasing because so is $(t^{-1}(1+P_{t}))_{t\in[T]}$ by assumption.
+Moreover, the first term satisfies:
+$\frac{1+P_{T}}{\eta_{T}}=\mathcal{O}\mathopen{}\mathclose{{\left(n^{1/2}\frac{T^{3/4}}{(1+P_{T})^{3/4}}\mathopen{}\mathclose{{\left(1+P_{T}}}\right)}}\right)\quad\text{so}\quad\frac{1+P_{T}}{T\eta_{T}}=\mathcal{O}\mathopen{}\mathclose{{\left(n^{1/2}\mathopen{}\mathclose{{\left(\frac{1+P_{T}}{T}}}\right)^{1/4}}}\right)\;.$
+For the second term,
+$\sum_{t=1}^{T}\frac{\eta_{t}}{\alpha_{t}^{2}}=\sum_{t=1}^{T}\frac{\sqrt{1+P_{t}}}{\sqrt{t}\alpha_{t}}=\mathcal{O}\mathopen{}\mathclose{{\left(n^{1/2}\sum_{t=1}^{T}\mathopen{}\mathclose{{\left(\frac{1+P_{t}}{t}}}\right)^{1/4}}}\right)\;..$
+Moreover, since $1+P_{t}\leqslant 1+P_{T}$ for any $t\in[T]$, we have:
+$\sum_{t=1}^{T}\frac{\eta_{t}}{\alpha_{t}^{2}}=\mathcal{O}\mathopen{}\mathclose{{\left(n^{1/2}\mathopen{}\mathclose{{\left(1+P_{T}}}\right)^{1/4}\sum_{t=1}^{T}\mathopen{}\mathclose{{\left(\frac{1}{t}}}\right)^{1/4}}}\right)=\mathcal{O}\mathopen{}\mathclose{{\left(n^{1/2}\mathopen{}\mathclose{{\left(1+P_{T}}}\right)^{1/4}T^{3/4}}}\right)\;,$
+where we used in the last line that $\sum_{t=1}^{T}t^{-\gamma}=\mathcal{O}\mathopen{}\mathclose{{\left(T^{1-\gamma}}}\right)$ for $\gamma\in(0,1)$.
+Dividing by $T$ again gives a rate of $\mathcal{O}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left(\frac{1+P_{T}}{T}}}\right)^{1/4}}}\right)$ for the second term.
+Finally, we obtain by the same reasoning that
+$n\sum_{t\in[T]}\alpha_{t}\leqslant n^{1/2}(1+P_{T})^{1/4}\sum_{t\in[T]}t^{-1/4}=\mathcal{O}\mathopen{}\mathclose{{\left(n^{1/2}(1+P_{T})^{1/4}T^{3/4}}}\right)\;,$
+and dividing by $T>0$ yields the desired rate. ∎
+G.4 Proof of Proposition˜2
 See 2
 Proof.
-We know by Proposition 2 that the regret of Algorithm 1 satisfies the following asymptotic bound:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{1+P_{T}}{\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{\alpha_{t}^{2}}+n\sum_{t\in[T]}\alpha_{t}+\xi T}\right]}\right)\;.$
-The rest of the proof proceeds exactly as the one of Corollary 1. ∎
-F.3 Proofs of Appendix B.
-See 4
-Proof.
-For the sake of conciseness we use the noation $\mathbb{E}_{t}$ to denote $\mathbb{E}\mathopen{}\mathclose{{}\left[\cdot\mid\mathcal{H}_{t-1}}\right]$ We have, for any $\theta$, the following decomposition of the static regret
-$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}\right\rangle}\right]$ (47)
+For the sake of conciseness we use the notation $\mathbb{E}_{t}$ to denote $\mathbb{E}\mathopen{}\mathclose{{\left[\cdot\mid\mathcal{H}_{t-1}}}\right]$ In what follows, we consider, for any $\theta$, the intermediary regret: $\mathrm{R}_{T}^{s}(\theta):=\sum_{t=1}^{T}F_{t}(\theta_{t})-F_{t}(\theta)$, where $\quad F_{t}:\theta\mapsto\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}}\right]$.
+We have, for any $\theta$, the following decomposition of the static regret
+$\mathrm{R}_{T}^{s}(\theta)$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}}\right\rangle}}\right]$ (49)
 $+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}}\right\rangle}}\right]-\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$ (50)
 $+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]-\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}}\right\rangle}}\right]$ (51)
-Then, taking the $\sup$ over $\Theta$ in Equations 47 and 49, and defining
+Then, taking the $\sup$ over $\Theta$ in Equations˜49 and 51, and defining
 $\mathrm{Reg}^{\mathrm{S}}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}}\right\rangle}}\right]-\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right],$
-in Equation 48 yields for any $\theta$:
-$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\rangle}\right]$
-Using the fact that $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{\infty}$ is dual to $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{1}$:
-$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]$
-Since for any $t\in[T]$, $\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{2}\leqslant D_{\mathcal{Z}}$ by H 1-(ii),
-$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}\;$
-First, notice that, by Lemma 3, and because for any $t,\alpha_{t}=\alpha$ we have for all $t$:
-$\mathbb{E}_{X_{1},\cdots,X_{T}}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]$ $\leqslant 2D_{\mathcal{Z}}\alpha\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}\right)}\right)\;.$ (50)
+in Equation˜50 yields for any $\theta$:
+$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$
+[…]
+$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]$
+Since for any $t\in[T]$, $\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{2}\leqslant D_{\mathcal{Z}}$ by H˜1-(ii),
+$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\mathbb{E}_{t}{\sup_{\theta\in\Theta}\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}\;$
+First, notice that, by Lemma˜2, and because for any $t,\alpha_{t}=\alpha$ we have for all $t$:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right)\;.$ (52)
 Second, taking the expectation (denoted by $\mathbb{E}$) over $(\sigma_{t})_{t\in[T]}$ and $X_{1},\cdots X_{T}$ on both sides gives:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}^{\mathrm{S}}_{T}(\theta)}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\mathrm{Reg}^{\mathrm{S}}_{T}}\right]+2D_{\mathcal{Z}}\alpha T\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}\right)}\right).$ (51)
-To bound the first term in the right-hand-side of Equation 51. by the definition of conditional expectation, we have:
-$\mathrm{Reg}^{\mathrm{S}}_{T}$ $=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}\right\rangle-\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle}\right]$
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]+2D_{\mathcal{Z}}\alpha T\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right)\;.$ (53)
+To bound the first term in the right-hand-side of Equation˜53. by the definition of conditional expectation, we have:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]$ $=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}}\right\rangle-\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$
 Then,
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}^{\mathrm{S}}_{T}}\right]$ $\leqslant\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}\right].$
-One recognises (up to a factor $T$) the left-hand side of Proposition 3 on the loss sequence $(\tilde{f}_{t})_{t\in[T]}$.
-Furthermore, we can use this proposition as, given our choice of $\mathcal{R}$, for any $t\in[T]$, $\tilde{f}_{t}$ is $L-$Lipschitz with $L=5D_{\mathcal{Z}}G(4\alpha)^{-1}$ almost surely by Lemma 2.
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]$ $\leqslant\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}}\right].$
+One recognises (up to a factor $T$) the left-hand side of Proposition˜4 on the loss sequence $(\tilde{f}_{t})_{t\in[T]}$.
+Furthermore, we can use this proposition as, given our choice of $\mathcal{R}$, for any $t\in[T]$, $\tilde{f}_{t}$ is $L$-Lipschitz with $L=5D_{\mathcal{Z}}G(4\alpha)^{-1}$ almost surely by Lemma˜1 (with $\chi=0)$.
 We then have:
-$\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{}\left[\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}\right]\leqslant\mathcal{O}\mathopen{}\mathclose{{}\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi m\frac{1}{\alpha}}\right)\;.$ (52)
-Dividing Equation 51 by $T$, and plugging Equation 52 into it concludes the proof. ∎
-F.4 Proofs of Appendix D.
-Proof of Lemma 1.
-The proof starts exactly in the same way as the one of Lemma 3, where proved in (24) that:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\mid\mathcal{H}_{t-1}}\right]\\ =\alpha_{t}+2\alpha_{t}\ln(d)\int_{(\alpha_{t}\ln(d))}^{(2\ln(d))}\mathbb{P}\mathopen{}\mathclose{{}\left(\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>2\frac{\alpha_{t}\ln(d)}{\varepsilon}\,\mid\,\mathcal{H}_{t-1}}\right)\frac{d\varepsilon}{\varepsilon^{2}}.$ (53)
-Moreover, we have for any $\varepsilon>0$:
-$1-C_{0}\varepsilon^{\beta}$ $\leqslant\mathbb{P}\mathopen{}\mathclose{{}\left(\forall\theta\in\Theta,\;\inf_{j\neq I_{t}}u_{j}(\theta,X_{t})-u_{I_{t}}(\theta,X_{t})\geqslant\varepsilon\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right)$
-$\leqslant\mathbb{P}\mathopen{}\mathclose{{}\left(\forall\theta\in\Theta,\;\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}{\mathcal{H}_{t-1}}\leqslant 2\alpha_{t}\ln(d)\varepsilon^{-1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right)$
-$\leqslant\mathbb{P}\mathopen{}\mathclose{{}\left(\;\sup_{\theta\in\Theta}\;\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}\leqslant 2\alpha_{t}\ln(d)\varepsilon^{-1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right)$
-where the first inequality stems from H 4, and the second one from Lemma 4. It follows that:
-$\mathbb{P}\mathopen{}\mathclose{{}\left(\sup_{\theta\in\Theta}\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>2\frac{\alpha_{t}\ln(d)}{\varepsilon}\mid\mathcal{H}_{t-1}}\right)$ $=1-\mathbb{P}\mathopen{}\mathclose{{}\left(\sup_{\theta\in\Theta}\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\leqslant 2\frac{\alpha_{t}\ln(d)}{\varepsilon}\mid\mathcal{H}_{t-1}}\right)$
-$\leqslant C_{0}\varepsilon^{\beta}\;,$ (54)
-Hence, plugging (54) in (24) gives:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right]$ $\leqslant\alpha_{t}+2\alpha_{t}\ln(d)C_{0}\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon\;.$
-And the result is established by bounding $\varphi(\beta)=\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon$ in the same way as in the proof of Lemma 3. ∎
-Proof of Proposition 5.
-Observe that the pseudo dynamic regret can be decomposed as follows:
-$\mathfrak{R}_{T}^{\mathrm{d}}$ $=\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}\right\rangle+\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle$ (55)
-$+\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\vartheta_{t})}\right\rangle-\inf_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle$ (56)
-$+\sum_{t\in[T]}\mathopen{}\mathclose{{}\left[\inf_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle-\inf_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}\right\rangle}\right]$ (57)
-Taking the $\sup$ over $\Theta$ and defining by ~$\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}:=\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle$ in (55), using H 3 in (56), and noticing that $\inf_{\theta\in\Theta}f_{t}(\theta)-\inf_{\theta\in\Theta}\tilde{f}_{t}(\theta)\leqslant\sup_{\theta\in\Theta}f_{t}(\theta)-\tilde{f}_{t}(\theta)$ in (57) leads to:
-$\mathfrak{R}_{T}^{\mathrm{d}}$$\leqslant\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\rangle+\xi T$
-Using the fact that $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{\infty}$ is dual to $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{1}$:
-$\mathfrak{R}_{T}^{\mathrm{d}}$$\leqslant\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}+\xi T$
-Observing that $\lVert\bar{g}_{t}(X_{t})\rVert_{\infty}\leqslant\lVert\bar{g}_{t}(X_{t})\rVert_{2}\leqslant D_{\mathcal{Z}}$ almost surely by H 1-(ii), and taking the expectation over $X_{1},\ldots,X_{T}$ and $u_{1},\ldots,u_{T}$ then gives:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right]$ $\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}}\right]+2D_{\mathcal{Z}}\sum_{t\in[T]}\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]+\xi T\;$ (58)
-We now proceed to bound the two first terms in (58). First, notice that by Lemma 1:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]$ $\leqslant\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)\;.$ (59)
-Second, noticing that $\tilde{f}_{t}$ is $K_{t}-$Lipschitz with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$ almost surely by Lemma 2, we have by Lemma 5:
-$\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}}\right]\leqslant\frac{1}{2\eta_{T}}(D_{\Theta}^{2}+D_{\Theta}P_{T})+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}^{2}G^{2}\eta_{t}}{32\alpha_{t}^{2}}\;.$ (60)
-Plugging (59) and (60) in (58) and dividing by $T$ on both sides gives the desired result. ∎
-F.5 Proofs of Appendix E
-Proof of Proposition 6.
+$\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{\left[\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}}\right]\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi}}\right)\;.$ (54)
+Dividing Equation˜53 by $T$, and plugging Equation˜54 gives: for all $\theta\in\Theta$:
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]$ $\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi}}\right)$
+$+2D_{\mathcal{Z}}\alpha\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right),$
+$=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi+\alpha\ln(d)}}\right).$
+Finally remark that, by the definition of the conditional expectation (thus of $F_{t}$), and because $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, we have for any $\theta$:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]$ $=\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{t=1}^{T}f_{t}(\theta_{t})-f_{t}(\theta)}}\right]$
+$=\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{t=1}^{T}f_{t}(\theta_{t})}}\right]-\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)}}\right]$
+Then taking the infimum over $\theta$ yields:
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $\leqslant\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi+\alpha\ln(d)}}\right).$
+This concludes the proof. The second equation consists in simply plugging the proposed value of $\eta,\alpha$ in this bound. ∎
+G.5 Proof of Proposition˜3
+See 3
+Proof.
+In this proof, for the sake of conciseness, we denote by $\mathbb{E}_{t}$ the conditional expectation $\mathbb{E}\mathopen{}\mathclose{{\left[\cdot\mid\mathcal{H}_{t-1}}}\right]$.
+Observe that the dynamic regret can be decomposed as follows:
+$\mathfrak{R}^{d}_{T}$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}}\right\rangle}}\right]+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]$ (55)
+$+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]-\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$ (56)
+$+\sum_{t\in[T]}\mathopen{}\mathclose{{\left[\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]-\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}}\right\rangle}}\right]}}\right]$ (57)
+First, we remark that for any $t$, given $\vartheta_{t}=\mathbf{O}(\tilde{f}_{t})$, we control (56) by Jensen:
+$\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]-\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$ $\leqslant\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\tilde{f}_{t}(\vartheta_{t})-\inf_{\theta\in\Theta}\tilde{f}_{t}(\theta)}}\right]$
+$\leqslant\xi.$
+Then, taking the $\sup$ over $\Theta$ for each summand of the first sum of (55) (which is valid as $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable), defining by $\mathrm{Reg}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]$, and noticing for (57) that $\inf_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)]-\inf_{\theta\in\Theta}\mathbb{E}_{t}[\tilde{f}_{t}(\theta)]\leqslant\sup_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)-\tilde{f}_{t}(\theta)]$ leads to:
+$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right]+\xi T$
+[…]
+$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]+\xi T$
+Since for any $t\in[T]$, $\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{2}\leqslant D_{\mathcal{Z}}$ almost surely by H˜1-(ii),
+$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]+\xi T\;$ (58)
+First, notice that, by Lemma˜2, for all $t$:
+$\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]$ $\leqslant\alpha_{t}\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)}}\right)\;.$ (59)
+Second, taking the expectation (denoted by $\mathbb{E}$) over $(u_{t})_{t\in[T]}$ and $X_{1},\cdots X_{T}$ on both sides gives:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}_{T}}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)}}\right)+\xi T\;.$ (60)
+To bound the first term in the right-hand side of Equation˜60. Remark that, by the definition of conditional expectation:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}_{T}}}\right]=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}}\right].$
+One recognises the definition of $\tilde{\mathfrak{R}^{d}_{T}}$ of Lemma˜8.
+Thus, by this lemma, we know that for any $t\in[T]$, because $\tilde{f}_{t}$ is $K_{t}$-Lipschitz with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$ almost surely by Lemma˜1, we have:
+$\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}}\right]\leqslant\frac{1}{2\eta_{T}}(D_{\Theta}^{2}+D_{\Theta}P_{T})+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}^{2}G^{2}\eta_{t}}{32\alpha_{t}^{2}}\;.$ (61)
+Plugging Equation˜61 into Equation˜60 and then dividing by $T>0$ on both sides gives the desired result. ∎
+G.6 Proofs of Appendix˜F
+Proof of Proposition˜5.
 The regret simply decomposes as follows:
-$\mathfrak{R}_{T}$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\underline{w}_{t}(\theta_{t})}\right\rangle}\right]-\sum_{t\in[T]}\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}\right\rangle}\right]$
+$\mathfrak{R}^{d}_{T}$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\underline{w}_{t}(\theta_{t})}}\right\rangle}}\right]-\sum_{t\in[T]}\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}}\right\rangle}}\right]$
 $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\underline{w}_{t}(\theta_{t})-w^{\star}_{t}(\theta_{t})}}\right\rangle}}\right]$
@@
 $\leqslant\kappa+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})}}\right\rangle}}\right]-\sum_{t\in[T]}\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}}\right\rangle}}\right]\;,$
-and the proof continues as in Proposition 2. ∎
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: [Yes] Yes, the main claims in the abstract and introduction accurately reflect the paper’s contributions and scope.
-They clearly summarize the key methods, findings, and the intended impact, aligning well with the content presented in the body of the paper.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9.
-Justification: All the assumptions underlying the analysis are highlighted and clearly discussed (in particular H 2, which has a dedicated discussion appendix in Appendix D).
-The scope of our results is also extensively discussed, see the comments after Proposition 1, Proposition 2, Corollary 1, Corollary 2.
-Likewise, the experimental setting is clearly described.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14.
-Justification: We checked the correctness of all our proofs.
-The assumptions are systematically recalled in the propositions and lemmas, see the statements of Proposition 1, Proposition 2, Corollary 1, Corollary 2.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: The experimental setting is extensively described in discussed in Appendix C.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: All our experiments are easily reproducible, as our experimental setting is precisely described in Appendix C.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: Yes, see Appendix C.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes] .
-  34. Justification: see the plots in Section 5.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: Our experiments are particularly light and can be run on any personal computer in reasonable time.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We made sure that the research conducted in the paper conforms in every aspcet with the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49.
-Justification: We discuss in the introduction (Section 1) the potential impact of online Decision-Focused learning in several fields of Operation Research, including logistics, healthcare operation or supply chain management.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: There is no need for such safeguards as we only use simple, synthetic data for our experiments.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We explicitly cite all the papers upon which the current work builds.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: Our algorithm is thoroughly explained in the main text of this paper (see Section 3).
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: We do not have any crowdsourcing experiment or research with human subjects in this paper.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: Same answer as before.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: No LLMs were used in this study.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+and the proof continues as in Theorem˜2. ∎

### v3

@@
 We instead investigate DFL in dynamic environments where the objective function and data distribution evolve over time.
-This setting is challenging for online learning because the objective function has zero or undefined gradients—which prevents the use of standard first-order optimization methods—and is generally non-convex.
+This setting is challenging for online learning because the objective function has zero or undefined gradients, which prevents the use of standard first-order optimization methods, and is generally non-convex.
 To address these difficulties, we (i) regularize the objective to make it differentiable and (ii) use perturbation techniques along with a near-optimal oracle to overcome non-convexity.
@@
 Finally, we showcase the effectiveness of our algorithms on a knapsack experiment, where they outperform two standard benchmarks.
-1 Introduction.
-Many real-world decision problems involve uncertainty, and a common approach to handling it is through the predict-then-optimize framework [10].
+1 Introduction
+Many real-world decision problems involve uncertainty, which is commonly handled through the predict-then-optimize framework (Bertsimas and Kallus, 2020).
 First, a prediction model is trained on historical data; then, its output is fed into an optimization problem to guide decision-making.
-This natural strategy has been successfully applied in many operation research (OR) problems, ranging from supply chain management [2, 28, 7, 10] and revenue management [26, 27, 18, 16] to healthcare operation [11, 5, 34, 61], see Mivsić and Perakis [56] for an extensive review.
+This natural strategy has been successfully applied in many operation research (OR) problems, ranging from supply chain management (Acimovic and Graves, 2015; Fisher et al., 2016; Ban and Rudin, 2019; Bertsimas and Kallus, 2020), revenue management (Farias et al., 2013; Ferreira et al., 2016; Cohen et al., 2017; Chen et al., 2022) to healthcare operation (Bertsimas et al., 2013; Aswani et al., 2019; Gupta et al., 2020; Rath et al., 2017), see Mivsić and Perakis (2020) for an extensive review.
 It is clear that this approach would yield optimal decisions if the predictions were perfectly accurate.
-However, in practice, prediction errors are inevitable—and even small inaccuracies can propagate through the optimization process, potentially leading to poor decisions.
-To address this limitation, an approach known as decision-focused learning [54]—also referred to as smart predict-then-optimize [25] or integrated learning-optimization [62]—has emerged.
+However, in practice, prediction errors are inevitable, and even small inaccuracies can propagate through the optimization process, potentially leading to poor decisions.
+To address this limitation, an approach known as decision-focused learning (Mandi et al., 2024), also referred to as smart predict-then-optimize (Elmachtoub and Grigas, 2022) or integrated learning-optimization (Sadana et al., 2025), has emerged.
 Instead of optimizing for prediction accuracy alone, this method trains the predictive model to directly minimize the downstream decision loss.
 By aligning the learning objective with the decision-making goal, it produces models that are more robust to prediction errors in practical applications.
-While decision-focused learning yields strong empirical performance [23, 71, 72, 73], theoretical development has so far been limited to the batch setting, where models are trained on pre-collected, independently and identically distributed (i.i.d.) data [74, 53, 64, 63].
-This assumption breaks down in many real-world scenarios involving dynamic environments [17, 57] and shifting data distributions [50, 60].
-Online learning [15, 37] provides a general way to cope with such non-stationarity of data-generating processes.
+While decision-focused learning yields strong empirical performance (Donti et al., 2017; Verma et al., 2022; 2023; Wang et al., 2023), theoretical development has so far been limited to the batch setting, where models are trained on pre-collected, independently and identically distributed (i.i.d.) data (Wilder et al., 2019; Mandi et al., 2022; Shah et al., 2022; Schutte et al., 2024).
+This assumption breaks down in many real-world scenarios involving dynamic environments (Cheung et al., 2019; Padakandla et al., 2020) and shifting data distributions (Lu et al., 2018; Quiñonero-Candela et al., 2022).
+Online learning (Cesa-Bianchi and Lugosi, 2006; Hazan, 2023) provides a general way to cope with such non-stationarity of data-generating processes.
 This framework considers a learner who makes decisions sequentially, at each round leveraging data collected from previous rounds to inform its next decision.
@@
 We develop a theoretical foundation for online decision-focused learning, enabling its application in non-stationary settings.
-This presents a significant technical challenge, as the inherent difficulties of decision-focused learning—such as the non-differentiability of the objective function or the lack of convexity of the losses due to the bi-level nature of the problem—compound those already present in online learning.
+This presents a significant technical challenge, as the inherent difficulties of decision-focused learning, such as the non-differentiability of the objective function or the lack of convexity of the losses due to the bi-level nature of the problem, compound those already present in online learning.
 Our contributions are as follows:
@@
 While both rely on regularizing the inner problem to make the resulting decision differentiable, they differ on the way they update the parameters of the prediction model.
-DF-FTPL uses the FTPL approach [43], and DF-OGD leverages a variant of Online Gradient Descent [81].
+DF-FTPL uses the FTPL approach (Hutter and Poland, 2005), and DF-OGD leverages a variant of Online Gradient Descent (Zinkevich, 2003).
 We establish sublinear convergence guarantees for both procedures, in the form of a static regret bound for the former and a dynamic regret bound for the latter.
@@
   3. (iii)
-    Finally, we assess the performance of our algorithms on a knapsack experiment inspired by [54].
+    Finally, we assess the performance of our algorithms on a knapsack experiment inspired by Mandi et al.
+(2024).
 Our simulations demonstrate that our approach outperforms the online version of two popular baselines, namely prediction-focused learning and Smart-Predict-then-Optimize, in both static and dynamic environments.
 Additional related work.
-Our work addresses the decision-focused learning problem, pioneered by [25].
-Several approaches have been proposed in the batch setting, among which differentiating the associated KKT conditions [33, 4, 23, 74, 52], smoothing via random perturbation [9], building surrogate losses via duality [25] or directional gradients [42] and relying on pairwise ranking techniques [53], see [54, 62] for additional references.
+Several decision-focused approaches have been proposed in the batch setting, among which differentiating the associated KKT conditions Gould et al.
+(2016); Amos and Kolter (2017); Donti et al.
+(2017); Wilder et al.
+(2019); Mandi and Guns (2020), smoothing via random perturbation Berthet et al.
+(2020), building surrogate losses via duality Elmachtoub and Grigas (2022) or directional gradients Huang and Gupta (2024) and relying on pairwise ranking techniques Mandi et al.
+(2022), see Mandi et al.
+(2024); Sadana et al.
+(2025) for additional references.
 However, the extension of decision-focused learning to the online setting remains unexplored.
-While several recent studies address online bi-level optimization [65, 69, 49], none of them are applicable to decision-focused learning as they rely on restrictive smoothness assumptions, that are incompatible with the structure of decision-focused problems.
+While several recent studies address online bi-level optimization Shen et al.
+(2023); Tarzanagh et al.
+(2024); Lin et al.
+(2023), none of them are applicable to decision-focused learning as they rely on restrictive smoothness assumptions, that are incompatible with the structure of decision-focused problems.
 In particular, the decision-focused objective is typically non-convex and features gradients that are either zero or undefined, owing to its underlying linear structure.
 This motivates the use of different online methods, specifically tailored to address these problems.
-On the one hand, lack of differentiability is usually tackled through zero-th order methods [39, 29], sub-gradient [24], proximal [22] or smoothing approaches [1].
-On the other hand, methods to address non-convexity in online learning often rely on the existence of a near optimal oracle [45, 3, 67, 75] or additional smoothness conditions [46, 32].
-All those results are derived for static regret [81] which compares the learned predictors to the best static strategy and is defined in Section˜2.
-A more challenging criterion, also described in Section˜2 is dynamic regret, introduced in [81] and later developed in [35, 12, 79, 78] among others, which takes into account the evolution of the environment.
-An alternative approach to obtain guarantees in the online non-convex setting is to weaken static regret to local regret [38, 6, 80, 36], which is the sum of the objective gradient norms evaluated in the iterates over time.
+On the one hand, lack of differentiability is usually tackled through zero-th order methods Héliou et al.
+(2020); Frezat et al.
+(2023), sub-gradient Duchi et al.
+(2011), proximal Dixit et al.
+(2019) or smoothing approaches Abernethy et al.
+(2014).
+On the other hand, methods to address non-convexity in online learning often rely on the existence of a near optimal oracle Kalai and Vempala (2005); Agarwal et al.
+(2019); Suggala and Netrapalli (2020); Xu and Zhang (2024) or additional smoothness conditions Lesage-Landry et al.
+(2020); Ghai et al.
+(2022).
+All those results are derived for static regret (Zinkevich, 2003) which compares the learned predictors to the best static strategy and is defined in Section˜2.
+A more challenging criterion, also described in Section˜2 is dynamic regret, introduced in Zinkevich (2003) and later developed in Hall and Willett (2013); Besbes et al.
+(2015); Zhao et al.
+; Zhao and Zhang (2021) among others, which takes into account the evolution of the environment.
+An alternative approach to obtain guarantees in the online non-convex setting is to weaken static regret to local regret Hazan et al.
+(2017); Aydore et al.
+(2019); Zhuang et al.
+(2020); Hallak et al.
+(2021), which is the sum of the objective gradient norms evaluated in the iterates over time.
 Minimizing local regret corresponds to encouraging convergence toward stationary points.
@@
 Notation.
-For a differentiable map $\varphi:\mathbb{R}^{m}\rightarrow\mathcal{Y}$ such that $\mathcal{Y}\subseteq\mathbb{R}$, $\nabla\varphi(x)$ denotes the gradient of $g$ at $x\in\mathbb{R}^{m}$ and $\nabla^{2}\varphi(x)$ denotes the Hessian of $g$.
+For a differentiable map $\varphi:\mathbb{R}^{m}\rightarrow\mathcal{Y}$ such that $\mathcal{Y}\subseteq\mathbb{R}$, $\nabla\varphi(x)$ denotes the gradient of $\varphi$ at $x\in\mathbb{R}^{m}$ and $\nabla^{2}\varphi(x)$ denotes the Hessian of $\varphi$.
 In the case where $\mathcal{Y}\subseteq\mathbb{R}^{d}$, $\nabla\varphi(x)$ denotes the Jacobian of $\varphi$.
@@
 $\min_{w\in\mathcal{W}}\ \langle\bar{g}_{t}(X_{t}),w\rangle\;,$ (1)
-where $X_{t}\in\mathcal{X}$ are covariates (possibly random) and $\mathcal{W}=\mathrm{Conv}(v_{1},\ldots,v_{K})$ is a bounded convex polytope of $\mathbb{R}^{d}$ with non-empty interior and vertices $\{v_{i}\}_{i=1}^{K}$.
-This feasible set appears naturally in many problems such as shortest-path [30], portfolio selection [47] or mixed strategy design in games [68].
-In (1), $\bar{g}_{t}:\mathcal{X}\to\mathbb{R}^{d}$, is a cost function, which satisfies for instance $\bar{g}_{t}(X_{t})=\mathbb{E}[Z_{t}\mid X_{t}]$ for some hidden state $Z_{t}\in\mathbb{R}^{d}$.
+where $\mathcal{W}=\mathrm{Conv}(v_{1},\ldots,v_{K})$ is a bounded convex polytope of $\mathbb{R}^{d}$ with non-empty interior and vertices $\{v_{i}\}_{i=1}^{K}$.
+This feasible set appears naturally in many problems such as shortest-path (Gallo and Pallottino, 1988), portfolio selection (Li and Hoi, 2014) or mixed strategy design in games (Syrgkanis et al., 2015).
+In (1), $X_{t}$ are random covariates and $\bar{g}_{t}:\mathcal{X}\to\mathbb{R}^{d}$, is a deterministic cost function, which satisfies for instance $\bar{g}_{t}(X_{t})=\mathbb{E}[Z_{t}\mid X_{t}]$ for some hidden state $Z_{t}\in\mathbb{R}^{d}$.
+At each period $t$, nature picks both a distribution for $X_{t}$ and a cost function $g_{t}$.
+This situation corresponds to the stochastic adversary setting (Rakhlin et al., 2011).
 Importantly, $X_{t}$ is revealed at the beginning of the round, but not $\bar{g}_{t}(X_{t})$ which is only available at the end of the round.
 While $\bar{g}_{t}(X_{t})$ is unknown to the decision-maker at the decision time, they have access to a family of models $g:\Theta\times\mathcal{X}\rightarrow\mathbb{R}^{d}$, parameterized by $\Theta\subset\mathbb{R}^{m}$, to predict it.
-Then, the general form of the decision-making dynamics we consider can be described as follows: at each round $t\in[T]$, for a horizon $T\in\mathbb{N}$, given the current parameter $\theta_{t}$,
+The general form of the decision-making dynamics we consider can be described as follows: at each round $t\in[T]$, for a horizon $T\in\mathbb{N}$, given the current parameter $\theta_{t}$,
   1. 1.
+    Nature picks a distribution for $X_{t}\in\mathcal{X}$ and a cost function $\bar{g}_{t}:\mathcal{X}\rightarrow\mathcal{Z}$.
+  2. 2.
     The decision-maker observes $X_{t}$ and compute its prediction as $g(\theta_{t},X_{t})$.
-  2. 2.
+  3. 3.
     Then, they take an action minimizing the resulting predicted cost
     $w_{t}=w^{\star}_{t}(\theta_{t})\in\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),w}}\right\rangle$ (2)
-  3. 3.
-    Finally, they observe $\bar{g}_{t}(X_{t})$ and update $\theta_{t+1}$ for the next round.
+  4. 4.
+    Finally, the decision-maker observes $\bar{g}_{t}(X_{t})$ and update $\theta_{t+1}$ for the next round.
 Formally, the decision-maker considers a joint process $(\theta_{t},w_{t})_{t\in[T]}$, starting from $\theta_{1}\in\Theta$ and a history $\mathcal{H}_{0}=\emptyset$, and defined by the following recursion.
-At round $t\in[T]$, the decision-maker takes the best action given the collected history $\mathcal{H}_{t-1}$, the current estimate $\theta_{t}$ and a new feature $X_{t}$ as in (2).
+At round $t\in[T]$, the decision-maker takes the best action given the current estimate $\theta_{t}$ and a new feature $X_{t}$ as in (2).
 Then, they observe $\bar{g}_{t}(X_{t})$ and update their history $\mathcal{H}_{t}=\mathcal{H}_{t-1}\cup\{(X_{t},\bar{g}_{t}(X_{t}),\theta_{t},w^{\star}_{t}(\theta_{t}))\}$.
@@
 Online decision-focused learning.
-From an online perspective, prediction-focused learning consists in selecting at each $t\in[T]$ an algorithm $\mathsf{Alg}_{t}$ to estimate $\operatorname*{\arg\!\min}_{\theta\in\Theta}\mathrm{R}_{t}(\theta)$, where $\mathrm{R}_{t}$ is a statistical risk based on the historical observations $\mathcal{H}_{t-1}$, typically chosen as the empirical risk $\mathrm{R}_{t}(\theta)=\sum_{i=1}^{t}\ell(\bar{g}_{i}(X_{i}),g(\theta,X_{i}))$, for some loss function $\ell$ (e.g., cross-entropy or squared error).
+From an online perspective, prediction-focused learning consists in selecting at each $t\in[T]$ an algorithm $\mathsf{Alg}_{t}$ to estimate $\operatorname*{\arg\!\min}_{\theta\in\Theta}\mathrm{R}_{t}(\theta)$, where $\mathrm{R}_{t}$ is a statistical risk based on the historical observations $\mathcal{H}_{t}$, typically chosen as the empirical risk $\mathrm{R}_{t}(\theta)=\sum_{i=1}^{t}\ell(\bar{g}_{i}(X_{i}),g(\theta,X_{i}))$, for some loss function $\ell$ (e.g., cross-entropy or squared error).
 Alternatively, we consider in this paper, the decision-focused learning approach, which selects parameters $\theta\in\Theta$ by directly minimizing the downstream objective instead of a risk function.
 Specifically, the decision-maker chooses $\mathsf{Alg}_{t}$ to solve $\operatorname*{\arg\!\min}_{\theta\in\Theta}\langle\bar{g}_{t}(X_{t}),w_{t}^{\star}(\theta)\rangle$ where $w_{t}^{\star}$ is defined in (2).
-From the previous formulation, we see that the decision-focused learning formulation corresponds to a bilevel optimization problem [19, 66, 44].
-We emphasize that no stationary assumptions are made on the process $\{X_{t},\bar{g}_{t}(X_{t})\}_{t\in[T]}$, allowing an adversarial selection of the features $X_{t}$ and cost functions $\bar{g}_{t}$.
+From the previous formulation, we see that the decision-focused learning formulation corresponds to a bilevel optimization problem (Colson et al., 2007; Sinha et al., 2017; Ji et al., 2021).
+We emphasize that no stationary assumptions are made on the process $\{X_{t},\bar{g}_{t}(X_{t})\}_{t\in[T]}$.
+In other words, the adversary may select any distribution for $X_{t}$ and any function $\bar{g}_{t}$ (as long as H˜2 below is satisfied).
 To fully accommodate this flexibility, we assess the optimality of $\{\theta_{t}\}_{t\in[T]}$ with two notions of regret.
@@
 denotes the loss incurred when taking an action based on the prediction parameter $\theta_{t}\in\Theta$.
-Inspired from [81], we consider the notions of static and dynamic regret to measure the effectiveness of a learning strategy $\mathopen{}\mathclose{{\left\{\theta_{t}}}\right\}_{t\in[T]}$, which are respectively:
+Following Zinkevich (2003), we consider the notions of static and dynamic regret to measure the effectiveness of a learning strategy $\mathopen{}\mathclose{{\left\{\theta_{t}}}\right\}_{t\in[T]}$, which are respectively:
 $\mathfrak{R}^{s}_{T}=\sum_{t\in[T]}f_{t}(\theta_{t})-\inf_{\theta\in\Theta}\sum_{t\in[T]}\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)}}\right]\quad\text{and}\quad\mathfrak{R}^{d}_{T}=\sum_{t\in[T]}F_{t}(\theta_{t})-\sum_{t\in[T]}\inf_{\theta\in\Theta}F_{t}(\theta),$ (4)
 where $\quad F_{t}:\theta\mapsto\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}}\right]$.
+In our dynamic regret, the sequence of actions is compared against a sequence of oracles, each minimizing the instantaneous loss.
+Without taking the conditional expectation over $\mathcal{H}_{t-1}$, each comparator could overfit to the specific realization of $X_{t}$ resulting in an unrealistically strong benchmark.
+By considering the conditional expectation of the loss, we effectively regularize the dynamic comparators, making them meaningful competitors.
 Note that our static regret compare to the best fixed strategy, with respect to the averaged losses.
 This notion makes sense here as we aim to control those regrets in expectation over the randomness of the process.
-In our dynamic regret, we consider the expectation of the loss given past information to prevent the dynamic comparator from overfitting the current realization of $X_{t}$.
 We make the following mild regularity assumptions for the rest of the analysis.
@@
 (i) $\Theta\subset\mathbb{R}^{m}$ is a compact and convex set with diameter $D_{\Theta}<\infty$.
-(ii) For any $\theta\in\Theta$ and $t\in[T]$, $\nabla_{\theta}g(\theta,X_{t})\in\mathbb{R}^{m\times d}$ is continuously differentiable and $\mathopen{}\mathclose{{\left\|\nabla_{\theta}g(\theta,X_{t})}}\right\|_{\mathrm{op}}\leqslant G<\infty$ almost-surely.
+(ii) for any $t\in[T]$, $\theta\mapsto g(\theta,X_{t})\in\mathbb{R}^{m\times d}$ is continuously differentiable almost surely and $\mathopen{}\mathclose{{\left\|\nabla_{\theta}g(\theta,X_{t})}}\right\|_{\mathrm{op}}\leqslant G<\infty$ for any $\theta\in\Theta$ almost-surely.
 (iii) For any $t\in[T]$, $\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|\leqslant D_{\mathcal{Z}}<\infty$ almost-surely.
-It is common in online learning to assume boundedness of the parameter space, model gradient and prediction space [14, 13].
+It is common in online learning to assume boundedness of the parameter space, model gradient and prediction space (Boyd, 2004; Bishop and Nasrabadi, 2006).
 Note that in the well-specified setting, i.e., $\bar{g}_{t}=g(\bar{\theta}_{t},\cdot)$ for some $\bar{\theta}_{t}$, these two former conditions automatically imply the latter.
-We emphasize however that we do not assume this realizability condition, in contrast to most of the literature on decision-focused learning [8].
+We emphasize however that we do not assume this realizability condition, in contrast to most of the literature on decision-focused learning (Bennouna et al., 2024).
 Without further assumptions on the sequence of costs $\bar{g}_{t}$ and models $g$, the problem can still be made arbitrarily hard.
-Therefore, we make an assumption coming from the classification [51, 70] and bandit literature [76, 59] about the margins of the cost function.
+Therefore, we make an assumption coming from the classification (Mammen and Tsybakov, 1999; Tsybakov, 2004) and bandit literature (Zeevi and Goldenshluger, 2009; Perchet and Rigollet, 2013) about the margins of the cost function.
 Recall that we denote by $\{v_{i}\}_{i=1}^{K}$ the vertices of $\mathcal{W}$.
-We define for any $x\in\mathcal{X}$, $u_{i}(\theta,x)=\mathopen{}\mathclose{{\left\langle g(\theta,x),v_{i}}}\right\rangle$ for $i\in[d]$.
+We define for any $x\in\mathcal{X}$ and $i\in[K]$, $u_{i}(\theta,x)=\mathopen{}\mathclose{{\left\langle g(\theta,x),v_{i}}}\right\rangle$.
 H 2.
-There exist $C_{0}>0$ and $\beta\in\mathopen{}\mathclose{{\left[0,1}}\right]$, such that almost surely for any $t\in[T]$, $\theta\in\Theta$ and $\varepsilon\in\mathopen{}\mathclose{{\left[0,1}}\right]$,
+There exist $C_{0}>0$ and $\beta\in(0,1]$, such that almost surely for any $t\in[T]$, $\theta\in\Theta$ and $\varepsilon\in\mathopen{}\mathclose{{\left[0,1}}\right]$,
 $\mathbb{P}\mathopen{}\mathclose{{\left(\inf_{j\neq I_{t}(\theta)}\{u_{j}(\theta,X_{t})-u_{I_{t}(\theta)}(\theta,X_{t})\}\geqslant\varepsilon\,\mid\mathcal{H}_{t-1}}}\right)\geqslant 1-C_{0}\varepsilon^{\beta}\;,$
@@
 For example, in Appendix˜A we show that it is satisfied when $g:(\theta,X)\mapsto X\theta$ where $X$ has i.i.d standard Gaussian columns.
-3 Online algorithms for decision-focused learning .
+3 Online algorithms for decision-focused learning
 On the need of regularization to get differentiation.
-From an online learning perspective, a natural approach to minimize the two regrets defined in (4) would be to update $\theta_{t+1}$ at each round $t$ using a variant of either the Follow-The-Leader algorithm [FTL, 45] or Online Gradient Descent [OGD, 81], applied to the objective function $f_{t}$ defined in (3), thus specifying the choice of $\mathsf{Alg}_{t}$.
+From an online learning perspective, a natural approach to minimize the two regrets defined in (4) would be to update $\theta_{t+1}$ at each round $t$ through an algorithm $\mathsf{Alg}_{t}$ being either a variant of either the Follow-The-Leader algorithm (FTL, Kalai and Vempala, 2005) or Online Gradient Descent (OGD, Zinkevich, 2003), applied to the objective function $f_{t}$ defined in (3).
 However, these algorithms are designed for single-level optimization problems and require access to (sub-)gradients of the objective.
@@
 Indeed, $w^{\star}_{t}$ minimizes a linear function over the convex polytope $\mathcal{W}$, thus $\theta\mapsto w_{t}^{\star}(\theta)$ ranges in the finite set of vertices $\{v_{1},\ldots,v_{K}\}$.
-To address this issue, following [74], we propose to add a regularizer $\mathcal{R}$ to the objective function in (2). Accordingly, we define for any $\theta\in\Theta$,
+To address this issue, following Wilder et al.
+(2019), we propose to add a regularizer $\mathcal{R}$ to the objective function in (2).
+Accordingly, we define for any $\theta\in\Theta$ and some $\alpha_{t}>0$
 $\tilde{w}_{t}(\theta)\in\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\{\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),w}}\right\rangle+\alpha_{t}\mathcal{R}(w)\}\;,$ (5)
 which is a regularized approximation of $w^{\star}_{t}$.
-When this surrogate is continuously differentiable—which holds for our choices of $\mathcal{R}$ in Section˜3—we can use $\nabla\tilde{w}_{t}(\theta)$ in our algorithmic routine.
+When this surrogate is continuously differentiable, which is the case for our choices of $\mathcal{R}$ in Section˜3, we can use $\nabla\tilde{w}_{t}(\theta)$ in our algorithmic routine.
 This approach amounts to minimizing the surrogate
@@
 In this case, $\tilde{w}_{t}(\theta)$ is differentiable in a neighborhood of any $\theta\in\Theta$ and admits a close-form Jacobian $\nabla\tilde{w}_{t}(\theta)$ by the implicit function theorem.
-A natural choice to force $\tilde{w}_{t}$ to remains in the interior of $\mathcal{W}$ is the corresponding log-barrier function:
+A natural choice to force $\tilde{w}_{t}$ to remains in the interior of $\mathcal{W}$ is the corresponding log-barrier function
 $\mathcal{R}:w\mapsto-\sum_{i=1}^{n}\ln(b_{i}-A_{i}^{\mathsf{\scriptscriptstyle T}}w)\;,$ (7)
@@
 The bi-level structure of the problem yields unexpected properties.
-In particular, even when the regularizer $\mathcal{R}$ is strongly-convex in (26)—which the case with both log-barriers and negative entropy—and the model $g$ is simple (such as $g(\theta,X)=X\theta$), $\tilde{f}_{t}:\theta\mapsto\mathopen{}\mathclose{{\left\langle\tilde{w}_{t}(\theta),\bar{g}_{t}(X_{t})}}\right\rangle$ may not be convex, but is lipschitz (see Lemmas˜5 and 1).
-This prevents us from directly using known online convex optimization algorithms [37].
-However, a recent line of research [3, 67, 75] has developed online algorithms in the non-convex case, provided the losses are Lipschitz-continuous—which is the case with log-barriers or negative entropy (see Lemmas˜1 and 5).
+In particular, even when the regularizer $\mathcal{R}$ is strongly-convex in (26), which the case with both log-barriers and negative entropy, and the model $g$ is simple (such as $g(\theta,X)=X\theta$), $\tilde{f}_{t}:\theta\mapsto\mathopen{}\mathclose{{\left\langle\tilde{w}_{t}(\theta),\bar{g}_{t}(X_{t})}}\right\rangle$ is lipschitz (see Lemmas˜5 and 1) but not convex in general.
+This prevents us from directly using known online convex optimization algorithms (Hazan, 2023).
+However, a recent line of research (Agarwal et al., 2019; Suggala and Netrapalli, 2020; Xu and Zhang, 2024) has developed online algorithms in the non-convex case, provided the losses are Lipschitz-continuous.
 These studies combine near-optimal oracles with perturbation techniques to establish sublinear bounds on the expected regret.
-In line with this literature, we also assume to have access to an approximate offline optimization oracle, a notion that appeared in a slightly modified form in Suggala and Netrapalli [67].
+In line with this literature, we also assume to have access to an approximate offline optimization oracle, a notion that appeared in a slightly modified form in Suggala and Netrapalli (2020).
 Definition 1.
-An $\xi$-approximate offline optimization oracle (or approximate oracle), adapted to a class $\mathcal{C}$ is a mapping $\mathbf{O}_{\xi}$ taking a function $f\in\mathcal{C}$ and outputting $\vartheta=\mathbf{O}_{\xi}(f)\in\Theta$ satisfying:
+An $\xi$-approximate offline optimization oracle (or approximate oracle), adapted to a class $\mathcal{C}$ is a mapping $\mathbf{O}_{\xi}$ taking a function $f\in\mathcal{C}$ such that $f:\Theta\rightarrow\mathbb{R}$, and outputting $\vartheta=\mathbf{O}_{\xi}(f)\in\Theta$ satisfying:
 $f(\vartheta)\leqslant\inf_{\theta\in\Theta}f(\theta)+\xi\;.$ (8)
-The notion of an approximate oracle formalizes that, in non-convex settings, we cannot rely on subroutines that provably attain a global minimizer of $\tilde{f}_{t}$ at time $t$ (as offline gradient descent would in convex problems).
-Instead, we must settle for local minimizers, whose quality is characterized by a parameter $\xi$—which vanishes in favorable loss landscapes.
-As a concrete example, consider the stochastic gradient descent (SGD) algorithm as $\mathbf{O}_{\xi}$.
-A large body of work has shown that, even when $\Theta$ is a class of deep neural networks, SGD can converge to local minimizers [31, 55, 58, 20], justifying that $\mathbf{O}_{\xi}$ is indeed an approximate oracle.
+This notion of an approximate oracle reflects the fact that, in non-convex settings, we cannot rely on subroutines that provably find a global minimizer of $\tilde{f}_{t}$ at time $t$ (as is possible in convex problems with offline gradient descent).
+Instead, we must be content with local minimizers.
+Their quality is measured by a parameter $\xi$, which becomes small in favorable loss landscapes.
+When $\mathcal{C}$ contains differentiable and Lipschitz functions, the stochastic gradient descent (SGD) algorithm can be used for the oracle $\mathbf{O}_{\xi}$.
+Indeed, a large body of work has shown that, even with models as complicated as deep neural networks, SGD can converge to local minimizers (Ghadimi and Lan, 2013; Mertikopoulos et al., 2020; Patel and Zhang, 2021; Cutkosky et al., 2023), justifying its use as an approximate oracle.
 A more detailed discussion is provided in Appendix˜B.
-We conclude this remark by mentioning that it is possible to conduct a non-convex analysis without an approximate oracle by considering the weaker notion of local regret defined as $\sum_{t\in[T]}\mathopen{}\mathclose{{\left\|\nabla F_{t}(\theta_{t})}}\right\|$; see [38, 6, 80, 36].
+We conclude this remark by mentioning that it is possible to conduct a non-convex analysis without an approximate oracle by considering the weaker notion of local regret defined as $\sum_{t\in[T]}\mathopen{}\mathclose{{\left\|\nabla F_{t}(\theta_{t})}}\right\|$; see (Hazan et al., 2017; Aydore et al., 2019; Zhuang et al., 2020; Hallak et al., 2021).
 However, such a definition is not meaningful in the DFL framework, where gradients are zero or undefined due to the structure of the problem.
@@
 Our first algorithm is displayed in Algorithm˜1, where $\mathrm{Exp}(\eta)$ refers to an exponential distribution with parameter $\eta>0$.
-It is inspired from the classic Follow-the-Leader approach [45], which consists in making a decision minimizing the sum of objective functions observed so far.
-When losses are non-convex, it is common to inject random noise at each step to regularize the total cost function, resulting in an approach known as Follow-the-Perturbed-Leader [43, 67].
+It is inspired from the classic Follow-the-Leader approach (Kalai and Vempala, 2005), which consists in making a decision minimizing the sum of objective functions observed so far.
+When losses are non-convex, it is common to inject random noise at each step to regularize the total cost function, resulting in an approach known as Follow-the-Perturbed-Leader (Hutter and Poland, 2005; Suggala and Netrapalli, 2020).
 We employ this strategy as $\mathsf{Alg}_{t}$ to update $\theta_{t+1}$ in Algorithm˜1.
@@
 7:end for
-Note that Algorithm˜1 requires determining $w^{\star}$, which can be efficiently achieved by classic numerical solvers in reasonable settings.
+Note that in line 3, the algorithm takes the best action $w^{\star}_{t}(\theta_{t})$ given the predicted cost $g(\theta_{t},X_{t})$ (the regularized action $\tilde{w}_{t}$ is only needed to compute $\tilde{f}_{t}$).
+In most practical settings, $w^{\star}_{t}$ can be computed efficiently with standard numerical solvers.
 However, we show in Appendix˜F that only being able to determine $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ such that $\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),\underline{w}_{t}(\theta_{t})}}\right\rangle-\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),w^{\star}_{t}(\theta_{t})}}\right\rangle\leqslant\kappa$ for some $\kappa>0$ only shifts the regret bounds of the next section by $\kappa$.
-We provide a theoretical guarantee on the convergence of Algorithm˜1 in the form of the following static regret bound.
+We provide a theoretical guarantee on the convergence of Algorithm˜1 with the following static regret bound.
 Theorem 1.
@@
 Furthermore, taking $\eta\propto m^{1/4}T^{-3/4}n^{-1/2}$ and $\alpha\propto m^{3/4}n^{1/2}T^{-1/4}$ yields:
-$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(m^{3/4}\sqrt{n}T^{-1/4}+\xi}}\right).$
+$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(m^{3/4}n^{3/2}T^{-1/4}+\xi}}\right).$
 The proof of Theorem˜1 can be found in Section˜G.2.
@@
 It is informative to compare our guarantee to existing bounds in the literature.
-[67], who also study non-convex online learning, achieves a rate of $\mathcal{O}(T^{-1/2})$ as long as their offline oracle satisfies $\xi=\mathcal{O}(T^{-1/2})$, at the cost of a degraded $m^{3/2}$ dependency on the dimension.
+Suggala and Netrapalli (2020), who also study non-convex online learning, achieves a rate of $\mathcal{O}(T^{-1/2})$ as long as their offline oracle satisfies $\xi=\mathcal{O}(T^{-1/2})$, at the cost of a degraded $m^{3/2}$ dependency on the dimension.
 Their faster rate in $T$ comes from the fact that they tackle a simpler, single-level problem as compared to our bi-level, non differentiable setting.
@@
 3.2 Decision-Focused Online Gradient Descent
-While DF-FTPL (Algorithm˜1) benefits from a converging static regret bound, the techniques involved in [67] are not enough to reach dynamic regret guarantees, which is particularly relevant in highly non-stationary environments where the optimal decisions may change significantly from one round to another.
-To this end, we go beyond the Follow-the-Leader approach and propose an original algorithm based on the celebrated Online Gradient Descent [81].
+While DF-FTPL (Algorithm˜1) benefits from a converging static regret bound, the techniques involved in Suggala and Netrapalli (2020) are not enough to reach dynamic regret guarantees, which is particularly relevant in highly non-stationary environments where the optimal decisions may change significantly from one round to another.
+To this end, we go beyond the Follow-the-Leader approach and propose an original algorithm based on the celebrated Online Gradient Descent (Zinkevich, 2003).
 This procedure, which we call DF-OGD, is presented in Algorithm˜2.
-In words, at each time $t\in[T]$, Algorithm˜2 first makes a decision $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ based on the current parameters $\theta_{t}\in\Theta$.
-Then, it observes $\bar{g}_{t}(X_{t})\in\mathcal{Z}$, determines the surrogate objective $\tilde{f}_{t}$ and computes one of its near minimizer via the offline oracle from Definition˜1.
-Next, to handle non-convexity, it determines the gradient $\nabla\tilde{w}_{t}(u_{t})$, evaluated in a point $u_{t}$ drawn uniformly at random in $\theta_{t},\vartheta_{t}]$.
-From there, it updates $\theta_{t+1}$ through the classical gradient step scheme of OGD.
+In Algorithm˜2, at each time $t\in[T]$, a decision $w_{t}^{\star}(\theta_{t})\in\mathcal{W}$ is made based on the current parameter $\theta_{t}\in\Theta$.
+The feedback $\bar{g}_{t}(X_{t})\in\mathcal{Z}$ is then observed and used to define the surrogate objective $\tilde{f}_{t}$.
+Then, a near-minimizer of this objective is computed using the offline oracle from Definition˜1.
+To address non-convexity, the gradient $\nabla\tilde{w}t(u_{t})$ is evaluated at a point $u_{t}$ drawn uniformly at random from $\theta_{t},\vartheta_{t}]$.
+Finally, the parameter is updated to $\theta{t+1}$ using the standard gradient step of OGD.
 […]
@@
 where $P_{T}=\sum_{t=1}^{T-1}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|$, $\mathbb{E}$ denotes the expectation on both data and the intrinsic randomness of DF-OGD and $\tilde{\mathcal{O}}$ contains polynomial dependency in $\ln(1/\alpha),\ln(\ln(d))$.
-Furthermore, assume $(t^{-1}(1+P_{t}))_{t\geqslant 1}$ is non-increasing almost surely with $P_{t}=\sum_{s=1}^{t}\mathopen{}\mathclose{{\left\|\vartheta_{s+1}-\vartheta_{s}}}\right\|$.
+Furthermore, assume $(t^{-1}(1+P_{t}))_{t\geqslant 1}$ is non-increasing almost surely with $P_{t}=\sum_{s=1}^{t-1}\mathopen{}\mathclose{{\left\|\vartheta_{s+1}-\vartheta_{s}}}\right\|$.
 Then, using $\alpha_{t}\propto n^{-1/2}t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto n^{-1/2}t^{-3/4}(1+P_{t})^{3/4}$ for any $t\in[T]$ leads to:
 $T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\mathbb{E}\mathopen{}\mathclose{{\left[\sqrt{n}(1+P_{T})^{1/4}T^{-1/4}}}\right]+\xi}}\right)\;.$
-The complete statement of the main bound, including constants is provided in Section˜G.3.
 Our dynamic regret bound naturally depends on $P_{T}$, which captures the problem’s variability by measuring the cumulative distance between approximate minimizers over time.
@@
 It is instructive to compare our guarantee with those established in recent studies on dynamic regret.
-For instance, [77] derive a $\mathcal{O}(\sqrt{T(P_{T}+1)})$ bound.
+For instance, Zhang et al.
+(2018) derive a $\mathcal{O}(T^{-1/2}\sqrt{T(P_{T}+1)})$ bound.
 However, their setting is not directly comparable to ours, as they consider a simpler single-level problem with differentiable and convex objectives.
 In contrast, our framework involves non-convex, non-differentiable losses due to the bi-level nature of decision-focused learning.
-More recently, [41] obtained a $\mathcal{O}((1+P^{\infty}_{T})^{1/3}T^{-1/3})$ bound, where $P^{\infty}_{T}:=\sum_{t=1}^{T}\|f_{t+1}-f_{t}\|_{\infty}$.
-Yet, this rate is achieved under substantially more favorable conditions: a single-level problem with losses that are strongly convex or Lipschitz, an additional assumption of “quasi-stationary” and a more challenging path involving the full landscapes of $f_{t}$s.
+More recently, Huang and Wang (2025) obtained a $\mathcal{O}((1+P^{\infty}_{T})^{1/3}T^{-1/3})$ bound, where $P^{\infty}_{T}:=\sum_{t=1}^{T}\|f_{t+1}-f_{t}\|_{\infty}$.
+Yet, this rate is achieved under substantially more favorable conditions: a single-level problem with losses that are strongly convex or Lipschitz, an additional assumption of “quasi-stationary” and a more challenging path involving the full landscapes of the $f_{t}$s.
 By comparison, our losses are neither Lipschitz (indeed, they may even be discontinuous due to the linearity of the lower-level problem) nor required to be stationary over time—the only restriction being H˜2 to hold.
@@
 4 Experiments
-In this section, we compare the performances of our algorithms DF-FTPL and DF-OGD to two important benchmarks, namely prediction-focused learning and SPO [25].
+In this section, we compare the performances of our algorithms DF-FTPL and DF-OGD to two important benchmarks, namely prediction-focused learning and SPO (Elmachtoub and Grigas, 2022).
 Setting.
 Our experimental setup is inspired by the knapsack example from Mandi et al.
-[54].
+(2024).
 More precisely, we consider a decision maker who must pick at each $t\in[T]$ an object $v_{t}\in\mathcal{V}$ among $K$ items denoted $\mathcal{V}=\{1,2,\ldots,K\}$ with respective costs $\bar{g}_{t}(X)=(\bar{g}_{t,1}(X),\ldots,\bar{g}_{t,K}(X))\in[0,1]^{K}$ depending on some covariates $X\in\mathcal{X}$.
@@
 This approach is formally described in Algorithm˜3 in Appendix˜E.
-Second, we compare our algorithms to an online version of the Smart Predict-then-Optimize (online SPO, see Algorithm˜4) approach from [25].
+Second, we compare our algorithms to an online version of the Smart Predict-then-Optimize (online SPO, see Algorithm˜4) approach from Elmachtoub and Grigas (2022).
 This algorithm introduces a differentiable and convex loss which upper-bounds the actual decision-focused loss.
@@
 We also mention the presence of additional numerical experiments in Appendix˜E, to study how model misspecification affects the relative performances of DFL and PFL.
-5 Conclusion.
+5 Conclusion
 Decision-focused learning offers a promising way to integrate prediction into decision making.
@@
 We believe that this work can be extended in several ways.
-First, we believe that alternative smoothing techniques could be successfully employed, such as Moreau-Yosida transform and proximal operators.
-Second, it would be valuable to investigate less adversarial environments, such as those involving i.i.d.
+First, it is plausible that faster rates (e.g., $T^{-1/2}$) could be obtained using more traditional online-learning techniques that bypass the bilevel framework, such as discretizing the parameter space and applying expert aggregation over the resulting bins.
+However, such approaches would likely incur a prohibitive dependence on the parameter dimension $m$ (e.g., exponential in $m$).
+Exploring this direction remains however an interesting avenue in future work.
+Second, we believe that alternative smoothing techniques could be successfully employed, such as Moreau-Yosida transform and proximal operators, to adress the more challenging of a general convex set $\mathcal{W}$.
+Third, it would be valuable to investigate less adversarial environments, such as those involving i.i.d.
 data, to explore whether stronger theoretical guarantees could be obtained and whether novel algorithmic designs might emerge.
 Finally, implementing our method in more ambitious experimental settings would be of great interest from a practical perspective.
-Appendix A H˜2 is satisfied in the Gaussian linear case.
+Acknowledgment
+Funded by the European Union (ERC, Ocean, 101071601).
+M.H.
+is partially supported by the European Research Council Starting Grant DYNASTY – 101039676.
+Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency.
+Neither the European Union nor the granting authority can be held responsible for them.
+Appendix A H˜2 is satisfied in the Gaussian linear case
 In this section, we prove that H˜2 holds true when $\mathcal{W}$ is the simplex of $\mathbb{R}^{d}$, $g$ is a linear model and $X$ has columns distributed according to a Gaussian distribution.
@@
 Even in the context of deep networks, it is plausible that $\mathbf{O}$ converges to a local minimizer.
-Indeed, a large body of work has analyzed SGD in non-convex settings, establishing convergence to stationary points either in expectation [31] or almost surely [55, 58, 20].
+Indeed, a large body of work has analyzed SGD in non-convex settings, establishing convergence to stationary points either in expectation [Ghadimi and Lan, 2013] or almost surely [Mertikopoulos et al., 2020, Patel and Zhang, 2021, Cutkosky et al., 2023].
 Such stationary points may correspond to local/global minima or saddle points.
@@
 [2020] proved that the trajectories of SGD almost surely avoid all strict saddle manifolds—i.e., sets of critical points $x$ where the Hessian has at least one negative eigenvalue.
-These manifolds include connected families of non-isolated saddle points, a phenomenon common in the loss landscapes of overparametrized neural networks [48].
+These manifolds include connected families of non-isolated saddle points, a phenomenon common in the loss landscapes of overparametrized neural networks [Li et al., 2018].
 Beyond SGD, similar guarantees extend to more general methods.
-In particular, the stochastic Riemannian Robbins–Monro method (a broad template encompassing various algorithms) has been shown to converge almost surely to local or global minima [40].
-Appendix C Supplementary results on the simplex.
+In particular, the stochastic Riemannian Robbins–Monro method (a broad template encompassing various algorithms) has been shown to converge almost surely to local or global minima [Hsieh et al., 2023].
+Finally, we highlight that, in most practical industrial settings, solving the downstream optimization problem is typically far more computationally demanding than updating the prediction model.
+It is therefore reasonable to assume that the oracle call represents only a small fraction of the overall computational cost.
+Appendix C Supplementary results on the simplex
 We now provide convergence guarantees for Algorithms˜1 and 2 when $\mathcal{W}$ is the simplex. To do so we choose another regularizer than done in the main document.
@@
 $\mathcal{W}_{0}=\mathopen{}\mathclose{{\left\{w\in\mathbb{R}^{d}:w\succcurlyeq 0,\;{\bf 1}^{\mathsf{\scriptscriptstyle T}}w=1}}\right\}\;.$
-This setting encompasses various decision-making scenarios such as portfolio selection [47] or mixed strategy design in multiplayer games [68]. In this case, we choose the negative entropy
+This setting encompasses various decision-making scenarios such as portfolio selection Li and Hoi [2014] or mixed strategy design in multiplayer games Syrgkanis et al.
+[2015].
+In this case, we choose the negative entropy
 $\mathcal{R}_{0}:w\mapsto\sum_{i\in[d]}w_{i}\ln(w_{i})\;,$
@@
 Approximate Optimization Oracle.
-The recent work of [67] proposed an online algorithm for nonconvex losses $(\ell_{t})_{t\in[T]}$ with static regret guarantees.
+The recent work of Suggala and Netrapalli [2020] proposed an online algorithm for nonconvex losses $(\ell_{t})_{t\in[T]}$ with static regret guarantees.
 They rely on an approximate optimization oracle $\mathcal{O}$ which takes as input any function $\ell$, a $d$-dimensional vector $\sigma$ and returns an approximate minimizer of $\ell-(\sigma,\cdot)$.
@@
 Follow The Perturbed Leader.
-Given access to an ($\xi,\chi$ )-approximate optimization oracle, [67] study the FTPL algorithm which is described by the following recursion.
+Given access to an ($\xi,\chi$ )-approximate optimization oracle, Suggala and Netrapalli [2020] study the FTPL algorithm which is described by the following recursion.
 Starting from $\hat{\mu}_{1}$, at each time steps $t$, $\hat{\mu}_{t}\in\mathcal{K}$ is updated as follows:
@@
 Following this update route, the addition of an exponential noise allowing to handle the non-convexity of the losses, they reach the following static regret bound.
-Proposition 4 (Theorem 1 of [67]).
+Proposition 4 (Theorem 1 of Suggala and Netrapalli [2020]).
 Let $D$ be the diameter of $\mathcal{K}$.
@@
 For any fixed $\eta$, FTPL (Equation˜14) with access to a $(\xi,\chi)$-approximate" optimization oracle satisfies the following static regret bound:
-$\frac{1}{T}\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\inf_{\mu}\sum_{t=1}^{T}\ell_{t}(\mu)\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}DL^{2}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi mL}}\right),$
+$\frac{1}{T}\mathopen{}\mathclose{{\left(\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\inf_{\mu}\sum_{t=1}^{T}\ell_{t}(\mu)}}\right)\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}DL^{2}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi mL}}\right),$
 where $\mathbb{E}$ denotes the expectation over $\sigma_{1}\cdots\sigma_{T}$.
 Appendix E Additional experimental material
-The motivating example from [54].
+The motivating example from Mandi et al. [2024].
 […]
 0112233445501122334455Value of Item 1Value of Item 2
-Figure 2: Figure 2 from [54].
+Figure 2: Figure 2 from Mandi et al. [2024].
 Details on the experimental setup.
@@
 8:end for
-Appendix F Regret bound with an approximate optimal solution.
+Higher dimension problem.
+Finally, the following figure shows the performance of our algorithms in a higher dimension problem, with $K=80$ items.
+We see that our algorithms still significantly beat the baselines, suggesting that they remain effective in moderately high dimension problems.
+Figure 4: Experiment results with 80 items.
+Appendix F Regret bound with an approximate optimal solution
 In this section, we show that when we cannot access $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ but only $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ such that $\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),\underline{w}_{t}(\theta_{t})-w^{\star}_{t}(\theta_{t})}}\right\rangle\leqslant\kappa$ for $\kappa>0$, our regret bounds are only shifted by $\kappa$.
@@
 Appendix G Proofs
-G.1 Proofs of preliminary lemmas.
+G.1 Proofs of preliminary lemmas
 G.1.1 Simplex
@@
 Assume H˜2, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $\theta\in\Theta$ and $t\in[T]$, for any $\theta\in\theta$
-$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)}}\right)\;.$
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha_{t}+2(\alpha_{t}\ln(d))^{\beta}C_{0}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}\;.$
 Proof.
@@
 $\mathbb{E}\mathopen{}\mathclose{{\left[\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\,\mid\,\mathcal{H}_{t-1}}}\right]$ $\leqslant\alpha_{t}+2\alpha_{t}\ln(d)C_{0}\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon\;.$
-We now proceed to bound $\varphi(\beta)=\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon$ for $\beta\in(0,1)$. With $a=\alpha_{t}\ln(d)$ and $b=2\ln(d)$, we have:
-$\varphi(\beta)=(1-\beta)^{-1}(a^{\beta-1}-b^{\beta-1})=(1-\beta)^{-1}[e^{-(1-\beta)\ln(a)}-e^{-(1-\beta)\ln(b)}]\;,$
-and since for any $x>0,\,1-x\leqslant e^{-x}\leqslant 1-x+x^{2}$:
-$\varphi(\beta)$ $\leqslant(1-\beta)^{-1}[1-(1-\beta)\ln(a)+(1-\beta)^{2}\ln^{2}(a)-1+(1-\beta)\ln(b)]$
-$=\ln\mathopen{}\mathclose{{\left(\frac{b}{a}}}\right)+(1-\beta)\ln^{2}(a)\;.$
-Substituting in the values of $a$ and $b$ yields the desired result. ∎
+We now proceed to bound $\varphi(\beta)=\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon$ for $\beta\in(0,1)$. With $a=\alpha_{t}\ln(d)$ and $b=2\ln(d)$, we have
+$\varphi(\beta)=(1-\beta)^{-1}(a^{\beta-1}-b^{\beta-1})=(1-\beta)^{-1}\bigl[e^{-(1-\beta)\ln(a)}-e^{-(1-\beta)\ln(b)}\bigr]\;.$
+Factoring out $e^{-(1-\beta)\ln(a)}=a^{\beta-1}$ yields
+$\varphi(\beta)=a^{\beta-1}\,(1-\beta)^{-1}\Bigl[1-e^{-(1-\beta)\ln(b/a)}\Bigr]\;.$
+Since for any $x\in\mathbb{R}$ one has $1-e^{-x}\leqslant\min\{x,1\}$, we obtain
+$\varphi(\beta)\leqslant a^{\beta-1}\,(1-\beta)^{-1}\min\Bigl\{(1-\beta)\ln(b/a),\,1\Bigr\}=a^{\beta-1}\min\Bigl\{\ln(b/a),\,(1-\beta)^{-1}\Bigr\}\;.$
+Substituting the values of $a$ and $b$ finally gives
+$\varphi(\beta)\leqslant(\alpha_{t}\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}\;,$
+which concludes the proof.
 G.1.2 General polytope
@@
 Since $h_{t}(\tilde{w}_{t}(\theta),\theta)=\min_{w\in\mathcal{W}_{1}}h_{t}(w,\theta)$, we deduce that $\tilde{w}_{t}(\theta)\in\text{int}(\mathcal{W}_{1})$.
-It follows from the first order condition and the implicit function theorem [21, Theorem 2] that $\tilde{w}_{t}:\theta\mapsto\tilde{w}_{t}(\theta)$ is differentiable, and
+It follows from the first order condition and the implicit function theorem [de Oliveira, 2014, Theorem 2] that $\tilde{w}_{t}:\theta\mapsto\tilde{w}_{t}(\theta)$ is differentiable, and
 $\nabla\tilde{w}_{t}(\theta)$ $=-(\nabla^{2}_{ww}h_{t}(\tilde{w}_{t}(\theta),\theta))^{-1}\nabla^{2}_{\theta w}h_{t}(\tilde{w}_{t}(\theta),\theta)$
@@
 Assume H˜2, $\mathcal{W}=\mathcal{W}_{1}$ and $\mathcal{R}=\mathcal{R}_{1}$. Then for any $t\in[T]$ and $\theta\in\Theta$,
-$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha_{t}+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha_{t}(\alpha_{t}\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}$
 Proof.
@@
 $\tilde{w}_{t}(\theta)=\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),w}}\right\rangle-\alpha_{t}\sum_{i=1}^{d}\ln(b_{i}-A_{i}w)\;,$ (26)
-which is uniquely determined by strong convexity of the objective. We know by [14, page 566] that:
+which is uniquely determined by strong convexity of the objective. We know by [Boyd, 2004, page 566] that:
 $\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle-\inf_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{\left\langle g(\theta,X_{t}),w}}\right\rangle\leqslant n\alpha_{t}\;,$
@@
 $\mathbb{E}_{u_{1:T}}\mathopen{}\mathclose{{\left[\sum_{t=1}^{T}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+2D_{\Theta}}{2\eta_{T}}\sum_{t=1}^{T}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|+\sum_{t=1}^{T}\frac{\eta_{t}K_{t}^{2}}{2}.$
-G.2 Proof of Theorem˜1.
+G.2 Proof of Theorem˜1
 See 1
@@
 First, notice that, by Lemma˜6, and because for any $t,\alpha_{t}=\alpha$ we have for all $t$:
-$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$ (39)
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha(\alpha\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right),\,(1-\beta)^{-1}\Bigr\}\;.$ (39)
 Second, taking the expectation (denoted by $\mathbb{E}$) over $(\sigma_{t})_{t\in[T]}$ and $X_{1},\cdots X_{T}$ on both sides gives:
-$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]+2D_{\mathcal{Z}}\alpha T\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$ (40)
-To bound the first term in the right-hand-side of Equation˜40. by the definition of conditional expectation, we have:
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]+2D_{\mathcal{Z}}T\mathopen{}\mathclose{{\left(\alpha+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha(\alpha\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right),\,(1-\beta)^{-1}\Bigr\}}}\right)\;.$ (40)
+We now bound the first term in the right-hand-side of Equation˜40. By the definition of conditional expectation, we have:
 $\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]$ $=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}}\right\rangle-\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$
-Then,
+and then,
 $\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]$ $\leqslant\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}}\right].$
@@
 $T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]$ $\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi}}\right)$
-$+2D_{\mathcal{Z}}\alpha\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right),$
+$+2D_{\mathcal{Z}}T\mathopen{}\mathclose{{\left(\alpha+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha(\alpha\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right),\,(1-\beta)^{-1}\Bigr\}}}\right)$
 $=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi+\alpha n}}\right).$
@@
 $\leqslant\xi.$
-Then, taking the $\sup$ over $\Theta$ for each summand of the first sum of (42) (which is valid as $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable), defining by $\mathrm{Reg}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]$, and noticing for (44) that $\inf_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)]-\inf_{\theta\in\Theta}\mathbb{E}_{t}[\tilde{f}_{t}(\theta)]\leqslant\sup_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)-\tilde{f}_{t}(\theta)]$ leads to:
-$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right]+\xi T$
+Then, taking the $\sup$ over $\Theta$ for each summand of the first sum of (42) (which is valid as $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable), defining by $\mathrm{Reg}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]$, and noticing for (44) that $\inf_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)]-\inf_{\theta\in\Theta}\mathbb{E}_{t}[\tilde{f}_{t}(\theta)]\leqslant|\inf_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)]-\inf_{\theta\in\Theta}\mathbb{E}_{t}[\tilde{f}_{t}(\theta)]|\leqslant\sup_{\theta\in\Theta}|\mathbb{E}_{t}[f_{t}(\theta)-\tilde{f}_{t}(\theta)]|\leqslant\sup_{\theta\in\Theta}\mathbb{E}_{t}[|f_{t}(\theta)-\tilde{f}_{t}(\theta)|]$ leads to:
+$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right|}}\right]+\xi T$
 […]
@@
 First, by Lemma˜6, we know have for any $t$:
-$\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$
+$\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]\leqslant\alpha_{t}+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha_{t}(\alpha_{t}\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}\;.$
 Thus combining the two last equations and taking the expectation (denoted by $\mathbb{E}$) over $(u_{t})_{t\in[T]}$ and $X_{1},\cdots X_{T}$ on both sides gives:
-$\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}_{T}}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)+\xi T\;.$ (46)
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}_{T}}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\mathopen{}\mathclose{{\left(\alpha_{t}+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha_{t}(\alpha_{t}\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}}}\right)+\xi T\;.$ (46)
 Now, to bound the first term in Equation˜46 note that the definition of conditional expectation implies:
@@
 $T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{\left[P_{T}}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{\left\|b-Aw}}\right\|_{\infty}^{2}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}}\right)^{2}$
-$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}(\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d))]+\xi\;.$ (48)
+$+2D_{\mathcal{Z}}\sum_{t\in[T]}\mathopen{}\mathclose{{\left(\alpha_{t}+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha_{t}(\alpha_{t}\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}}}\right)+\xi\;.$ (48)
 Rewriting Section˜G.3 using $\tilde{\mathcal{O}}$ concludes the proof.
@@
 $+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]-\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}}\right\rangle}}\right]$ (51)
-Then, taking the $\sup$ over $\Theta$ in Equations˜49 and 51, and defining
+Then, taking absolute values, using the triangular inequality and taking the $\sup$ over $\Theta$ in Equations˜49 and 51, as well as defining
 $\mathrm{Reg}^{\mathrm{S}}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}}\right\rangle}}\right]-\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right],$
 in Equation˜50 yields for any $\theta$:
-$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$
+$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right|}}\right]$
 […]
@@
 First, notice that, by Lemma˜2, and because for any $t,\alpha_{t}=\alpha$ we have for all $t$:
-$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right)\;.$ (52)
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha+2(\alpha\ln(d))^{\beta}C_{0}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right),\,(1-\beta)^{-1}\Bigr\}\;,\;.$ (52)
 Second, taking the expectation (denoted by $\mathbb{E}$) over $(\sigma_{t})_{t\in[T]}$ and $X_{1},\cdots X_{T}$ on both sides gives:
-$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]+2D_{\mathcal{Z}}\alpha T\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right)\;.$ (53)
+$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]+2D_{\mathcal{Z}}T\mathopen{}\mathclose{{\left(\alpha+2(\alpha\ln(d))^{\beta}C_{0}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right),\,(1-\beta)^{-1}\Bigr\}}}\right)\;.$ (53)
 To bound the first term in the right-hand-side of Equation˜53. by the definition of conditional expectation, we have: